A Complete Bibliography of Publications in International Journal on Digital Libraries

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

07 May 2024 Version 1.00

Title word cross-reference

3 [DVP09]. n [TLJ^{+}02]. R [Bar21]. R^{4} [Bar21]. × [Hie00]. X^{2} [MSW^{+}05].

-D [DVP09]. -gram-based [TLJ^{+}02].

1 [MTLH17]. 14th [FRS10]. 1787 [CARM20]. 17th [AP14]. 19 [GCC24].

2013 [SF15]. 2014 [KR15]. 2015 [KMW16]. 2016 [FW19, JCRK18]. 2017 [MMJ20, TK20]. 2020 [ADH^{+}22, JMI^{+}22]. 2021 [ISK^{+}24]. 2022 [JKPC24]. 20th [FKRN19]. 21 [BDvdS06]. 25964 [AIL16].

3.0 [SOM22]. 3D [BKS^{+}06, CVHS20]. academia [CVDT22]. academic [ASFF19, KH17, KHVT23]. access [BJWW04, CDN^{+}98, FBPC17, GCC24, GMP^{+}23, HB16, HBC^{+}19, OHBR15, WLC^{+}19]. Accessibility [GMP^{+}23, TMTA07]. Accessing [GRB^{+}05]. ACM [FW19, MMJ20, KSH^{+}99]. ACM/IEEE [FW19, MMJ20]. ACM/IEEE-CS [FW19, MMJ20]. acquisition [DFT^{+}12, TMTA07]. across [GN16, WBL^{+}07a, WBL^{+}07b]. ActiveGraph [MHML05]. activities [CLvHM18, GABD07]. activity [PETS12, VC17]. activity-based [PETS12]. actors [BKS^{+}05]. acyclic [AN15]. adaptation [NMT10, SAPN00]. Adaptive [ZYZ^{+}00, CESF99, ITSY06, KJ05, KTS00]. Adding [EMS^{+}98]. aDORe [VCH08]. advanced [BJWW04, Pap06, TKM15]. advances [SC23]. Advancing [Sch04, VC17]. aesthetic [HCH04]. affiliation [BHD22]. affordable [NPS05]. African [NA13]. agenda [CBRCW05]. agent [ADT00, DLR02, GCAN00, YRW^{+}05]. agent-based [ADT00, GCAN00]. aggregation [FRH^{+}20, HMG20, Kel24]. aggregators [Kel24]. AI [BGAE24]. AI-based [BGAE24]. Alexandria [FFF^{+}00, GT17]. Alexandrian [NS00]. algorithmic [Lee07]. algorithms [CIMR12, NHP^{+}21]. Alignment [jCZhC16, KSG10, OFM^{+}00, WLC^{+}19]. Alliance [GMP^{+}23]. ALOCOM [VD08]. alone [Zim07]. American [Buc23]. Amharic [ARA23]. among [KMM19]. analyses [MSS12, RM19]. Analysing [BHS^{+}22, IMR12]. analysis [BDAB18, BM20, BM05, CSL^{+}97, DWD12, EWSB20, FISN20, GMPT00, HMG20, ISY20, JKT19, JCMB00, KGJK14, KF98, KB22, KAMKJ19, KWK^{+}22, KGE23, KHVT23, LLSG08, PMKD19, RZ20, SFT22]. Analytical [SB23]. Analyzing [KMM19, SJ21]. Anatomy [ASFF19]. ancient [FM17, HLZ05]. anisotropic [BM20]. annotated [JP05]. annotating [KAMKJ19, dRF18]. annotation [BP19, CC04, HJK^{+}09, SM12, TMTA07, TPMG15, WGZ^{+}06]. annotations [ABDF07, HJK^{+}09]. anonymity [BDY02]. anonymity-revoking [BDY02]. answering [LSC19, SGM^{+}22]. answers [LSC19]. API [LCK^{+}18]. API-based [LCK^{+}18]. Appearance [CLM06]. applicability [TNZR13]. application [BJWW04, MTLH17, SPLF02]. Applications [KW19, DLR02, GST19, LKR02, MTC^{+}16, NAO08, PAM09, WHK^{+}97]. approach [ACZ06, ADL18, AFMF21, BM16, BBF^{+}10, BDF^{+}15, BGAE24, Bor04, Fig21, Fis21, HC06b, KGE23, OHBR15, SPLF02, SAS^{+}14, TMBC22, TLJ^{+}02, TPMG15, Wan99, dSCGL20]. approaches [BP19, FJ20, MSS12, MSHC04, SH00]. Approximate [OSD23, CIMR12]. approximately [TLJ^{+}02]. ARCA [BCMC23]. archaeological [Fig21, MR21, MBS21, Nic17a, TMBC22]. archaeology [HGG^{+}21]. Archéologie [MR21]. architectural [MKFK18, SSNM97]. architecture [BCGP97, DBG^{+}16, FFF^{+}00, GCAN00, GBR17, LPSW06, PCDM^{+}22a, PCDM^{+}22b, RBdP^{+}15, TCK^{+}97, VCH08]. architectures [TS20]. archivability [BKWN16, BM16]. archival [Wau06]. archive [AN15, AWNV14, ANV^{+}16, FMN18, HR97, HLZ05, MT00, STvO^{+}18, ST04, VC17, dBOS16, AAWN14, Cap07, MMK^{+}19]. archives [AWN16b, Aya22, Can12, EKV^{+}10, FHKN20, FISN20, FBPC17, GMPT00, GRD20, HK12, MKPI^{+}22, OSCF09, RBdP^{+}15, SP04, SJ09, YRW^{+}05, RPM^{+}15]. ArchiveWeb [FMN18]. Archiving [BG12, CGM05, CGS17, FKX18, LCK^{+}18, RWS06]. art [BWSH12, jCZhC16, MKFK18]. artefacts [CVHS20]. Artexte [NMT10]. article [BACB20, CSSS21]. articles [HDG19, KGJK14, KBFG19, PCDM^{+}22a, PCDM^{+}22b, SGM^{+}22]. Artificial [FD98]. arts [NMT10, PEC^{+}15]. Asian [LCN^{+}04]. aspect [KGE23]. assess [CCGS14]. Assessing [BNLD22, LSC19, PB20]. assessment [DSGF06, FRT21, Kor16]. AssetCollector [STP07]. assets [BDvdS06, OHBR15, STP07]. assignment [RD10]. assist [RZ20]. Assisting [TKM15]. assurance [Fad02]. atlas [LGT97]. attention [CSMW22]. attitudes [CNHB16]. attribute [SSK^{+}24]. attributes [BB17]. attributions [Fig21]. Audio [HC06a, Can12, HB16, OSCF09]. Audio-based [HC06a]. audiovisual [MNZD22a, MNZD22b, dBOS16]. audit [SJ09]. augmenting [WBL^{+}07a, WBL^{+}07b]. authentication [Sae00]. authenticity [DVP09]. author [SGLF15]. authoring [KVZ17]. authority [RGÓ^{+}15]. authorization [KFW04]. Automated [HDG19, MSS12]. Automatic [ADL18, FAM00, GMPT00, MG06, NAO08, OD09, TS20, TPMG15]. Automatically [LRT21, BGAE24]. Automating [KVZ17]. autonomously [LSC19]. Avoiding [JNV18]. award [SF15]. award-nominated [SF15]. aware [BTN12, HMG20, KPK^{+}24, LLS16].

ba [RNFD16]. Babel [GT17]. Bag [Whi18]. balancing [KTS00]. based [AS02, ADT00, BNC12, dSBG21, BCGM00, BDF^{+}15, BGAE24, Cap07, CLvHM18, DIJ^{+}06, DSGF06, DBBK00, FRD23, Fis21, Fre14, GCAN00, HMG20, HC06a, HN07, HBC^{+}19, JDAD22, JGL^{+}97, KG21, KRZ18, KDP14, KD21, KI04, KNH97, KGE23, LA00, LCK^{+}18, MBM19, MG06, MMK^{+}19, MRY06, PETS12, PKK18, PRS97, RTM05, RD10, RM19, RZ20, SH00, Sat06, SB14, SZOS22, SMD00, SDLÅ14, SPO08, SM21, TBC^{+}15, TLJ^{+}02, TCK^{+}97, WWFW98, WIMM24, dBOS16]. baselines [BHD22]. bases [NHP^{+}21]. behavior [ASFF19]. Behaviour [KPC^{+}21, SB04]. behaviours [CNHB16]. Bench [WRSB23]. benchmark [OOKS13, WRSB23]. Benchmarking [ADT^{+}19]. beneath [DBT^{+}15]. BERT [JDAD22]. BERT-based [JDAD22]. Best [ADH^{+}22, ABDT^{+}17, TK20]. between [GMA12, JBFNMD15, OOKS13]. Beyond [CBC^{+}23a, CBC^{+}23b, DBT^{+}15, Kel24, MMK^{+}17, WPS^{+}04]. Bias [HMG20, HDG19, STvO^{+}18]. Bias-aware [HMG20]. biblio [WIMM24]. biblio-record-based [WIMM24]. bibliographic [ADT^{+}19, KWK^{+}22, RZ20, SB14]. bibliometric [MFC^{+}18]. bibliometric-enhanced [MFC^{+}18]. biodiversity [dSTMGF06]. BioKleisli [DOTW97]. biomedical [BKG^{+}97, DOTW97, NHP^{+}21, WHK^{+}97]. biosphere [DBT^{+}15]. BIRNDL [MFC^{+}18]. Blogosphere [HTR15]. books [CLM06, HB16, YKP20]. Boolean [JMS99]. borders [WPS^{+}04]. both [Sae00]. Brazilian [FCR20]. Bridging [JBFNMD15]. briefing [ACM^{+}98, Mar04]. Bring [Gri23a]. Bringing [Gri23b, GGM23]. Broadcasting [MMK^{+}19]. BroadScale [YSZ06]. Broker [Wit07]. browse [SDLÅ14]. browser [MTT15]. Browsing [KK99, LKA^{+}17, MSHC04, NMWP99]. BTG [AIL16]. BTI [AIL16]. BTP [AIL16]. Building [FHKN20, FS00, GABD07, MBS21, BKG^{+}09, MBM19]. buildings [RNFD16]. business [MAC^{+}15]. businesses [Fad02].

C [FAM00]. C-value [FAM00].

C-value/NC-value [FAM00]. Can [MM13]. capacity [LKR02]. Capisco [HBC^{+}19]. capital [CCCJ10]. capture [MAC^{+}15, SPLF02]. capturing [BNC12]. card [BDY02]. cards [GT17]. care [KM13]. caring [Bar07]. case [BPL22, BG12, BGL05, CPR23, FMNH04, HSPQ04, LHD08, PETS12, RSF08, RZ20, STM^{+}20, SB04, SFTD14, VM17, Wu16, dSRL19, dRF18]. cases [BHS^{+}22]. catalogs [CRK18, Sto00]. categorization [DWD12, SH00, dSCGL20]. category [KGE23]. CDX [ANV^{+}16]. centered [Aya22, BDF^{+}15, Pap06]. centre [HWT17]. centric [FISN20, GRD20]. certify [SJ09]. CH [WRSB23]. CH-Bench [WRSB23]. chain [CCGS14]. challenge [Fad02, WTH17]. Challenges [Di 23, HWT17, KYK23, TBC^{+}15]. Champions [NA13]. Change [GTP^{+}99, EWSB20]. changes [BG12]. changing [CS98]. chaos [FD98]. character [ARA23]. Characterising [WHCF20]. Characteristics [AWN16a, Hus19]. characterizing [MSS12]. Chi [HCHH05]. Children [VH21]. Chinese [jCZhC16, ZG04]. Choice [BACB20]. CIDOC [CVHS20, FM17, HDE17, LHD08, MTLH17, Nic17a, Nic17b, NH17, SFTD14]. Citation [CPR23, FJ20, CG18, CRK18, KH17, LMZ^{+}18, NAO08, SK15]. citations [LMZ^{+}18, SFT22]. citeable [HWT17]. cited [Whi18]. citizens [SSK^{+}24]. city [SSK^{+}24]. CL [JCRK18]. claims [PB20]. classification [CSMW22, Ke15, LMZ^{+}18, LKA^{+}17, MG06, SM21, XI24]. Classifier [SOM22]. classifiers [JDAD22]. classifying [SOM22, SZOS22, SI21]. classroom [GWB^{+}08, WPS^{+}04]. CLEAR [BM16]. click [SDLÅ14]. clinically [BKG^{+}97]. clinically-relevant [BKG^{+}97]. cluster [LA00, SGLF15]. clustered [ZYZ^{+}00]. clustering [Ke15, MRY06, SDLÅ14]. CNN [SZOS22]. CNN-based [SZOS22]. co [Whi18]. co-cited [Whi18]. Cognitive [CMM04]. Coins [KBG23]. collaboration [CCvdM08, MBS21, VSVS08]. Collaborative [CMM^{+}07, RAA^{+}12, SASV12, ST04]. collaboratively [FMN18]. collaboratory [TBF^{+}04]. COLLATE [TBF^{+}04]. collecting [LCK^{+}18]. collection [JP05, KBG23, KAMKJ19, SDLÅ14]. collections [AOP18, Buc23, DFT^{+}12, FMN18, GRD20, JDAD22, MSW^{+}05, MMK^{+}19, NPD20, NMWP99, PB20, SLM04, STP07]. colors [BDAB18]. Columbia [ACM^{+}98]. combination [SGLF15]. combining [EWSB20]. comments [MHTH15]. commerce [DLR02, Fad02, LKR02, MPLR02, SPLF02]. Commercial [Par04]. Commons [MHTH15]. Communication [CMT13]. community [CCCJ10, KMM19, LSC19]. Comparing [KBFG19, SSK^{+}24]. Comparison [RD10, BKS^{+}06, ISY20, MMT04]. complex [LPSW06, NV06]. component [SPO08]. component-based [SPO08]. components [ABDT^{+}17]. composite [HC06b]. composition [AIL16]. comprehensive [NPD20, SK15]. compression [AS02]. Computational [LMZ^{+}18, dlCAS^{+}08, MSS12]. Computer [LW09, SB04]. Computing [GMA12, BGL05, Pap06, Sat06]. concept [GBR17, HBC^{+}19, RZ20, SAB^{+}05, TS20]. concept-based [HBC^{+}19]. concepts [Pap06, PMKD19, PAM09]. conceptual [jCZhC16, Cho14, SM12]. Conference [AP14, FRS10, FKRN19, FW19, KR15, MMJ20, WC12]. conflict [KRB21]. confronts [BWE07]. Connecting [WSMK07]. connectivity [JKT19]. considerations [GBR17]. consistent [ACZ06]. Constitutional [CARM20]. constraint [ACZ06]. construct [AFMF21]. constructing [HC07, SKT98]. construction [BHS^{+}22, JDAD22]. constructs [Wes10]. contained [DBG^{+}16]. contemporary [ABDF07, GT17]. Content [dSBG21, MMK^{+}19, WWFW98, AWNV14, ALT07, BSZ04, DBBK00, GCC24, HPSB10, MNZD22a, MNZD22b, MG06, PB20, RWS06, RTM05, SS98, VD08, WB98, dBOS16]. Content-based [MMK^{+}19, WWFW98, MG06, RTM05]. content-oriented [ALT07]. contents [BcCC06, CSMW22]. context [BTN12, BA18, CDN^{+}98, PRS97]. context-aware [BTN12]. context-based [PRS97]. contextualization [CG18]. continual [Kor16]. contributive [LS12]. contributors [Fre14]. control [BJWW04, Kor16, KM13]. Controlled [BKG^{+}97, jCZhC16]. Convention [CARM20]. conversion [NMT10]. Cooper [JNV18]. copies [CN15]. coping [SAPN00]. Core [dSRL19, dSRL19]. corpus [AFMF21, SKS^{+}99]. Correction [CBC^{+}23b, MNZD22a, PCDM^{+}22a, BTN12].

Correspondence [Aya22]. cost [BMG10, HBC^{+}19]. costing [PETS12]. councilors [SSK^{+}24]. course [KDP14]. coverage [AWNV14]. COVID [GCC24]. COVID-19 [GCC24]. crawler [FLF18]. create [KJ05]. created [BKS^{+}15]. Creating [WB98, YG17, Bor04, Sch04, SLM04]. creation [AFS00, GMPT00, KBG23]. creative [IMR12]. criteria [dSBG21, BKSR10]. CRM [FM17, HDE17, LHD08, MTLH17, Nic17a, Nic17b, NH17, RNFD16, SFTD14, VM17]. CRMarchaeo [Ron17]. CRMba [Ron17]. CRMdig [CVHS20]. CRMgeo [HDE17]. Cross [SFT22, CFP21, DFT^{+}12, VF21, WLC04]. cross-disciplinary [VF21]. cross-domain [CFP21]. cross-language [WLC04]. Cross-lingual [SFT22]. cross-modal [DFT^{+}12]. crossroads [Ioa05]. crowdsourced [MS14]. Crowdsourcing [ABB^{+}19, AFMF21]. CS [FW19, MMJ20, KSH^{+}99]. CSO [SOM22]. CULTURA [SAS^{+}14]. Cultural [CFP21, FRH^{+}20, HN21, AOP18, Bar21, CVHS20, cCWWK05, CCvdM08, GCC24, LHD08, LRT21, MMK^{+}17, OHBR15, STP07, Wu16]. culture [vKdBV19]. culturomics [TBC^{+}15]. cuneiform [Buc23]. Curating [KAMKJ19]. curation [Can12]. Current [ADH^{+}22, TK20, ABDT^{+}17]. curriculum [OYP^{+}16]. customized [ACZ06]. CWSpace [RWS06].

D [DVP09]. DAITSS [Cap07]. DANS [PETS12]. Darmstadt [Sch04]. DART [TMTA07]. Data [FRT21, HN21, KPC^{+}21, AQM^{+}97, ABB^{+}19, AGC05, ABDT^{+}17, Bar07, BS97, BDAB18, BBF^{+}10, BDF^{+}15, BT16, BWE07, BDS^{+}15, Buc23, CCMC17, CFP21, CRK18, DG19, DVP09, ELH^{+}17, HWT17, HGG^{+}21, ISY20, JGL^{+}97, KW19, KVZ17, KRB21, MTLH17, MR21, MSY10, MBS21, PETS12, RBdP^{+}15, RZ20, RGÓ^{+}15, SBF^{+}17, VC17, VF21, WTH17, Zim07, ABB^{+}19, PWB05, SMK^{+}14]. Data-Seeking [KPC^{+}21]. database [MPLR02, TLJ^{+}02, vKdBV19]. databases [ACC^{+}97, Gre97]. Dataset [TMTA07, BHD22]. datasets [FJ20]. Daubechies [WWFW98]. DDC [LKA^{+}17]. dealing [HPSB10]. decentralised [LS12]. decision [FBR07]. decisions [BGAE24]. Declaration [BDvdS06]. decorated [VM17]. deductive [LHD08]. deep [KGE23, PKK18]. DeepMetaGen [KGE23]. defined [ARA23]. Delaunay [MRY06]. delivery [GMP05]. DELOS [GPT05]. DelosDLMS [IMSS08]. deluge [BWE07]. demand [YRW^{+}05]. Dendro [dSRL19]. Dependency [MT09, TKM15]. depth [KWK^{+}22]. derive [ABDF07]. description [HC06a, VCC00]. descriptions [GABD07, LRT21, Wes10]. descriptor [dSRL19]. Design [BCMC23, SS06, ABDF07, BDF^{+}15, BGL05, GHJ^{+}09, KG21, RAA^{+}12, SRGK16, WLS06, Wau06, WRSB23, NA13]. Designing [Clu19, CSSS21, Dil99, dOGM99]. designs [HN07]. Detecting [AWN16b, TLJ^{+}02]. detection [AFMF21, BNLD22, Fis21, MMT04]. deteriorating [Lan15]. determine [IMR12]. determining [AS02]. develop [ABDT^{+}17]. Developing [LTCC13]. Development [PRS97]. developments [SB23]. devices [GN16]. DiCoMo [BMG10]. did [BKSR10]. diets [Clu19]. differences [KMM19]. different [BPL22, NA13]. diffusion [BM20]. Digital [AP14, ADH^{+}22, ACM^{+}98, BCGP97, Cap07, cCWWK05, Cra23, Di 23, DBG^{+}20, FFF^{+}00, FRS10, FMNH04, FKRN19, FW19, Gri23a, Ioa05, IMA^{+}05, LCN^{+}04, MMJ20, NF04, NS00, PEC^{+}15, RPM^{+}15, Tós04, ZMM^{+}12, AB04, ACDG10, AOP18, BTN12, BNC12, Bar21, BLG^{+}07, BPL22, BS97, BKG^{+}09, BDvdS06, BSC^{+}99, BCLS00, BBF^{+}10, BDF^{+}15, BCMC23, BBJ04, Bor04, BKS^{+}05, BWE07, BDS^{+}15, CCCJ10, Can12, CAA^{+}07, CLvHM18, CGR08, CPT04, CDN^{+}98, Cho14, CESF99, CBC^{+}23a, CBC^{+}23b, Cru99, DFT^{+}12, DWD12, DOTW97, Dil99, DVP09, EMS^{+}98, FS10, FD98, FBR07, FRT21, FCR20, FTA^{+}07, GABD07, GMPT00, GFW08, GT17, GWB^{+}08, GPT05, Gri15, HSPQ04, HR97, HB16, HBC^{+}19, HCH04, HCHH05, HC06b, JDAD22, JGL^{+}97, JCMB00, KW06, KK99, KLL08, KTS00, KFRF^{+}04, KW19, KSH^{+}99, KGJK14, KBG23, KFW04, KDP14]. digital [Kor16, KD21, KM13, KI04, KSAP09, LS12, Lan15, Lee07, LLSG08, LHD08, MT00, Mar04, MHM08, MT09, MHML05, MSS12, MM13, MAC^{+}15, MFC^{+}18, McD06, MSHC04, MIW^{+}08, MSY10, MBS21, NV06, NMT10, NA13, NPS05, OD09, OYP^{+}16, OSD23, OSCF09, PD19, PETS12, PTB^{+}19, Par04, PWM^{+}05, PB20, PAM09, RBE04, RSF08, RH05, Sae00, SASV12, SP04, SSNM97, Sch04, SSB^{+}12, SMD00, SBC04, SC05, Smi97, SJ09, SH05, SAS^{+}14, SB04, SAPN00, SM12, SKT98, dSTMGF06, TK20, TCK^{+}97, TKM15, VCH08, Wan99, WLC04, WLS06, WGZ^{+}06, Wat23, WB98, WR18, WDDT04, WHK^{+}97, WSMK07, WC12, ZJ98, ZYZ^{+}00, dOGM99, dlCAS^{+}08, BPL22, CCFI10, KR15, RAA^{+}12]. digitised [BNLD22]. digitization [BMG10]. digitized [GMP05]. DILIGENT [CAA^{+}07]. dimensional [KK99, NH16, ZM19]. Directory [GTP^{+}99]. disambiguation [PP20, SGLF15]. disciplinary [VF21]. discontent [Par04]. discourse [CG18, SMD00, dRF18]. Discovering [PTB^{+}19, KN14]. discovery [KFRF^{+}04, KPK^{+}24, PWB05, RSF08, WL97]. discrimination [TNZR13]. discussion [jCZhC16]. disintermediation [IMR12]. display [BCGM00]. dissemination [PETS12]. dissertations [RSF08]. distance [SDLÅ14]. distance-based [SDLÅ14]. Distant [Gri23a, Gri23b, GGM23, WRSB23]. distributed [JGL^{+}97, KW06, Sch04]. diverse [Zim07]. diversity [BDS^{+}15, SAPN00]. divides [MHM08]. DLs [ORR12]. do [KM13]. document [BM20, CG18, CD18, DFT^{+}12, FRD23, JCRK18, KYK23, Ke15, KGJK14, KF98, KNH97, KPK^{+}24, MMT04, MSW^{+}05, OSD23, PRS97, SKS^{+}99, TPMG15, XI24]. documentation [CVHS20, NMT10, RNFD16]. Documenting [Nic17a, vKdBV19, VM17]. documents [ACC^{+}97, AS02, BNLD22, HPSB10, SOM22, SMD00, ZBS^{+}10, dRF18]. domain [AWNV14, BDAB18, CFP21, jCZhC16, PTB^{+}19, SGLF15, SDLÅ14]. domain-specific [SGLF15, SDLÅ14]. Drag [BSC^{+}99]. Drag-and-drop [BSC^{+}99]. Drawing [BGL05]. drift [AN15, TS20]. driven [BTN12]. drop [BSC^{+}99]. Drupal [VM17]. DSpace [RWS06]. Dublin [dSRL19]. duplicate [MMT04, TLJ^{+}02]. during [GCC24]. Dutch [vKdBV19].

e-commerce

[DLR02, Fad02, LKR02, SPLF02]. e-learning [FMNH04]. e-payment [BDY02]. e-research [TMTA07]. earth [PWB05, CAA^{+}07]. East [HLZ05]. ECDL [FRS10]. ECHO [SP04]. ecological [BLG^{+}07]. ecology [BWE07]. ED [KM13]. editions [DIJ^{+}06]. editor [BM05, FKX18, KR15, MTT15, MSB^{+}17, NV06, PD19]. Editorial [Ano00, HN21, HDH05, KMW16, RCD07]. editors [NF04, AR04, FRS10]. Education [BGL05, MIW^{+}08, NA13, WPS^{+}04]. Educational [CGR08, KDP14, Kor16, KD21, MHM08, MSS12, MIW^{+}08, RWS06, RD10, RZ20, SAB^{+}05, ZJ98]. Effective [CC04]. Effectiveness [BCGM00, BACB20, BKS^{+}06, STM^{+}20].

Efficient [YSZ06, CSSS21, WRSB23]. electronic [ADT00, CLM06, DIJ^{+}06, MPLR02, RSF08, Sto00]. element [WL97]. elements [dRF18]. embedded [BWE07]. embedding [EESB20]. embeddings [EWSB20, MBVL18]. Emerging [CBRCW05, Wat23]. empirical [Kor16]. empirically [SLS19]. empowering [VC17]. enabled [KSAP09]. encapsulated [Wau06]. encoder [SZOS22]. Encoding [DG19, McD06, RM19]. Encounters [CNHB16]. encyclopedia [KSG10]. end [GTP^{+}99]. Energy [STM^{+}20]. engagement [AOP18]. engaging [CBC^{+}23a, CBC^{+}23b]. English [OTA00, SFT22, TNZR13]. English-text [OTA00]. Enhanced [MSHC04, MFC^{+}18]. enhancement [ABB^{+}19]. enrich [BB17, CRK18, KDP14]. enrichment [SBF^{+}17]. entities [FISN20, Hus19]. entity [FISN20, HTR15, KPK^{+}24, PCDM^{+}22a, PCDM^{+}22b]. entity-centric [FISN20]. entity-interaction-aware [KPK^{+}24]. environment [ACM^{+}98, CESF99, PP20, SAS^{+}14, ST04, ZMM^{+}12, dBOS16]. environmental [HWT17, WTH17]. environments [JBFNMD15, PWM^{+}05, Sat06]. Envisioning [Buc23]. epidemic [Fis21]. Epigraphy [Gri23a, Gri23b, SC23]. epistemic [FS10]. EPrints [NMT10]. equal [BKS^{+}15]. Erratum [WBL^{+}07a]. Error [KF98]. errors [dOVA^{+}23]. eScience [Bar07, CMM^{+}07, CBB07, WSMK07]. establishing [CFP21]. ethnic [Can12]. EthnoMuse [SM12]. ethnomusicological [SM12]. EU [GMP^{+}23]. EU-DELOS [GPT05]. European [FRS10, TBF^{+}04]. evaluate [BM16, Fig21]. Evaluating [AN15, JDAD22, SAS^{+}14, SDLÅ14, dBOS16, dOVA^{+}23, ADT^{+}19, BKG^{+}09, Dil99, LRT21]. Evaluation [FTA^{+}07, Wes10, BLG^{+}07, BCMC23, BGL05, GHJ^{+}09, GWB^{+}08, JP05, KG21, KN14, KM13, OD09, PCSH10, PTB^{+}19, PRS97, SK15]. event [BNLD22, Fis21, GRD20, NPD20]. event-based [Fis21]. event-centric [GRD20]. events [FLF18, SI21]. everyday [CNHB16]. evidence [KRZ18]. evidence-based [KRZ18]. evolution [CGS17, FISN20, HTR15, Mar04]. evolving [Lee07]. Exact [CIMR12]. Examining [ALT07, Hus19, MHM08]. execution [LAP06]. exhibitions [HCH04]. expanded [OTA00]. expansion [CVDT22]. Expectation [DSGF06]. Expectation-based [DSGF06]. experience [Wau06, WPS^{+}04]. experienced [NF04]. Experiences [ELH^{+}17, SLM04, ABB^{+}19, HCHH05]. experiment [DWD12]. experimental [BKS^{+}06, SBF^{+}17]. expert [CVDT22]. expertise [dSCGL20]. explicit [DMRA97]. Exploiting [Kel24]. exploration [KH17, MSW^{+}05, SKT98, VH21]. exploratory [BKSR10, BDF^{+}15, CLvHM18, TMBC22]. Explorer [HC07]. Exploring [DLR02, MTT15, BCMC23, CCCJ10, FMN18]. express [MR21]. Expressing [NH17]. Expressiveness [PMKD19]. extended [FHKN20, KB22]. Extending [FCR20, Nic17b, SSB^{+}12, FMN18]. extensible [GCAN00]. extension [HDE17, RNFD16]. Extracting [JTCA97, dRF18, GRD20, HN07]. Extraction [BSZ04, BP19, Fre14, NAO08, RT18].

face [AB04]. facet [Bar21, CCCJ10]. faceted [JCRK18]. facets [IMR12]. facilitate [GABD07, KAMKJ19]. facilitating [OHBR15]. facsimiles [SASV12]. factually [ACZ06]. FAIR [Bar21, HN21, HGG^{+}21, VF21]. fast [AS02]. Feature [SI21, ADL18]. features [ABDF07, BM20, CSSS21]. Federated [Has97]. federation [VCH08]. federations [KNH97]. Fedora [LPSW06]. feedback [HPSB10]. field [MKFK18, OYP^{+}16, PEC^{+}15]. fieldwork [BLG^{+}07]. file [GN16]. film [SP04]. films [TBF^{+}04]. finding [HKS^{+}15]. finetuning [XI24]. first [DWD12]. Físchlár [SLM04]. fisheye [JP05]. Flickr [MHTH15]. Florida [Cap07, MBS21]. Fluid [SH05]. flute [KAMKJ19]. Focused [Di 23, FLF18, AP14, FKRN19, FW19, JKT19, MMJ20, ORR12, SF15]. focusing [Nic17b]. focussed [FRS10, HK12, RPM^{+}15]. Font [BB17]. foreign [CBC^{+}23a, CBC^{+}23b]. Foreword [Ano97]. forgotten [HKS^{+}15]. form [LCK^{+}18]. formal [GFW08, RZ20]. format [Cap07]. formative [Kor16]. forms [Zim07]. formulation [VH21]. Foundation [RBdP^{+}15]. foundations [CCFI10, dlCAS^{+}08]. four [NH16, SLM04]. four-dimensional [NH16]. framework [BHD22, BSC^{+}99, Cho14, CARM20, DFT^{+}12, Dil99, GMP05, HJK^{+}09, HCH04, KW06, KTS00, KG21, KM13, MHM08, MMK^{+}17, SZOS22, dSTMGF06, dOGM99]. France [MR21]. free [JTCA97]. frequencies [EWSB20]. full [ZG04]. full-text [ZG04]. functional [Kel24]. functionality [CLM06, ZBS^{+}10]. functions [KK99]. Fusion [TS20]. future [Cra23, WSMK07]. fuzzy [Fig21, KG21, OTA00]. fuzzy-based [KG21].

Gallery [BGL05]. game [LTCC13]. gap [JBFNMD15]. Gazetteer [GGM23]. gazetteers [NH16]. gender [KMM19, dRF18]. generalized [TPMG15]. generate [BGAE24, KGE23]. generated [NMWP99]. Generating [ACZ06, LHD08]. generation [CPT04, GPT05]. generic [MSW^{+}05, VD08]. geographic [CSL^{+}97, dOGM99, PCSH10]. geometric [BM20]. geospatial [CSL^{+}97]. German [KSG10, MMK^{+}19]. Germany [Sch04]. Getulio [RBdP^{+}15]. GigaDB [ELH^{+}17]. GIS [SS06]. Global [GTP^{+}99]. goals [SAB^{+}05]. gone [KN14]. gram [TLJ^{+}02]. graph [FRD23, GN16, JDAD22, KDA24, SM21, BHS^{+}22]. graph-based [FRD23, SM21]. graphical [HC07, JMS99]. graphics [LW09]. graphs [KPK^{+}24, MTT15]. Greece [MNZD22a, MNZD22b]. Greenstone [MSHC04]. Grid [CAA^{+}07, PWB05]. grids [AGC05]. grounded [Aya22]. group [Sch04]. groups [GPT05]. Guest [BM05, FKX18, KR15, MSB^{+}17, NV06, PD19, FRS10].

habitat [BWE07]. Hand [BGL05]. handwritten [BP19]. Harmonizing [Ron17]. Harvesting [KDP14]. HathiTrust [DBG^{+}20]. haz.hu [Tós04]. health [TCK^{+}97]. healthcare [KM13]. Held [BGL05]. help [MBVL18]. Heritage [CFP21, AOP18, Bar21, CVHS20, CCvdM08, FRH^{+}20, GCC24, LHD08, LRT21, MMK^{+}17, OHBR15, Wu16, HN21]. heterogeneous [DFT^{+}12, OHBR15, Wu16, YSZ06]. Heuristic [BP19]. heuristics [SGLF15]. Hidden [CSSS21]. hierarchical [AIL16, BHD22, CSMW22, RM19, dSCGL20, BSZ04]. hierarchies [NMWP99]. High [STM^{+}20, JGL^{+}97, MS14, PB20, SPLF02]. high-quality [PB20]. high-yield [SPLF02]. higher [MIW^{+}08, NA13]. historic [TBF^{+}04]. Historical [BM20, EKV^{+}10, GGM23, ABDF07, BDAB18, cCWWK05, DBG^{+}16, JKT19, MMK^{+}19, PCDM^{+}22a, PCDM^{+}22b, RBdP^{+}15, SP04, WGZ^{+}06]. History [SMK^{+}14, FISN20, GN16, MKFK18, SJ21]. history-related [SJ21]. hit [GTP^{+}99]. Holistic [FRD23]. Homer [SB23]. hospitals [Has97]. HuCI [CPR23]. Human [WIMM24, Aya22, dSBG21, RD10]. human-centered [Aya22]. Human-in-the-loop [WIMM24]. humanities [CRK18, CPR23, CBB07, MBS21, SAS^{+}14, Wat23, BPL22, CPR23]. Humdrum [DG19]. hybrid [KSG10, Wan99]. hypermedia [HCH04, ZBS^{+}10]. hypertensive [Clu19]. hypertext [BNC12]. hypertext-based [BNC12]. HyperView [FS00].

ICADL [ISK^{+}24, JMI^{+}22, JKPC24]. ideas [GWB^{+}08]. Identification [YKP20, CSSS21, HDG19]. identifier [KTT^{+}14]. identifiers [KB22]. Identifying [CS98, MBVL18]. idf [Hie00]. IEEE [KSH^{+}99]. IEEE-CS [FW19, MMJ20]. IFLA [FCR20]. II [Gri23a]. IIIF [FRH^{+}20]. IJDL [HK12, NV06]. Image [MKFK18, CC04, CCGS14, DIJ^{+}06, GCAN00, JP05, Par04, SH00, WWFW98, GCAN00]. image-based [DIJ^{+}06]. imagery [cCWWK05, WGZ^{+}06]. images [BP19, BKG^{+}97, Par04, SS98]. imaging [NPS05, WHK^{+}97]. IMAGO [BPL22]. impact [BNLD22, BKS^{+}15, BKWN16, KM13, PTB^{+}19, SFT22, dOVA^{+}23, KM13]. Implementation [HWT17, RAA^{+}12, SS06, WLS06, ZMM^{+}12]. implementing [dOGM99]. important [IMR12]. improve [WR18]. Improved [SB14, CVDT22]. Improving [BT16, EWSB20, KRB21]. in-depth [KWK^{+}22]. in-text [BA18]. incremental [KJ05]. Index [CPR23, KG21]. indexing [BKG^{+}97, DBBK00, Hus19, TS20, WWFW98, ZG04]. indigenous [MS14]. individuals [Clu19]. infection [KM13]. inference [LHD08]. InfoMonitor [CGM05]. Information [Ke15, PCSH10, ACZ06, ACM^{+}98, ASFF19, Aya22, BCLS00, BM05, BB17, CCCJ10, CMM04, DBG^{+}16, DMRA97, GHJ^{+}09, Has97, Hie00, IMA^{+}05, JTCA97, KFRF^{+}04, LA00, LLS16, MMK^{+}17, MFC^{+}18, MKPI^{+}22, MSB^{+}17, NAO08, RT18, RTM05, SKS^{+}99, SBF^{+}17, Smi97, SFTD14, SKT98, dSTMGF06, VSVS08, Wau06, Wu16, dOVA^{+}23, dSCGL20]. information-seeking [GHJ^{+}09]. information-technology [IMA^{+}05]. Information-theoretic [Ke15]. infrastructure [CAA^{+}07, CPT04, Les15, NMSK07, RBE04, Wat23, WTH17, YRW^{+}05]. infrastructures [BDS^{+}15, DBT^{+}15, IMA^{+}05, CMT13]. Inheriting [GT17]. initial [GBR17]. InscriptiFact [HLZ05]. inscriptions [HLZ05]. Insights [JCRK18, WRSB23]. inspection [HSPQ04]. Institute [OOKS13]. institutional [GCC24]. instruction [dlCAS^{+}08]. INSYDER [RTM05]. integrated [ELH^{+}17, HCHH05, HC06b, KNH97, TCK^{+}97]. Integrating [AFS00, CAA^{+}07]. integration [Has97, MMK^{+}17, Wu16]. integrity [SJ09]. intelligence [FD98, Fis21]. intelligent [FBR07, HCH04]. intensive [EMS^{+}98]. Interaction [FS10, DFT^{+}12, KPK^{+}24, SAPN00]. interactions [dSRL19]. Interactive [BTN12, LTCC13, SKS^{+}99, BCGM00, LA00, SAB^{+}05]. interdisciplinary [HDG19]. interests [CS98]. interface [BSC^{+}99, DLR02, HC07, JMS99, Wit07, dOGM99]. interfaces [BM05, Cru99]. intermediate [RT18]. International [AP14, FKRN19, LCN^{+}04, NS00]. Internet [AAWN14, KNH97]. interoperability [BT16, CFP21, SMK^{+}14, TKM15, WBL^{+}07a, WBL^{+}07b]. interpretation [ADT^{+}19]. interpreting [DBG^{+}16]. Introduction [AP14, Ano98, AGS05, FKRN19, FW19, Gri15, LCN^{+}04, MMJ20, MFC^{+}18, ORR12, RPM^{+}15, SF15, BM05, CGR08, FKX18, FRS10, GST19, KLL08, KR15, MTC^{+}16, MSB^{+}17, NV06, PD19]. Investigating [CLvHM18]. investigation [GCC24]. investments [Fad02]. involving [XI24]. IR [MBVL18]. Irish [DBG^{+}16, KAMKJ19, RGÓ^{+}15]. Island [YG17]. ISO [AIL16]. Issue [Di 23, Gri23b, HN21, JMI^{+}22, NS00, AP14, AR04, BM05, CGR08, Cru99, FKX18, FRS10, FKRN19, FW19, GST19, Gri15, HK12, ISK^{+}24, JKPC24, KLL08, KMW16, KR15, LCN^{+}04, MMJ20, MTC^{+}16, MFC^{+}18, MSB^{+}17, NV06, ORR12, PD19, Pap06, RPM^{+}15, SF15]. Issues [LHD08, WHK^{+}97, jCZhC16, SSK^{+}24]. Item [BDvdS06]. items [Kor16].

Japanese [KTT^{+}14, NAO08]. JavaScript [BKWN16]. JCDL [FW19, MMJ20, SF15, WC12]. JDL [Pap06]. JoDL [BM05]. John [Tós04]. Joint [FW19, MMJ20, WC12]. Journal [LCN^{+}04, NS00, KBFG19, NF04, STM^{+}20]. just [ZBS^{+}10]. just-in-time [ZBS^{+}10]. justification [Hie00].

keep [GN16]. Key [ABDT^{+}17, ABDF07]. Keyframe [MRY06]. Keyframe-based [MRY06]. keyword [BCGM00, ISY20]. keyword-based [BCGM00]. KMIR [MSB^{+}17]. Knowledge [BDS^{+}15, BHS^{+}22, KMM19, RSF08, BB17, CSL^{+}97, EMS^{+}98, FMNH04, JDAD22, KDA24, Les15, LHD08, MSB^{+}17, NHP^{+}21, SRGK16, TBC^{+}15, TMBC22, YG17, Zim07, DBT^{+}15, GST19, MTC^{+}16, PMKD19, ZM19]. knowledge-based [TBC^{+}15]. KOS [PMKD19, ZM19].

label [CSMW22, XI24]. labeled [SI21]. labeling [OSD23, dBOS16]. labels [SM21]. Laboratoire [MR21]. Language [MFP18, AQM^{+}97, AWNV14, CBRCW05, KWK^{+}22, ŁB15, MFC^{+}18, MS14, WLC04]. languages [CBC^{+}23a, CBC^{+}23b]. large [FISN20, JGL^{+}97, KLL08, KRB21, NMWP99, NHP^{+}21, SFT22]. large-scale [KRB21, NHP^{+}21, SFT22]. latent [KDP14, WIMM24]. Latin [ARA23]. layers [FHKN20]. layout [BM20]. leaderboards [KDA24, KDA24]. learned [GHJ^{+}09]. learning [EMS^{+}98, FMNH04, GABD07, KYK23, PKK18, SAB^{+}05, VSVS08, VD08, WIMM24]. learning-based [PKK18]. LectureLounge [WPS^{+}04]. legislative [CSMW22]. LEMO [HJK^{+}09]. Lessons [GHJ^{+}09]. level [AWNV14, CLvHM18]. leveled [Dil99]. leveraging [KGE23]. Lexically [NMWP99]. Lexically-generated [NMWP99]. liberal [PEC^{+}15]. libOR [VSVS08]. Libraries [AP14, ADH^{+}22, FRS10, FKRN19, FW19, Gri23a, KR15, MMJ20, AB04, BTN12, Bar21, BPL22, BS97, BSC^{+}99, BBJ04, BKS^{+}05, BWE07, BDS^{+}15, CLvHM18, CGR08, CPT04, CDN^{+}98, Cho14, CMM^{+}07, CESF99, Cra23, Cru99, DWD12, Dil99, FS10, FD98, FRT21, FMNH04, FTA^{+}07, GABD07, GPT05, HSPQ04, HBC^{+}19, Ioa05, JGL^{+}97, KK99, KLL08, KTS00, KW19, KSH^{+}99, KFW04, KM13, KI04, LS12, LLSG08, LCN^{+}04, LHD08, MFC^{+}18, MSHC04, MIW^{+}08, MSY10, MKFK18, NF04, NA13, OOKS13, OSD23, PD19, Par04, RBE04, RSF08, RH05, Sae00, SSNM97, Sch04, SC05, Smi97, SB04, SAPN00, SKT98, TK20, Wan99, WLC04, WSMK07, WC12, ZJ98, dlCAS^{+}08, LCN^{+}04, NS00]. Library [Di 23, FCR20, Tós04, ABB^{+}19, ACDG10, BNC12, BLG^{+}07, BCLS00, BBF^{+}10, BDF^{+}15, BCMC23, Bor04, CAA^{+}07, CRK18, CBC^{+}23a, CBC^{+}23b, DFT^{+}12, DOTW97, EMS^{+}98, GMPT00, GFW08, GT17, GWB^{+}08, HR97, HB16, IMA^{+}05, JDAD22, JCMB00, KFRF^{+}04, KBG23, KDP14, Kor16, KD21, Mar04, MHM08, MHML05, MSS12, MM13, McD06, MBS21, OYP^{+}16, PTB^{+}19, PAM09, SP04, SMD00, SBC04, SLM04, SM12, dSTMGF06, TCK^{+}97, WLS06, WLC^{+}19, WB98, WR18, WDDT04, WHK^{+}97, ZYZ^{+}00, dOGM99, BCGP97, CCFI10, Cra23, DBG^{+}20, FFF^{+}00, RAA^{+}12]. lies [DBT^{+}15]. life [CNHB16]. lifecycle [BWSH12]. Lightweight [SPO08, GMA12]. like [FMN18]. lingual [AFMF21, SFT22]. Linguistic [SSNM97]. linguistically [SLS19]. linguistics [LMZ^{+}18]. linkage [LMZ^{+}18]. Linked [RGÓ^{+}15, BT16, MTLH17, RBdP^{+}15, ABB^{+}19, SMK^{+}14]. Linking [BPL22, SAB^{+}05, ITSY06, MSS12, PCDM^{+}22a, PCDM^{+}22b]. links [AAWN14]. listening [ABB^{+}19]. lists [KHVT23, dSRL19]. literature [BGLB16, HDG19, KS22, LLSG08, LMZ^{+}18, PRS97, WIMM24]. Little [BWE07]. Liverpool [WHCF20]. load [KTS00, ZYZ^{+}00]. locality [LLS16]. locality-aware [LLS16]. locate [Zim07]. Location [HB16, Sat06, ITSY06]. Location-based [Sat06]. Location-triggered [HB16]. LOD [WR18]. LODmilla [MTT15]. log [JCMB00, KJ05]. logics [VCC00]. Long [DVP09, BWSH12, OSD23, PETS12, SJ09]. Long-term [DVP09, BWSH12, PETS12, SJ09]. loop [WIMM24]. loosely [KNH97]. Lorel [AQM^{+}97]. Lost [HKS^{+}15]. low [HBC^{+}19, ZG04]. low-cost [HBC^{+}19]. LZW [AS02].

M [RZ20]. Machine [WGZ^{+}06, EMS^{+}98, PMKD19, RD10]. machine-based [RD10]. maintain [Sae00]. maintenance [Lee07]. make [CN15]. management [AGC05, BcCC06, BWSH12, FRD23, FMNH04, JGL^{+}97, KDP14, KTT^{+}14, MT09, SKS^{+}99, SMK^{+}14, VF21, WIMM24]. Managing [MPLR02, Clu19, STP07]. manent [FM17]. map [SAB^{+}05]. Mapping [LKA^{+}17, MNZD22a, MNZD22b, PI04, MMK^{+}17, Nic17b, OHBR15]. mappings [ARA23, GMA12]. Maps [DMRA97, BB17, MSB^{+}17, YG17]. marketplace [ADT00]. markup [DIJ^{+}06]. Master [GTP^{+}99, GMP^{+}23]. matching [CIMR12]. material [RZ20]. materials [cCWWK05, SM12, WGZ^{+}06]. matrix [HMG20]. matrix-based [HMG20]. meaningful [EESB20]. measure [Aya22, Gre97]. measuring [AN15, BKS^{+}15]. Media [LTCC13, AWN16a, ASFF19, FISN20, HDG19, LCK^{+}18, NMT10]. medical [PRS97]. medicine [KRZ18]. meet [MPLR02]. Meeting [WTH17]. MEI [DG19, RM19]. MEI-based [RM19]. MELHISSA [PCDM^{+}22a, PCDM^{+}22b]. members [KMM19]. Memento [Kel24]. mementos [BKS^{+}15]. Memory [YG17]. mention [YKP20]. MeSH [NKTP21]. message [AR04]. Meta [Smi97, KGE23]. Meta-information [Smi97]. meta-reviews [KGE23]. Metadata [Gre97, PWM^{+}05, SMK^{+}14, AFS00, BTN12, BCGP97, FRH^{+}20, GCC24, Hus19, KDP14, KWK^{+}22, Lee07, LTCC13, LKA^{+}17, MTT15, NMT10, OD09, OHBR15, RWS06, WLC^{+}19, WB98, Zim07]. Metadocuments [KFRF^{+}04]. metaphors [GT17]. method [AS02, BM16, CRK18, FAM00, KHVT23, RZ20, SOM22, SGLF15]. Methodologies [OSCF09]. methodology [EKV^{+}10]. methods [BKS^{+}06, ISY20, KYK23, KN14, MSS12]. METS [McD06]. migration [Cap07]. minimal [GMA12]. mining [HGG^{+}21, KDA24, KJ05, RZ20, WLC04, WDDT04]. missing [BKS^{+}15, SM21]. mitigating [dOVA^{+}23]. mixed [KHVT23]. mixed-method [KHVT23]. mixture [CD18]. mobile [BGL05, HB16]. modal [DFT^{+}12]. model [ABDT^{+}17, BMG10, Cho14, KFW04, MSY10, OHBR15, PETS12, SASV12, VD08, ZG04, dRF18, FCR20, GWB^{+}08]. modeling [DMRA97, LW09, MBVL18, TPMG15]. modelling [CARM20, GBR17, PC17]. models [BKS^{+}05, CD18, EESB20, Ron17]. modern [KYK23, TNZR13]. modularization [HCH04]. moments [BNC12]. monitoring [KTS00, LAP06]. mono [AFMF21]. mono-lingual [AFMF21]. MOPPET [ADT00]. morphology [SB23]. motivated [SLS19]. Moved [KN14]. moving [FD98]. MPEG [BDvdS06]. MPEG-21 [BDvdS06]. Multi [CSMW22, dSBG21, Dil99, FAM00, SI21, XI24, ZM19]. multi-dimensional [ZM19]. Multi-label [CSMW22, XI24]. multi-labeled [SI21]. multi-leveled [Dil99]. multi-video [dSBG21]. multi-word [FAM00]. Multidimensional [BS97]. multidisciplinary [VF21]. Multilabel [SM21]. multilevel [CC04]. multilingual [BNLD22, PCDM^{+}22a, PCDM^{+}22b]. Multimedia [BcCC06, ACM^{+}98, BSC^{+}99, CSL^{+}97, HJK^{+}09, KFW04, ZJ98]. multiple [BA18]. multispectral [NPS05]. Multitext [SB23]. Museum [LTCC13, HCH04, HCHH05, NPS05, PWM^{+}05, WHCF20, Wu16]. Museums [WHCF20, SH05, Wu16, SMK^{+}14]. music [Can12, DFT^{+}12, DG19, FBPC17, FCR20, IMR12, ORR12, RM19, WLS06, WLC^{+}19]. musical [BNC12, BP19]. musicological [vKdBV19]. musicology [DBG^{+}20, KW19, PD19]. myself [CN15].

Name [KTT^{+}14, SGLF15]. Named [HTR15, Hus19]. names [RGÓ^{+}15]. Nan [HCHH05]. narrative [KPK^{+}24]. narratives [MBM19]. NASA [GTP^{+}99]. national [BDAB18, Bor04, KM13, WHCF20, Wu16]. Natural [MFP18, ŁB15, MFC^{+}18, SMK^{+}14]. Nature [MT00]. navigation [Sto00]. NC-value [FAM00]. NCSTRL [KSH^{+}99]. NDLTD [KSH^{+}99, RSF08]. Near [HLZ05]. nearest [OSD23, SGLF15]. negatives [Lan15]. negotiation [SBC04]. neighbor [OSD23]. Netherlands [OOKS13]. network [KMM19]. networked [Buc23, SRGK16, KMM19]. networks [BWE07, CSMW22, KNH97, RD10, SSB^{+}12]. Neuma [RAA^{+}12]. Neumann [Tós04]. neural [KGE23, PKK18]. new-generation [GPT05]. News [ACM^{+}98, dSBG21, CNHB16, HDG19, HMG20]. newspaper [GMPT00]. NKOS [KMM19]. NLP [EESB20]. noise [BNLD22]. nominated [SF15]. non [SH00]. non-topical [SH00]. Note [HN21]. notions [EESB20]. novel [SGM^{+}22]. NSF [GPT05]. NSF/EU [GPT05]. NSF/EU-DELOS [GPT05].

OAI [SSB^{+}12, Wit07]. OAISRB [Wit07]. object [ACC^{+}97, ELH^{+}17, KW06, Wau06]. objects [HC06b, JGL^{+}97, LPSW06, McD06, NV06, PI04, VD08]. observation [CAA^{+}07]. occluded [SS98]. occurrence [Fre14]. OCR [BNLD22, dOVA^{+}23]. Off [VM17, AWN16b]. Off-the-shelf [VM17]. off-topic [AWN16b]. on-demand [YRW^{+}05]. Online [KHVT23, LGT97, BLG^{+}07, GCC24, MIW^{+}08, SLS19, WHCF20]. onomasiologically [SLS19]. ontological [GBR17, VF21, WB98]. ontologies [AFS00, GMA12, MAC^{+}15, SH05]. Ontology [PC17, Clu19, GFW08, SMD00]. ontology-based [SMD00]. Open [Di 23, RT18, ADT00, FRD23, GMP^{+}23, RBdP^{+}15, ST04, TBC^{+}15, BHS^{+}22, Mar04]. OpenDLib [CPT04]. operational [WTH17]. operations [VSVS08]. opinions [SSK^{+}24]. opportunistic [JP05]. opportunities [HWT17]. Optimizing [Fad02]. order [FD98]. Organization [CCCJ10, GST19, KMM19, MTC^{+}16, PMKD19, ZM19, SRGK16]. organizing [BNC12, RZ20]. OrgBR [RZ20]. OrgBR-M [RZ20]. oriented [ALT07, BBF^{+}10, KRZ18, LMZ^{+}18, Pap06, SRGK16, WRSB23]. original [CCCJ10]. origins [AS02]. ORKG [KDA24]. ORKG-Leaderboards [KDA24]. orthography [SB23]. overhead [ZG04]. overload [BACB20]. overview [CCGS14]. OWL [CVP^{+}15].

P2P [SSB^{+}12]. package [Wau06]. packages [HC07]. packaging [RWS06]. page [BG12, GMPT00, LW09]. pages [AWN16b, HKS^{+}15, KN14, NMSK07]. Palace [Wu16]. palaeography [SB23]. Paleography [Gri23a, Gri23b]. pandemic [GCC24]. PANIC [HC06b]. paper [BGLB16, HN07, KS22, SK15]. paper-based [HN07]. Papers [ADH^{+}22, JMI^{+}22, ISK^{+}24, JKPC24, SFT22, SF15, SK15, TK20, VM17]. Parallel [OFM^{+}00]. ParsCit [PKK18]. parser [PKK18]. parsing [SB14]. Part [Gri23a]. partially [SS98]. Past [Gri23a, Gri23b, CBRCW05, GGM23, SI21]. patent [NAO08]. Pathways [WBL^{+}07b, WBL^{+}07a]. patterns [ASFF19, BG12, jCZhC16, SB14]. payment [BDY02]. peer [BGAE24, KGE23, STM^{+}20]. peer-review [STM^{+}20]. PEERRec [BGAE24]. Performance [LKR02, AGC05, DG19, JGL^{+}97, KTS00, MPLR02, WRSB23]. period [NH16]. Perseus [Cra23]. Persian [RT18]. persistence [KB22]. persistent [HK12, KB22]. personal [BNC12, GN16]. Personalisation [SC05]. personalization [KI04]. personalized [SSNM97]. personalizing [dlCAS^{+}08]. persons [JKT19]. perspective [HWT17, KHVT23]. perspectives [GCC24]. photographic [Lan15]. physical [SB04]. Physics [STM^{+}20]. pick [BKSR10]. pipeline [Lan15]. place [RGÓ^{+}15]. places [GGM23]. plagiarism [AFMF21, MFP18]. Planning [LAP06, BKG^{+}09, LKR02, SS06]. plans [BKG^{+}09]. platform [LHD08, dSRL19]. platforms [ASFF19]. plausibility [PB20]. PMH [SSB^{+}12, Wit07]. point [KM13]. policies [AN15]. policy [BJWW04, ZMM^{+}12, BSZ04]. political [SSK^{+}24]. popular [FCR20]. popularity [JKT19]. portal [ACDG10]. posts [SJ21]. potential [BKG^{+}09, DBG^{+}20, FS10, Kel24, SK15]. Practical [Lee07]. Practice [AP14, ADH^{+}22, FKRN19, TK20]. practices [ABDT^{+}17, PC17, VC17]. pragmatic [dSCGL20]. pre [KBFG19]. pre-print [KBFG19]. predict [BGAE24]. prediction [SSK^{+}24]. preferences [ACDG10]. prehistoric [HGG^{+}21]. PREMIS [CVP^{+}15]. presence [dSCGL20]. Present [Gri23a, Gri23b, GGM23]. presentations [HCH04, ZJ98]. Preservation [RH05, Bar21, BKG^{+}09, BWSH12, Can12, Cap07, CN15, DVP09, FBPC17, FBR07, HC06b, JBFNMD15, Lee07, MT09, MAC^{+}15, PETS12, SSB^{+}12, TKM15, ZMM^{+}12]. preserve [NMSK07]. preserving [DBG^{+}16, DVP09, RBE04]. press [PCDM^{+}22a, PCDM^{+}22b]. prevalence [SFT22]. previews [GTP^{+}99]. primary [Aya22, BBF^{+}10]. principles [Bar21, VF21]. print [KBFG19]. privacy [RBE04, Sae00]. privacy-preserving [RBE04]. private [SKT98]. probabilistic [Hie00]. Process [Di 23, GBR17, NA13, MAC^{+}15]. processability [PMKD19]. Processing [MFP18, KYK23, ŁB15, MFC^{+}18, YRW^{+}05]. product [BWSH12, Sto00]. production [TMBC22, dBOS16]. Profiling [AWNV14, ANV^{+}16]. project [RWS06, TMTA07, ACM^{+}98, BGL05]. Promoting [AOP18]. propagating [SFTD14]. prosody [SB23]. prototype [MT00]. provenance [Buc23, HC07, SFTD14, HC07]. providers [MNZD22a, MNZD22b]. Providing [Wit07, WLC^{+}19]. provision [TCK^{+}97]. publication [HC07, MTT15, PP20, WTH17]. publications [ADL18, KS22, ŁB15, MFP18]. published [KBFG19]. publisher [FS00]. publishing [ABDT^{+}17, CCMC17, ELH^{+}17, HWT17, SMK^{+}14, VC17, WTH17]. PVAF [PP20].

QoS [MPLR02]. qualitative [MSS12, NHP^{+}21]. quality [Aya22, BHS^{+}22, CCGS14, DSGF06, Fad02, FRT21, KTS00, KG21, Kor16, KRB21, LSC19, LRT21, MS14, OD09, PB20]. Quantifying [STvO^{+}18]. quantitative [BM16, CVHS20]. queries [GTP^{+}99, IMR12]. query [AQM^{+}97, GTP^{+}99, JMS99, KI04, KWK^{+}22, KPK^{+}24, VH21, WLC04]. Querying [ACC^{+}97, MMM97, AFS00, FHKN20]. question [LSC19, SGM^{+}22, SZOS22]. quiescent [CCCJ10].

radiology [HR97, JTCA97]. Ranking [dSRL19, BS97]. ray [LGT97]. RDF [AFS00]. re [MHM08]. re-examining [MHM08]. reading [FM17, KHVT23, WRSB23]. Real [JGL^{+}97, JBFNMD15, KN14]. Real-time [JGL^{+}97, KN14]. realization [BCMC23]. Reasoning [BJWW04, TKM15]. recognition [FAM00, HTR15]. recommendation [BACB20, CSSS21, FJ20, ISY20, KD21, KS22, NHP^{+}21, SK15, WR18]. recommendations [BACB20, BGAE24, GPT05]. recommender [BGLB16, SC05]. record [WIMM24]. recordings [KAMKJ19]. Records [Gri23a, Gri23b, ADT^{+}19, DBG^{+}16, LRT21, RGÓ^{+}15, TLJ^{+}02]. rediscovery [CBRCW05]. Reduction [OTA00]. Reference [BKS^{+}05, FCR20, ABDT^{+}17, LGT97, MBVL18, PKK18, SB14]. references [BA18, CRK18]. related [BACB20, SJ21]. related-article [BACB20]. relation [JDAD22]. relations [AIL16, PMKD19, SZOS22]. relationship [OSD23]. relationships [LPSW06]. Relevance [HPSB10, BKSR10, SSK^{+}24]. relevant [BKG^{+}97]. reliability [Fig21, MR21, NH17]. remote [YRW^{+}05]. remote-sensing [YRW^{+}05]. Renwick [BGL05]. repeated [SB14]. replacement [KN14]. replicated [Has97]. report [GGM23]. reports [JTCA97]. repositories [GMP^{+}23, JBFNMD15, KSAP09, KRB21, OD09, VCH08, WBL^{+}07a, WBL^{+}07b]. repository [BLG^{+}07, Cap07, MBS21, NMT10, ZMM^{+}12]. representation [CSL^{+}97, CMM04, FRD23, PRS97, SBF^{+}17, SM12]. Representing [BDvdS06, CVHS20, NH16]. Republican [KBG23]. requests [LAP06]. requirements [ACDG10, BHS^{+}22, MPLR02, PWM^{+}05, SPLF02, SSNM97]. Research [ADH^{+}22, BGLB16, BHS^{+}22, BDF^{+}15, CAA^{+}07, CCMC17, CSSS21, Cho14, CBRCW05, DBG^{+}20, ELH^{+}17, EKV^{+}10, PETS12, RH05, SOM22, SMD00, SAS^{+}14, TBF^{+}04, TK20, TMTA07, VC17, VSVS08, Wat23, dSCGL20, vKdBV19]. Research-paper [BGLB16]. Researcher [KTT^{+}14]. researchers [DOTW97, KTT^{+}14, VC17]. resolution [BHD22, KRB21]. Resolver [KTT^{+}14]. Resource [Wit07, KM13, SGM^{+}22, WL97, vKdBV19]. resources [BKS^{+}15, CCCJ10, GN16, MNZD22a, MNZD22b, MIW^{+}08, Sch04, SAB^{+}05]. responses [CMM04]. responsibility [Fre14]. restoration [Can12, Lan15]. result [VH21]. Results [OYP^{+}16, BCGM00, PEC^{+}15]. Retrieval [WR18, ALT07, BCGM00, BKG^{+}97, BM05, BB17, CSL^{+}97, DBBK00, GCAN00, Hie00, KNH97, KPK^{+}24, LA00, MFC^{+}18, MSW^{+}05, MG06, MMK^{+}19, MKPI^{+}22, MSB^{+}17, OTA00, STvO^{+}18, SBF^{+}17, WGZ^{+}06, Whi18, ZG04, dOVA^{+}23, PCSH10]. retrospective [Fis21]. Reuse [MFP18, ABB^{+}19, GABD07, Zim07]. review [BGAE24, HDG19, KS22, STM^{+}20, ZM19]. reviews [KGE23]. revisited [HPSB10]. revoking [BDY02]. rhythm [CIMR12]. rich [GABD07]. RISM [KW19]. robust [ŁB15]. role [Hus19]. roles [BKS^{+}05, NA13]. Roman [KBG23]. Rule [KI04]. Rule-based [KI04]. rules [ARA23, BJWW04].

sandbox [PEC^{+}15]. satisfaction [ACZ06]. scalability [SPO08]. Scalable [CDN^{+}98, Kor16, GMP05, JBFNMD15, SOM22]. Scale [AGC05, KLL08, KRB21, NHP^{+}21, SFT22, VCH08]. scaling [YSZ06]. scarce [dSCGL20]. Scheduling [ZJ98]. schema [AFS00, LTCC13]. Schema.org [FRH^{+}20]. scheme [LLS16, MMT04]. schemes [Ke15]. SchenQL [KWK^{+}22]. Scholarly [KLL08, PC17, ASFF19, JDAD22, KB22, MKFK18, SGM^{+}22, SK15, WBL^{+}07a, WBL^{+}07b]. Scholars [NS00]. scholarship [Gri15, PEC^{+}15, WSMK07]. ScholOnto [SMD00]. schools [MHM08]. SciAnnotDoc [dRF18]. Science [CFP21, Di 23, BWE07, BDS^{+}15, FRD23, Gre97, Nic17a]. ScienceQA [SGM^{+}22]. Sciences [KPC^{+}21, PWB05, VC17]. Scientific [CG18, KYK23, KS22, ADL18, BPL22, BCLS00, BBF^{+}10, CD18, HC07, ISY20, JCRK18, JDAD22, KGJK14, KBFG19, LLSG08, ŁB15, PP20, PB20, Sch04, STM^{+}20, WSMK07, dRF18, CMT13]. scientist [SB04]. SciSumm [JCRK18]. score [BP19, FRT21]. Scores [RAA^{+}12]. Scripta [FM17]. scripts [ARA23]. SEA [GMP^{+}23]. Search [SS98, ACM^{+}98, BKSR10, BDF^{+}15, BKS^{+}06, CLvHM18, IMR12, JP05, KH17, KRZ18, OTA00, STvO^{+}18, VH21, WR18]. Search-by-content [SS98]. Searchable [WZ05]. searches [BCGM00, CC04]. searching [GCC24, LLS16, LKA^{+}17, PRS97, WWFW98]. Seattle [LTCC13]. Section [CD18]. secure [MMT04]. security [AR04]. seed [Whi18]. Seeking [KPC^{+}21, GHJ^{+}09, RTM05]. SEEM [VF21]. segmenting [HN07]. Selected [JMI^{+}22, ISK^{+}24, JKPC24]. selection [ADL18, dSBG21, BCGM00, SI21]. selective [BKG^{+}97]. Self [XI24]. Self-training [XI24]. Semantic [SBF^{+}17, AGC05, DBG^{+}16, EWSB20, FHKN20, GABD07, JP05, MG06, RT18, WLC^{+}19, XI24, HC06b, KSAP09, MT09, PWB05, RBE04, RPM^{+}15, ZM19]. semantic-space [XI24]. semantics [CC04, DMRA97, MAC^{+}15, SB23]. semi [XI24]. semi-supervised [XI24]. semiotic [FM17]. semistructured [AQM^{+}97]. SemWebDL [RBE04]. sense [TNZR13]. sensing [YRW^{+}05]. sensor [BWE07]. sentiment [KGE23]. series [BDF^{+}15]. server [CGM05, MPLR02, SMD00]. servers [KJ05, ZYZ^{+}00]. service [DSGF06, EMS^{+}98, ITSY06, KTS00, LAP06, Pap06]. service-oriented [Pap06]. services [GN16, HC06b, KW06, Sat06, SBC04, TCK^{+}97, WB98]. sessions [SDLÅ14]. set [CVDT22, WL97]. Shared [JCRK18]. SharedCanvas [SASV12]. Sharing [Bar07, KVZ17, SRGK16, VSVS08, ZYZ^{+}00]. sharing-oriented [SRGK16]. Sheldon [JNV18]. shelf [VM17]. shifts [ALT07]. should [CN15]. Sifting [MHTH15]. signatures [BSZ04]. significant [cCWWK05]. signification [BA18]. silicon [SC23]. similar [LLS16]. similarity [BKS^{+}06, EESB20]. Sitemaps [FRH^{+}20]. sites [FS00]. sliding [AN15]. small [Fad02]. smart [BDY02]. Social [KPC^{+}21, AWN16a, ASFF19, FISN20, Gre97, LCK^{+}18, VC17]. Soft [Sto00]. software [NMT10]. solution [BLG^{+}07, Lee07, VF21]. Some [GBR17]. song [vKdBV19]. Sound [OOKS13]. sources [Fig21]. South [MBS21]. space [KK99, NH16, SKT98, WIMM24, XI24, ZG04]. SpaceGlue [ITSY06]. spaces [GMP05, ITSY06]. spans [MBVL18]. spatial [BNC12, BS97]. spatiotemporal [HDE17]. Special [Cru99, Gri23b, HN21, ISK^{+}24, JMI^{+}22, JKPC24, NS00, AR04, BM05, CGR08, FKX18, GST19, Gri15, KLL08, KMW16, KR15, LCN^{+}04, MTC^{+}16, MFC^{+}18, MSB^{+}17, NV06, PD19, Pap06]. specific [SGLF15, SDLÅ14]. specification [JMS99]. Speech [MFP18]. SPIRE [BCLS00]. spoilers [JNV18]. spoken [GRB^{+}05]. Stance [SSK^{+}24]. standard [RD10, RM19]. standardized [McD06]. standards [CFP21]. standing [RNFD16]. Stanford [BCGP97]. State [BWSH12]. State-of-the-art [BWSH12]. statements [Fre14]. status [Mar04]. sticky [AN15]. Stimulating [VSVS08]. stone [SC23]. storage [AGC05, YSZ06, Wit07]. stories [AWN16a]. storytelling [Buc23]. stratagem [CLvHM18]. strategies [dSBG21, BKG^{+}09, BHS^{+}22]. Strategy [LA00, KD21, SRGK16]. Strategy-based [LA00]. Strathprints [DWD12]. stress [LKR02]. string [PKK18]. structure [HPSB10, KMM19, KGJK14, KF98, PTB^{+}19, RT18]. Structured [DMRA97, CMM04, SSB^{+}12, SLS19]. structures [jCZhC16, LKA^{+}17]. structuring [HC06a]. study [ABDF07, Aya22, BPL22, BG12, BGL05, CNHB16, FMNH04, GWB^{+}08, HSPQ04, KSH^{+}99, Kor16, LHD08, NHP^{+}21, RSF08, RZ20, STM^{+}20, SB04, SFTD14, VM17, WHCF20, Wu16, dSRL19]. styles [CMM04]. styling [HCH04]. stylistic [KAMKJ19]. subject [NMWP99, TS20]. subseafloor [DBT^{+}15]. substantial [Hus19]. subtitles [Hus19]. successes [HWT17]. summaries [Wes10]. summarization [ADL18, ANV^{+}16, dSBG21, CG18, CD18, JCRK18, LMZ^{+}18, MRY06, RT18]. sunk [Fad02]. supervised [BP19, XI24]. Supplementing [ZBS^{+}10]. supplied [Wes10]. Support [DIJ^{+}06, Di 23, DLR02, FBR07, Lee07, PB20, SMK^{+}14, WL97]. Supporting [KH17, EKV^{+}10, KFRF^{+}04, TBF^{+}04, Wan99]. survey [BGLB16]. Sustainability [Cho14]. sustainable [RH05]. Symbiosis [OOKS13]. system [BDF^{+}15, BCMC23, BDY02, CSSS21, CSL^{+}97, FBR07, HN07, HCHH05, KSG10, KDP14, KPK^{+}24, KTT^{+}14, Lee07, MSW^{+}05, MS14, NPS05, RTM05, SS06, STM^{+}20, SBC04, SLM04, STP07, SKT98, Wan99]. Systematic [BKG^{+}09, KDA24]. Systems [KMM19, PMKD19, Wu16, ZM19, ABDF07, BGLB16, GHJ^{+}09, KS22, MPLR02, PCSH10, SRGK16, SC05, dSTMGF06, WRSB23, YSZ06, GST19, MTC^{+}16].

tags [ŁB15]. Tape [FBPC17]. target [AN15]. Task [KRZ18, JCRK18]. Task-oriented [KRZ18]. tasks [KH17]. Teaching [PAM09]. technical [SC23]. Techniques [SJ09, EMS^{+}98, LW09]. technologies [CAA^{+}07, CBRCW05, FRH^{+}20, KNH97, MT09, PWB05, TCK^{+}97, TMTA07, WR18]. technology [IMA^{+}05, WHK^{+}97, Sch04]. telematic [TCK^{+}97]. telos [Sch04]. template [KGE23]. template-based [KGE23]. temporal [AN15, FISN20, SZOS22]. term [BWSH12, DVP09, Hie00, Ke15, PETS12, SJ09]. terminology [BKG^{+}97, SLS19]. terms [FAM00, OTA00, SOM22, dSRL19]. Territoires [MR21]. test [OYP^{+}16]. Text [SH00, WDDT04, BA18, DWD12, JTCA97, OTA00, OFM^{+}00, RT18, ZG04]. Text-based [SH00]. texts [FM17]. textual [DBBK00, LRT21]. tf [Hie00]. TF*IDF [Whi18]. their [BA18, Hus19, KBFG19, LPSW06, MKFK18]. theoretic [Ke15]. Theory [AP14, ADH^{+}22, FKRN19, Aya22, GFW08, TK20]. thesauri [AFS00, KNH97, NH16]. Thesaurus [KNH97, dBOS16]. thesaurus-based [dBOS16]. theses [GMP^{+}23, RSF08]. thing [GBR17]. Thinking [Bar21]. three [KK99, MSHC04]. three-dimensional [KK99]. Time [JKT19, BBF^{+}10, BDF^{+}15, JGL^{+}97, JNV18, KN14, NH16, ZBS^{+}10, Dil99]. Time-focused [JKT19]. time-oriented [BBF^{+}10]. TimeMaps [AWN16b]. timeouts [KTS00]. Tipple [HB16]. tolerant [KF98]. tool [LGT97, MHML05, MBM19]. tools [CFP21, EKV^{+}10, KVZ17, ŁB15, OSCF09]. top [AWNV14]. top-level [AWNV14]. topic [AWN16b, ALT07, MBVL18, TPMG15]. topical [SH00]. topics [SOM22]. Tours [MR21]. TPDL [AP14, FKRN19, ADH^{+}22, KMW16, TK20]. traceability [MR21]. track [GN16]. Tracking [FISN20, CS98]. tradition [NS00]. traditional [KAMKJ19]. trails [HC07]. training [XI24]. transaction [JCMB00, KD21]. transcription [MS14]. Transitioning [BLG^{+}07]. translation [CBC^{+}23a, CBC^{+}23b]. translations [WLC04]. Transliterating [ARA23]. travel [JNV18]. TRECVid [OOKS13]. triggered [HB16]. trust [SBC04]. turn [MM13]. turning [LW09]. tweets [YKP20]. Twitter [SJ21].

U.S. [MHM08]. ubiquitous [Sat06]. unacceptable [AB04]. unarchived [HKS^{+}15]. Uncovering [CCvdM08]. underlying [KDP14]. Understanding [ACDG10, MHM08]. Unified [Fis21, WLC^{+}19]. units [Gre97]. universe [DMRA97]. universities [GMP^{+}23]. University [Sch04, MBS21]. unobtrusively [CGM05]. unseen [AB04]. Unsupervised [KGJK14, KGE23, SOM22, dBOS16]. untapped [Kel24]. updating [SFTD14]. Usability [BBJ04, HSPQ04, KSH^{+}99, DVP09]. usage [KDP14, MSS12, SFT22]. usage-based [KDP14]. use [BHS^{+}22, GWB^{+}08, MHM08, MIW^{+}08, MKFK18, Zim07, dRF18]. useful [MHTH15]. User [GCC24, SAPN00, ARA23, ACDG10, AOP18, BTN12, BDF^{+}15, BCMC23, CMM04, Cru99, DLR02, EKV^{+}10, JMS99, KD21, Pap06, SDLÅ14, Wes10, WRSB23, dSRL19, dOGM99]. user-centered [BDF^{+}15, Pap06]. user-defined [ARA23]. user-driven [BTN12]. user-oriented [WRSB23]. user-supplied [Wes10]. user-transaction-based [KD21]. users [WHCF20]. Using [DWD12, GN16, KJ05, MAC^{+}15, NMSK07, ARA23, ADL18, ABDT^{+}17, BM16, BDvdS06, BG12, BM20, BT16, CC04, ELH^{+}17, FRH^{+}20, HMG20, HR97, Hie00, HB16, HC06b, KM13, LHD08, MT09, MRY06, PAM09, SFTD14, SK15, WWFW98, WR18, YG17, dSRL19, dRF18].

validated [SLS19]. value [FAM00].

value/NC [FAM00]. Vargas [RBdP^{+}15]. vectors [LKR02]. venues [PP20]. verge [CGR08]. VERS [Wau06]. version [FHKN20]. versions [KBFG19]. versus [GCC24]. very [KLL08]. VeTo [CVDT22]. VI [VF21]. via [CG18]. Video [LS12, dSBG21, LTCC13, MG06, MMK^{+}19, MKPI^{+}22, MRY06, OOKS13, SLM04, Wes10, Mar04]. videos [HC06a]. views [JP05]. virtual [HLZ05, ST04, ZBS^{+}10]. VisInfo [BDF^{+}15]. vision [CMT13, OOKS13]. Visions [TBC^{+}15]. Visual [DBBK00, MSW^{+}05, BDAB18, BBF^{+}10, CESF99, GHJ^{+}09, MKPI^{+}22, NMT10, RTM05, Wes10]. visual-information-seeking [RTM05]. visualising [MBM19, BKSR10]. visualization [BM05, KH17, LA00, MHML05, SKS^{+}99]. visualizations [SAB^{+}05]. visualizing [CARM20]. vital [DBG^{+}16]. VIVA [MKPI^{+}22]. vocabularies [jCZhC16]. vocabulary [BT16]. volumetric [SKS^{+}99].

wake [ASFF19]. walks [AN15]. War [MTLH17]. watermark [HN07]. wavelet [DWD12]. wavelets [WWFW98]. way [MR21]. weaving [HJK^{+}09]. web [AN15, AWNV14, Aya22, BG12, CGS17, FHKN20, FMN18, FKX18, FRH^{+}20, HKS^{+}15, KJ05, KN14, KB22, LAP06, LCK^{+}18, NMSK07, ACDG10, AWN16b, ANV^{+}16, BDAB18, CGM05, FS00, GRD20, HJK^{+}09, HC06b, KJ05, KSAP09, MT09, MM13, MMM97, PWB05, RTM05, RBE04, STvO^{+}18, SBC04, WLC04, WZ05, ZM19]. website [KG21, WHCF20, BM16]. weighting [Hie00, Ke15, Whi18]. Who [AAWN14]. Wide [CGM05, MMM97]. widely [JGL^{+}97]. wiki [JNV18]. Wikidata [MBM19]. Wikidata-based [MBM19]. Wikipedia [JKT19, ŁB15]. within [AWN16b, CDN^{+}98, CCGS14, Has97]. Word [Fre14, EWSB20, FAM00, GRB^{+}05, MBVL18, TNZR13]. words [WZ05]. work [FMN18, Fre14]. workflow [CCGS14, HWT17, KDA24, WTH17, GWB^{+}08]. workflows [ADT00]. working [Cap07, GPT05, Sch04]. Workload [LLSG08]. works [Whi18]. world [JBFNMD15, CGM05, GGM23, MTLH17, MMM97]. would [FMN18]. wrappers [FS00]. wrapping [VCC00]. WW1LOD [MTLH17]. www.neumann [Tós04]. www.neumann-haz.hu [Tós04].

X [LGT97]. X-ray [LGT97]. X3ML [MMK^{+}17]. XML [ALT07, DIJ^{+}06, HPSB10, MSW^{+}05].

years [TNZR13]. yield [SPLF02]. YouTube [MHTH15]. Yugoslav [BDAB18].

Z39.50 [VCC00]. zero [GTP^{+}99]. zero-hit [GTP^{+}99].

References

AlNoamany:2014:WWL

[AAWN14] Yasmin AlNoamany, Ahmed AlSum, Michele C. Weigle, and Michael L. Nelson. Who and what links to the Internet Archive. International Journal on Digital Libraries, 14 (3-4):101–115, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                                    tronic). URL https://link.
                                                    springer.com/article/10.
                                                    1007/s00799-014-0111-5.

Adams:2004:UUF

[AB04] Anne Adams and Ann Blandford. The unseen and unacceptable face of digital libraries. International Journal on Digital Libraries, 4(2):71–81, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                                    tronic). URL https://link.
                                                    springer.com/article/10.
                                                    1007/s00799-003-0071-7.

Adamou:2019:CLD

[ABB^{+}19] Alessandro Adamou, Simon Brown, Helen Barlow, Carlo Allocca, and Mathieu d’Aquin. Crowdsourcing Linked Data on listening experiences through reuse and enhancement of library data. International Journal on Digital Libraries, 20 (1):61–79, March 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                                    tronic). URL https://link.
                                                    springer.com/article/10.
                                                    1007/s00799-018-0235-0.

Agosti:2007:HCS

[ABDF07] Maristella Agosti, Giorgetta Bonfiglio-Dosio, and Nicola Ferro. A historical and contemporary study on annotations to derive key features for systems design. International Journal on Digital Libraries, 8(1):1–19, November 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s00799-007-0010-0.

Austin:2017:KCD

[ABDT^{+}17] Claire C. Austin, Theodora Bloom, Sünje Dallmeier-Tiessen, Varsha K. Khodiyar, Fiona Murphy, Amy Nurnberger, Lisa Raymond, Martina Stockhause, Jonathan Tedds, Mary Vardigan, and Angus Whyte. Key components of data publishing: using current best practices to develop a reference model for data publishing. International Journal on Digital Libraries, 18(2):77–92, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s00799-016-0178-2.

Abiteboul:1997:QDO

[ACC^{+}97] Serge Abiteboul, Sophie Cluet, Vassilis Christophides, Tova Milo, Guido Moerkotte, and Jérôme Siméon. Querying documents in object databases. International Journal on Digital Libraries, 1(1):5–19, April 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s007990050001.

Agosti:2010:UUR

[ACDG10] Maristella Agosti, Franco Crivellari, Giorgio Maria Di Nunzio, and Silvia Gabrielli. Understanding user requirements and preferences for a digital library Web portal. International Journal on Digital Libraries, 11(4):225–238, December 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s00799-011-0075-7.

Aho:1998:CDN

[ACM^{+}98] Alfred V. Aho, Shih-Fu Chang, Kathleen R. McKeown, Dragomir R. Radev, John R. Smith, and Kazi A. Zaman. Columbia Digital News Project: an environment for briefing and search over multimedia information. International Journal on Digital Libraries, 1(4): 377–385, March 1998. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.
                                        springer.com/article/10.
                                        1007/s007990050030.

Abidi:2006:GCY

[ACZ06] Syed Sibte Raza Abidi, Yong Han Chong, and Yan Zeng. Generating customized yet factually consistent information: a constraint satisfaction approach. International Journal on Digital Libraries, 6(3): 247–259, June 2006. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.
springer.com/article/10.
1007/s00799-006-0003-4.

Aalberg:2022:CRT

[ADH^{+}22] Trond Aalberg, Fabien Duchateau, Mark Hall, Tanja Mercŭn, andT homasRisse.Currentresearchontheoryandpracticeofdigitallibraries : BestpapersfromT P DL2019&2020.InternationalJournalonDigitalLibraries, 23

(1) : 1−−3, March2022.CODEN????ISSN1432−

1300(print), 1432−5012(electronic).URLhttpslinkspringercomarticle101007s00799022003225.

AlSaied:2018:ASS

[ADL18] Hazem Al Saied, Nicolas Dugué, and Jean-Charles Lamirel. Automatic summarization of scientific publications using a feature selection approach. International Journal on Digital Libraries, 19(2-3):203–215, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-017-0214-x.

Arpinar:2000:OEM

[ADT00] Sena Arpinar, Asuman Dogac, and Nesime Tatbul. An open electronic marketplace through agent-based workflows: MOP-PET. International Journal on Digital Libraries, 3(1):36–59, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00021474.

Aalberg:2019:BEI

[ADT^{+}19] Trond Aalberg, Fabien Duchateau, Naimdjon Takhirov, Joffrey Decourselle, and Nicolas Lumineau. Benchmarking and evaluating the interpretation of bibliographic records. International Journal on Digital Libraries, 20(2):143–165, June 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0233-2.

Asghari:2021:CAC

[AFMF21] Habibollah Asghari, Omid Fatemi, Salar Mohtaj, and Heshaam Faili. A crowdsourcing approach to construct mono-lingual plagiarism detection corpus. International Journal on Digital Libraries, 22(1):49–61, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-020-00294-4.

Amann:2000:IOT

[AFS00] Bernd Amann, Irini Fundulaki, and Michel Scholl. Integrating ontologies and thesauri for RDF schema creation and metadata querying. International Journal on Digital Libraries, 3(3):221–236, October 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990000037.

Anastasiadis:2005:SPS

[AGC05] Stergios V. Anastasiadis, Syam Gadde, and Jeffrey S. Chase. Scale and performance in semantic storage management of data grids. International Journal on Digital Libraries, 5(2):84–98, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0086-8.

Ashish:2005:I

[AGS05] Naveen Ashish, Carole Goble, and Amit Sheth. Introduction. International Journal on Digital Libraries, 5(2):71, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0084-x.

Alexiev:2016:CIH

[AIL16] Vladimir Alexiev, Antoine Isaac, and Jutta Lindenthal. On the composition of ISO 25964 hierarchical relations (BTG, BTP, BTI). International Journal on Digital Libraries, 17(1):39–48, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0162-2.

Ashoori:2007:ETS

[ALT07] Elham Ashoori, Mounia Lalmas, and Theodora Tsikrika. Examining topic shifts in content-oriented XML retrieval. International Journal on Digital Libraries, 8(1):39– 60, November 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-007-0026-5.

Ainsworth:2015:ESS

[AN15] Scott G. Ainsworth and Michael L. Nelson. Evaluating sliding and sticky target policies by measuring temporal drift in acyclic walks through a web archive. International Journal on Digital Libraries, 16 (2):129–144, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0120-4.

Anonymous:1997:F

[Ano97] Anonymous. Foreword. International Journal on Digital Libraries, 1(3):189–191, December 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990050015.

Anonymous:1998:I

[Ano98] Anonymous. Introduction. International Journal on Digital Libraries, 1(4):309–310, March 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990050025.

Anonymous:2000:E

[Ano00] Anonymous. Editorial. International Journal on Digital Libraries, 3(3):207, October 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990000039.

Alam:2016:WAP

[ANV^{+}16] Sawood Alam, Michael L. Nelson, Herbert Van de Sompel, Lyudmila L. Balakireva, Harihar Shankar, and David S. H. Rosenthal. Web archive profiling through CDX summarization. International Journal on Digital Libraries, 17(3):223–238, September 2016. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-016-0184-4.

Agosti:2018:PUE

[AOP18] Maristella Agosti, Nicola Orio, and Chiara Ponchia. Promoting user engagement with digital cultural heritage collections. International Journal on Digital Libraries, 19(4): 353–366, November 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0245-y.

Aalberg:2014:IFI

[AP14] Trond Aalberg and Christos Papatheodorou. Introduction to the focused issue on the 17th International Conference on Theory and Practice of Digital Libraries (TPDL 2013). International Journal on Digital Libraries, 14 (3-4):81–82, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0119-x.

Abiteboul:1997:LQL

[AQM^{+}97] Serge Abiteboul, Dallan Quass, Jason McHugh, Jennifer Widom, and Janet L. Wiener. The Lorel query language for semistructured data. International Journal on Digital Libraries, 1(1):68–88, April 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050005.

Atluri:2004:EMS

[AR04] Vijayalakshmi Atluri and Indrakshi Ray. Editors’ message for the special issue on security. International Journal on Digital Libraries, 4(3):137–138, November 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-004-0087-7.

Abebaw:2023:TLA

[ARA23] Zeleke Abebaw, Andreas Rauber, and Solomon Atnafu. Transliterating Latin to Amharic scripts using user-defined rules and character mappings. International Journal on Digital Libraries, 24(1):63–75, March 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-023-00346-5.

Agnew:2002:FMD

[AS02] G. B. Agnew and A. Sivanandan. A fast method for determining the origins of documents based on LZW compression. International Journal on Digital Libraries, 3 (4):297–301, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100043.

Alhoori:2019:ASI

[ASFF19] Hamed Alhoori, Mohammed Samaka, Richard Furuta, and Edward A. Fox. Anatomy of scholarly information behavior patterns in the wake of academic social media platforms. International Journal on Digital Libraries, 20(4):369–389, December 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0255-9.

AlNoamany:2016:CSM

[AWN16a] Yasmin AlNoamany, Michele C. Weigle, and Michael L. Nelson. Characteristics of social media stories. International Journal on Digital Libraries, 17(3):239–256, September 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-016-0185-3.

AlNoamany:2016:DTP

[AWN16b] Yasmin AlNoamany, Michele C. Weigle, and Michael L. Nelson. Detecting off-topic pages within TimeMaps in Web archives. International Journal on Digital Libraries, 17(3): 203–221, September 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0183-5.

AlSum:2014:PWA

[AWNV14] Ahmed AlSum, Michele C. Weigle, Michael L. Nelson, and Herbert Van de Sompel. Profiling web archive coverage for top-level domain and content language. International Journal on Digital Libraries, 14(3-4):149–166, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0118-y.

Ayala:2022:CPM

[Aya22] Brenda Reyes Ayala. Correspondence as the primary measure of information quality for web archives: a human-centered grounded theory study. International Journal on Digital Libraries, 23(1):19–31, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00314-x.

Bertin:2018:CMT

[BA18] Marc Bertin and Iana Atanassova. The context of multiple in-text references and their signification. International Journal on Digital Libraries, 19(2-3):127–138, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-017-0225-7.

Beierle:2020:COR

[BACB20] Felix Beierle, Akiko Aizawa, Andrew Collins, and Joeran Beel. Choice overload and recommendation effectiveness in related-article recommendations. International Journal on Digital Libraries, 21(3):231–246, September 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-019-00270-7.

Baru:2007:SCE

[Bar07] Chaitanya Baru. Sharing and caring of eScience data. International Journal on Digital Libraries, 7(1-2):113–116, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-007-0029-2.

Barbuti:2021:TDL

[Bar21] Nicola Barbuti. Thinking digital libraries for preservation as digital cultural heritage: by R to R^{4} facet of FAIR principles. International Journal on Digital Libraries, 22 (3):309–318, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00291-7.

Brath:2017:FAE

[BB17] Richard Brath and Ebad Banissi. Font attributes enrich knowledge maps and information retrieval. International Journal on Digital Libraries, 18(1):5–24, March 2017. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0168-4.

Bernard:2010:VDL

[BBF^{+}10] Jürgen Bernard, Jan Brase, Dieter Fellner, Oliver Koepler, Jörn Kohlhammer, Tobias Ruppert, Tobias Schreck, and Irina Sens. A visual digital library approach for time-oriented scientific primary data. International Journal on Digital Libraries, 11 (2):111–123, June 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0072-x.

Blandford:2004:UDL

[BBJ04] Ann Blandford, George Buchanan, and Matt Jones. Usability of digital libraries. International Journal on Digital Libraries, 4 (2):69–70, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0070-8.

Borbinha:2006:MCM

[BcCC06] José Borbinha, Ching chih Chen, and Stavros Christodoulakis. Multimedia contents and management. International Journal on Digital Libraries, 6(1):1–2, February 2006. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-005-0126-z.

Berenci:2000:EKB

[BCGM00] Ezio Berenci, Claudio Carpineto, Vittorio Giannini, and Stefano Mizzaro. Effectiveness of keyword-based display and selection of retrieval results for interactive searches. International Journal on Digital Libraries, 3 (3):249–260, October 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990000035.

Baldonado:1997:SDL

[BCGP97] Michelle Baldonado, Chen-Chuan K. Chang, Luis Gravano, and Andreas Paepcke. The Stanford Digital Library metadata architecture. International Journal on Digital Libraries, 1(2):108–121, September 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050008.

Bergman:2000:SDL

[BCLS00] Lawrence D. Bergman, Vittorio Castelli, Chung-Sheng Li, and John R. Smith. SPIRE: a digital library for scientific information. International Journal on Digital Libraries, 3(1):85–99, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00021478.

Bernasconi:2023:DRU

[BCMC23] Eleonora Bernasconi, Miguel Ceriani, Massimo Mecella, and Tiziana Catarci. Design, realization, and user evaluation of the ARCA system for exploring a digital library. International Journal on Digital Libraries, 24(1):1–22, March 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00343-0.

Ben-David:2018:CNW

[BDAB18] Anat Ben-David, Adam Amram, and Ron Bekkerman. The colors of the national Web: visual data analysis of the historical Yugoslav Web domain. International Journal on Digital Libraries, 19(1):95–106, March 2018. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0202-6.

Bernard:2015:VDL

[BDF^{+}15] Jürgen Bernard, Debora Daberkow, Dieter Fellner, Katrin Fischer, Oliver Koepler, Jörn Kohlhammer, Mila Runnwerth, Tobias Ruppert, Tobias Schreck, and Irina Sens. VisInfo: a digital library system for time series research data based on exploratory search — a user-centered design approach. International Journal on Digital Libraries, 16(1):37–59, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0134-y.

Borgman:2015:KIS

[BDS^{+}15] Christine L. Borgman, Peter T. Darch, Ashley E. Sands, Irene V. Pasquetto, Milena S. Golshan, Jillian C. Wallis, and Sharon Traweek. Knowledge infrastructures in science: data, diversity, and digital libraries. International Journal on Digital Libraries, 16(3-4):207–227, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0157-z.

Bekaert:2006:RDA

[BDvdS06] Jeroen Bekaert, Emiel De Kooning, and Herbert van de Sompel. Representing digital assets using MPEG-21 Digital Item Declaration. International Journal on Digital Libraries, 6(2):159–173, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0133-0.

Bo:2002:ARP

[BDY02] Yang Bo, Liu Dongsu, and Wang Yumin. An anonymity-revoking e-payment system with a smart card. International Journal on Digital Libraries, 3(4):291–296, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100047.

BenSaad:2012:AWU

[BG12] Myriam Ben Saad and Stéphane Gançarski. Archiving the web using page changes patterns: a case study. International Journal on Digital Libraries, 13(1):33–49, December 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-012-0094-z.

Bharti:2024:PAB

[BGAE24] Prabhat Kumar Bharti, Tirthankar Ghosal, Mayank Agarwal, and Asif Ekbal. PEERRec: an AI-based approach to automatically generate recommendations and predict decisions in peer review. International Journal on Digital Libraries, 25(1):55–72, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00375-0.

Boehner:2005:DED

[BGL05] Kirsten Boehner, Geri Gay, and Claire Larkin. Drawing evaluation into design for mobile computing: a case study of the Renwick Gallery’s Hand Held Education Project. International Journal on Digital Libraries, 5 (3):219–230, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0107-7.

Beel:2016:RPR

[BGLB16] Joeran Beel, Bela Gipp, Stefan Langer, and Corinna Breitinger. Research-paper recommender systems: a literature survey. International Journal on Digital Libraries, 17 (4):305–338, November 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0156-0.

Backes:2022:THA

[BHD22] Tobias Backes, Daniel Hienert, and Stefan Dietze. Towards hierarchical affiliation resolution: framework, baselines, dataset. International Journal on Digital Libraries, 23(3):267–288, September 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00326-1.

Brack:2022:ARO

[BHS^{+}22] Arthur Brack, Anett Hoppe, Markus Stocker, Sören Auer, and Ralph Ewerth. Analysing the requirements for an Open Research Knowledge Graph: use cases, quality requirements, and construction strategies. International Journal on Digital Libraries, 23 (1):33–55, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00306-x.

Bettini:2004:RAP

[BJWW04] Claudio Bettini, Sushil Jajodia, X. Sean Wang, and Duminda Wijesekera. Reasoning with advanced policy rules and its application to access control. International Journal on Digital Libraries, 4(3): 156–170, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0078-8.

Bidgood:1997:CTC

[BKG^{+}97] W. Dean Bidgood, Louis Y. Korman, Alan M. Golichowski, P. Lloyd Hildebrand, Angelo Rossi Mori, Bruce Bray, Nicholas J. G. Brown, Kent A. Spackman, S. Brent Dove, and Katherine Schoeffler. Controlled terminology for clinically-relevant indexing and selective retrieval of biomedical images. International Journal on Digital Libraries, 1 (3):278–287, December 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050022.

Becker:2009:SPD

[BKG^{+}09] Christoph Becker, Hannes Kulovits, Mark Guttenbrunner, Stephan Strodl, Andreas Rauber, and Hans Hofman. Systematic planning for digital preservation: evaluating potential strategies and building preservation plans. International Journal on Digital Libraries, 10(4):133–157, December 2009. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-009-0057-1.

Borbinha:2005:RMD

[BKS^{+}05] José Borbinha, John Kunze, Angela Spinazzè, Peter Mutschke, Hans-Jörg Lieder, Michael Mabe, Larry Dixson, Howard Besser, Becky Dean, and Warwick Cathro. Reference models for digital libraries: actors and roles. International Journal on Digital Libraries, 5 (4):325–330, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0095-7.

Bustos:2006:EEC

[BKS^{+}06] Benjamin Bustos, Daniel Keim, Dietmar Saupe, Tobias Schreck, and Dejan Vranić. An experimental effectiveness comparison of methods for 3D similarity search. International Journal on Digital Libraries, 6 (1):39–54, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0122-3.

Brunelle:2015:AMC

[BKS^{+}15] Justin F. Brunelle, Mat Kelly, Hany SalahEldeen, Michele C. Weigle, and Michael L. Nelson. Not all mementos are created equal: measuring the impact of missing resources. International Journal on Digital Libraries, 16(3-4):283–301, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0150-6.

Beresi:2010:WDY

[BKSR10] Ulises Cerviño Beresi, Yunhyong Kim, Dawei Song, and Ian Ruthven. Why did you pick that? Visualising relevance criteria in exploratory search. International Journal on Digital Libraries, 11(2):59–74, June 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-011-0067-7.

Brunelle:2016:IJA

[BKWN16] Justin F. Brunelle, Mat Kelly, Michele C. Weigle, and Michael L. Nelson. The impact of JavaScript on archivability. International Journal on Digital Libraries, 17 (2):95–117, June 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0140-8.

Barros:2007:TEF

[BLG^{+}07] Evandrino G. Barros, Alberto H. F. Laender, Marcos André Gonçalves, Ricardo G. Cota, and Francisco A. R. Barbosa. Transitioning from the ecological fieldwork to an online repository: a digital library solution and evaluation. International Journal on Digital Libraries, 7(1-2):109–112, October 2007. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-007-0024-7.

Borner:2005:JSI

[BM05] Katy Börner and Javed Mostafa. JoDL special issue on information visualization interfaces for retrieval and analysis — guest editor’s introduction. International Journal on Digital Libraries, 5(1):1–2, March 2005. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-005-0114-3.

Banos:2016:QAE

[BM16] Vangelis Banos and Yannis Manolopoulos. A quantitative approach to evaluate Website Archivability using the CLEAR+ method. International Journal on Digital Libraries, 17(2):119–141, June 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0144-4.

BinMakhashen:2020:HDL

[BM20] Galal M. BinMakhashen and Sabri A. Mahmoud. Historical document layout analysis using anisotropic diffusion and geometric features. International Journal on Digital Libraries, 21(3):329–342, September 2020. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-020-00280-
w.

Bia:2010:DDC

[BMG10] Alejandro Bia, Rafael Muñoz, and Jaime Gómez. DiCoMo: the digitization cost model. International Journal on Digital Libraries, 11(2):141–153, June 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0073-9.

Bainbridge:2012:SHB

[BNC12] David Bainbridge, Brook J. Novak, and Sally Jo Cunningham. A spatial hypertextbased, personal digital library for capturing and organizing musical moments. International Journal on Digital Libraries, 12 (2-3):89–103, August 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-012-0090-3.

Boros:2022:AIO

[BNLD22] Emanuela Boros, Nhu Khoa Nguyen, Gaël Lejeune, and Antoine Doucet. Assessing the impact of OCR noise on multilingual event detection over digitised documents. International Journal on Digital Libraries, 23(3):241–266, September 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00325-2.

Borbinha:2004:ACN

[Bor04] José Luis Borbinha. An approach to creating a national digital library. International Journal on Digital Libraries, 4 (1):19–22, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0067-3.

Bell:2019:HSA

[BP19] Eamonn Bell and Laurent Pugin. Heuristic and supervised approaches to handwritten annotation extraction for musical score images. International Journal on Digital Libraries, 20(1):49–59, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0249-7.

Bartalesi:2022:LDS

[BPL22] Valentina Bartalesi, Nicolò Pratelli, and Emanuele Lenzi. Linking different scientific digital libraries in Digital Humanities: the IMAGO case study. International Journal on Digital Libraries, 23(4): 303–317, December 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00331-4.

Beard:1997:MRD

[BS97] Kate Beard and Vyjayanti Sharma. Multidimensional ranking for data in digital spatial libraries. International Journal on Digital Libraries, 1(2):153–160, September 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050011.

Bergman:1999:DDM

[BSC^{+}99] Lawrence D. Bergman, Jerre Shoudt, Vittorio Castelli, Chung-Sheng Li, and Loey Knapp. Drag-and-drop multimedia: an interface framework for digital libraries. International Journal on Digital Libraries, 2 (2-3):178–189, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050046.

Bull:2004:HEP

[BSZ04] Laurence Bull, David McG. Squire, and Yuliang Zheng. A Hierarchical Extraction Policy for content extraction signatures. International Journal on Digital Libraries, 4 (3):208–222, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0082-z.

Binding:2016:IIU

[BT16] Ceri Binding and Douglas Tudhope. Improving interoperability using vocabulary linked data. International Journal on Digital Libraries, 17(1):5–21, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0166-y.

Bainbridge:2012:ICA

[BTN12] David Bainbridge, Michael B. Twidale, and David M. Nichols. Interactive context-aware user-driven metadata correction in digital libraries. International Journal on Digital Libraries, 13(1):17–32, December 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-012-0100-5.

Buchanan:2023:ENP

[Buc23]  Sarah A. Buchanan.  Envisioning net-

worked provenance data storytelling with American cuneiform collections. International Journal on Digital Libraries, 24(3): 149–158, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00335-0.

Borgman:2007:LSC

[BWE07] Christine L. Borgman, Jillian C. Wallis, and Noel Enyedy. Little science confronts the data deluge: habitat ecology, embedded sensor networks, and digital libraries. International Journal on Digital Libraries, 7 (1-2):17–30, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0022-9.

Brunsmann:2012:SAL

[BWSH12] Jörg Brunsmann, Wolfgang Wilkes, Gunter Schlageter, and Matthias Hemmje. Stateof-the-art of long-term preservation in product lifecycle management. International Journal on Digital Libraries, 12(1):27–39, July 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-012-0081-4.

Candela:2007:DID

[CAA^{+}07] Leonardo Candela, Fuat Akal, Henri Avancini, Donatella Castelli, Luigi Fusco, Veronica Guidetti, Christoph Langguth, Andrea Manzi, Pasquale Pagano, Heiko Schuldt, Manuele Simi, Michael Springmann, and Laura Voicu. DILIGENT: integrating digital library and Grid technologies for a new Earth observation research infrastructure. International Journal on Digital Libraries, 7(1-2):59–80, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0023-8.

Canazza:2012:DCE

[Can12] Sergio Canazza. The digital curation of ethnic music audio archives: from preservation to restoration. International Journal on Digital Libraries, 12(2-3):121–135, August 2012. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-012-0088-x.

Caplan:2007:FDA

[Cap07] Priscilla Caplan. The Florida Digital Archive and DAITSS: a working preservation repository based on format migration. International Journal on Digital Libraries, 6 (4):305–311, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0009-6.

Cole:2020:FMV

[CARM20] Nicholas Cole, Alfie Abdul-Rahman, and Grace Mallon. A framework for modelling and visualizing the US Constitutional Convention of 1787. International Journal on Digital Libraries, 21(2):191–213, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0263-9.

Crane:2007:EH

[CBB07] Gregory Crane, Alison Babeu, and David Bamman. eScience and the humanities. International Journal on Digital Libraries, 7 (1-2):117–122, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0031-8.

Crane:2023:BTE

[CBC^{+}23a] Gregory Crane, Alison Babeu, Lisa M. Cerrato, Amelia Parrish, Carolina Penagos, Faroosh Shamsian, James Tauber, and Jake Wegner. Beyond translation: engaging with foreign languages in a digital library. International Journal on Digital Libraries, 24 (3):163–176, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00349-2. See

correction [CBC^{+}23b].

Crane:2023:CBT

[CBC^{+}23b] Gregory Crane, Alison Babeu, Lisa M. Cerrato, Amelia Parrish, Carolina Penagos, Faroosh Shamsian, James Tauber, and Jake Wegner. Correction: Beyond translation: engaging with foreign languages in a digital library. International Journal on Digital Libraries, 24(3):177, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/

article/10.1007/s00799-023-00365-2. See

[CBC^{+}23a].

Crane:2005:ELT

[CBRCW05] Gregory Crane, Kalina Bontcheva, Jeffrey A. Rydberg-Cox, and Clifford Wulfman. Emerging language technologies and the rediscovery of the past: a research agenda. International Journal on Digital Libraries, 5 (4):309–316, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0096-6.

Cheng:2004:EIA

[CC04] Pu-Jen Cheng and Lee-Feng Chien. Effective image annotation for searches using multilevel semantics. International Journal on Digital Libraries, 4(4):258–271, November 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-004-0090-z.

Cabanac:2010:ODR

[CCCJ10] Guillaume Cabanac, Max Chevalier, Claude Chrisment, and Christine Julien. Organization of digital resources as an original facet for exploring the quiescent information capital of a community. International Journal on Digital Libraries, 11(4):239–261, December 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-011-0076-6.

Candela:2010:DLF

[CCFI10] Leonardo Candela, Donatella Castelli, Edward A. Fox, and Yannis Ioannidis. On Digital Library foundations. International Journal on Digital Libraries, 11(1):37–39, March 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-011-0063-y.

Ciocca:2014:HAI

[CCGS14] Gianluigi Ciocca, Silvia Corchs, Francesca Gasparini, and Raimondo Schettini. How to assess image quality within a workflow chain: an overview. International Journal on Digital Libraries, 15(1):1–25, November 2014. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-014-0124-0.

Candela:2017:RDP

[CCMC17] Leonardo Candela, Donatella Castelli, Paolo Manghi, and Sarah Callaghan. On research data publishing. International Journal on Digital Libraries, 18(2):73–75, June 2017. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-017-0213-y.

Cousins:2008:UCH

[CCvdM08] J. Cousins, S. Chambers, and E. van der Meulen. Uncovering cultural heritage through collaboration. International Journal on Digital Libraries, 9(2):125–138, November 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-008-0041-1.

Chen:2005:DIS

[cCWWK05] Ching chih Chen, Howard D. Wactlar, James Z. Wang, and Kevin Kiernan. Digital imagery for significant cultural and historical materials. International Journal on Digital Libraries, 5(4):275–286, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0097-5.

Conroy:2018:SMM

[CD18] John M. Conroy and Sashka T. Davis. Section mixture models for scientific document summarization. International Journal on Digital Libraries, 19(2-3):305–322, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-017-0218-6.

Cheng:1998:SAW

[CDN^{+}98] X. Cheng, R. Dolin, M. Neary, S. Prabhakar, K. V. Ravi Kanth, D. Wu, D. Agrawal, A. El Abbadi, M. Freeston, A. Singh, T. Smith, and J. Su. Scalable access within the context of digital libraries. International Journal on Digital Libraries, 1(4):358–376, March 1998. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050029.

Costabile:1999:AVE

[CESF99] M. F. Costabile, F. Esposito, G. Semeraro, and N. Fanizzi. An adaptive visual environment for digital libraries. International Journal on Digital Libraries, 2 (2-3):124–143, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050042.

Castelli:2021:HSC

[CFP21] Lisa Castelli, Achille Felicetti, and Fabio Proietti. Heritage science and cultural heritage: standards and tools for establishing cross-domain data interoperability. International Journal on Digital Libraries, 22(3): 279–287, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-019-00275-2.

Cohan:2018:SDS

[CG18] Arman Cohan and Nazli Goharian. Scientific document summarization via citation contextualization and scientific discourse. International Journal on Digital Libraries, 19 (2-3):287–303, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0216-8.

Cooper:2005:IUA

[CGM05] Brian F. Cooper and Hector Garcia-Molina. InfoMonitor: unobtrusively archiving a World Wide Web server. International Journal on Digital Libraries, 5(2):106–119, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-003-0052-x.

Cassel:2008:EDL

[CGR08] Lillian Cassel, Sarah Giersch, and Mimi Recker. Educational digital libraries on the verge: introduction to the special issue. International Journal on Digital Libraries, 9(1):1–2, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0040-2.

Costa:2017:EWA

[CGS17] Miguel Costa, Daniel Gomes, and Mário J. Silva. The evolution of web archiving. International Journal on Digital Libraries, 18(3):191–205, September 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0171-9.

Chowdhury:2014:SDL

[Cho14] Gobinda Chowdhury. Sustainability of digital libraries: a conceptual model and a research framework. International Journal on Digital Libraries, 14(3-4):181–195, August 2014. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-014-0116-0.

Chan:2012:EAR

[CIMR12] Joseph Wun-Tat Chan, Costas S. Iliopoulos, Spiros Michalakopoulos, and M. Sohel Rahman. Exact and approximate rhythm matching algorithms. International Journal on Digital Libraries, 12(2-3):149–158, August 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0085-0.

Crestani:2006:AFE

[CLM06] Fabio Crestani, Monica Landoni, and Massimo Melucci. Appearance and functionality of electronic books. International Journal on Digital Libraries, 6(2):192–209, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0113-9.

Clunis:2019:DOM

[Clu19] Julaine Clunis. Designing an ontology for managing the diets of hypertensive individuals. International Journal on Digital Libraries, 20(3):269–284, September 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0253-y.

Carevic:2018:IES

[CLvHM18] Zeljko Carevic, Maria Lusky, Wilko van Hoek, and Philipp Mayr. Investigating exploratory search activities based on the stratagem level in digital libraries. International Journal on Digital Libraries, 19 (2-3):231–251, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0226-6.

Chen:2004:CSU

[CMM04] Sherry Y. Chen, George D. Magoulas, and Robert D. Macredie. Cognitive styles and users’ responses to structured information representation. International Journal on Digital Libraries, 4(2):93–107, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0073-5.

Collins:2007:CEL

[CMM^{+}07] Linn Marks Collins, Mark L. B. Martinez, Ketan K. Mane, James E. Powell, Chad M. Kieffer, Tiago Simas, Susan K. Heckethorn, Kathryn R. Varjabedian, Miriam E. Blake, and Richard E. Luce. Collaborative eScience libraries. International Journal on Digital Libraries, 7(1-2):31–33, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0020-y.

Castelli:2013:VTS

[CMT13] Donatella Castelli, Paolo Manghi, and Costantino Thanos. A vision towards Scientific Communication Infrastructures. International Journal on Digital Libraries, 13 (3-4):155–169, September 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-013-0106-7.

Cartledge:2015:WSM

[CN15] Charles Cartledge and Michael L. Nelson. When should I make preservation copies of myself? International Journal on Digital Libraries, 16(3-4):183–205, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0155-1.

Cunningham:2016:WNE

[CNHB16] Sally Jo Cunningham, David M. Nichols, Annika Hinze, and Judy Bowen. What’s news? Encounters with news in everyday life: a study of behaviours and attitudes. International Journal on Digital Libraries, 17(3):257–271, September 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0187-1.

Colavizza:2023:CHC

[CPR23] Giovanni Colavizza, Silvio Peroni, and Matteo Romanello. The case for the Humanities Citation Index (HuCI): a citation index by the humanities, for the humanities. International Journal on Digital Libraries, 24(4):191–204, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00327-0.

Castelli:2004:OIN

[CPT04] Donatella Castelli, Pasquale Pagano, and Costantino Thanos. OpenDLib: an infrastructure for new generation digital libraries. International Journal on Digital Libraries, 4(1):45–47, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0061-9.

Crane:2023:PDL

[Cra23] Gregory Crane. The Perseus Digital Library and the future of libraries. International Journal on Digital Libraries, 24 (2):117–128, June 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00333-2.

Colavizza:2018:RRM

[CRK18] Giovanni Colavizza, Matteo Romanello, and Frédéric Kaplan. The references of references: a method to enrich humanities library catalogs with citation data. International Journal on Digital Libraries, 19 (2-3):151–161, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0210-1.

Cruz:1999:SIU

[Cru99] Isabel F. Cruz. Special issue on user interfaces for digital libraries. International Journal on Digital Libraries, 2(2-3):67, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050037.

Crabtree:1998:ITC

[CS98] I. Barry Crabtree and Stuart J. Soltysiak. Identifying and tracking changing interests. International Journal on Digital Libraries, 2(1):38–53, October 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050035.

Chen:1997:GKR

[CSL^{+}97] Hsinchun Chen, Terence R. Smith, Mary L. Larsgaard, Linda L. Hill, and Marshall Ramsey. A geographic knowledge representation system for multimedia geospatial retrieval and analysis. International Journal on Digital Libraries, 1(2):132–152, September 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050010.

Caled:2022:MLC

[CSMW22] Danielle Caled, Mário J. Silva, Bruno Martins, and Miguel Won. Multi-label classification of legislative contents with hierarchical label attention networks. International Journal on Digital Libraries, 23 (1):77–90, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-021-00307-w.

Chaudhuri:2021:HFI

[CSSS21] Arpita Chaudhuri, Nilanjan Sinhababu, Monalisa Sarma, and Debasis Samanta. Hidden features identification for designing an efficient research article recommendation system. International Journal on Digital Libraries, 22(2):233–249, June 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00301-2.

Chatzopoulos:2022:VIE

[CVDT22] Serafeim Chatzopoulos, Thanasis Vergoulis, Theodore Dalamagas, and Christos Tryfonopoulos. VeTo+: improved expert set expansion in academia. International Journal on Digital Libraries, 23(1):57–75, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-021-00318-7.

Catalano:2020:RQD

[CVHS20] Chiara Eva Catalano, Valentina Vassallo, Sorin Hermon, and Michela Spagnuolo. Representing quantitative documentation of 3D cultural heritage artefacts with CIDOC CRMdig. International Journal on Digital Libraries, 21(4):359–373, December 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00287-3.

Coppens:2015:PO

[CVP^{+}15] Sam Coppens, Ruben Verborgh, Sébastien Peyrard, Kevin Ford, Tom Creighton, Rebecca Guenther, Erik Mannens, and Rik Van de Walle. PREMIS OWL. International Journal on Digital Libraries, 15 (2-4):87–101, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0136-9.

Djeraba:2000:VTC

[DBBK00] Chabane Djeraba, Marinette Bouet, Henri Briand, and Ali Khenchaf. Visual and textual content based indexing and retrieval. International Journal on Digital Libraries, 2(4):269–287, May 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/PL00021471.

Debruyne:2016:SAP

[DBG^{+}16] Christophe Debruyne, Oya Deniz Beyan, Rebecca Grant, Sandra Collins, Stefan Decker, and Natalie Harrower. A semantic architecture for preserving and interpreting the information contained in Irish historical vital records. International Journal on Digital Libraries, 17(3):159–174, September 2016. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0180-8.

Downie:2020:HDL

[DBG^{+}20] J. Stephen Downie, Sayan Bhattacharyya, Francesca Giannetti, Eleanor Dickson Koehl, and Peter Organisciak. The HathiTrust Digital Library’s potential for musicology research. International Journal on Digital Libraries, 21(4):343–358, December 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00283-7.

deBoer:2016:EUT

[dBOS16] Victor de Boer, Roeland J. F. Ordelman, and Josefien Schuurman. Evaluating unsupervised thesaurus-based labeling of audiovisual content in an archive production environment. International Journal on Digital Libraries, 17(3):189–201, September 2016. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0182-6.

Darch:2015:WLB

[DBT^{+}15] Peter T. Darch, Christine L. Borgman, Sharon Traweek, Rebekah L. Cummings, Jillian C. Wallis, and Ashley E. Sands. What lies beneath?: Knowledge infrastructures in the subseafloor biosphere and beyond. International Journal on Digital Libraries, 16(1):61–77, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0137-3.

Damm:2012:DLF

[DFT^{+}12] David Damm, Christian Fremerey, Verena Thomas, Michael Clausen, Frank Kurth, and Meinard Müller. A digital library framework for heterogeneous music collections: from document acquisition to cross-modal interaction. International Journal on Digital Libraries, 12(2-3):53–71, August 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0087-y.

Devaney:2019:EMP

[DG19] Johanna Devaney and Hubert Léveillé Gauvin. Encoding music performance data in Humdrum and MEI. International Journal on Digital Libraries, 20(1):81–91, March 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-017-0229-3.

DiNunzio:2023:FID

[Di 23] Giorgio Maria Di Nunzio. Focused issue on digital library challenges to support the open science process. International Journal on Digital Libraries, 24(4): 185–189, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00388-9.

Dekhtyar:2006:SXM

[DIJ^{+}06] Alex Dekhtyar, Ionut E. Iacob, Jerzy W. Jaromczyk, Kevin Kiernan, Neil Moore, and Dorothy Carr Porter. Support for XML markup of image-based electronic editions. International Journal on Digital Libraries, 6 (1):55–69, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0123-2.

Dillon:1999:TML

[Dil99] Andrew Dillon. TIME — a multi-leveled framework for evaluating and designing digital libraries. International Journal on Digital Libraries, 2(2-3):170–177, September 1999. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050045.

delaChica:2008:CFP

[dlCAS^{+}08] Sebastian de la Chica, Faisal Ahmad, Tamara Sumner, James H. Martin, and Kirsten Butcher. Computational foundations for personalizing instruction with digital libraries. International Journal on Digital Libraries, 9(1):3–18, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0037-x.

Desharnais:2002:EAS

[DLR02] P. Desharnais, J. Lu, and T. Radhakrishnan. Exploring agent support at the user interface in e-commerce applications. International Journal on Digital Libraries, 3 (4):284–290, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990100040.

Delcambre:1997:SMM

[DMRA97] Lois M. L. Delcambre, David Maier, Radhika Reddy, and Lougie Anderson. Structured maps: modeling explicit semantics over a universe of information. International Journal on Digital Libraries, 1(1):20–35, April 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050002.

deOliveira:1999:FDI

[dOGM99] Juliano Lopes de Oliveira, Marcos André Gonçalves, and Claudia Bauzer Medeiros. A framework for designing and implementing the user interface of a geographic digital library. International Journal on Digital Libraries, 2(2-3):190–206, September 1999. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050047.

Davidson:1997:BDL

[DOTW97] S. B. Davidson, C. Overton, V. Tannen, and L. Wong. BioKleisli: a digital library for biomedical researchers. International Journal on Digital Libraries, 1(1):36–53, April 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050003.

deOliveira:2023:EMI

[dOVA^{+}23] Lucas Lima de Oliveira, Danny Suarez Vargas, Antônio Marcelo Azevedo Alexandre, Fábio Corrêa Cordeiro, Diogo da Silva Magalhães Gomes, Max de Castro Rodrigues, Regis Kruel Romeu, and Viviane Pereira Moreira. Evaluating and mitigating the impact of OCR errors on information retrieval. International Journal on Digital Libraries, 24(1):45–62, March 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00345-6.

deRibaupierre:2018:EDE

[dRF18] Hélène de Ribaupierre and Gilles Falquet. Extracting discourse elements and annotating scientific documents using the SciAnnot-Doc model: a use case in gender documents. International Journal on Digital Libraries, 19(2-3):271–286, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0227-5.

Barbieri:2021:CSC

[dSBG21] Tamires Tessarolli de Souza Barbieri and Rudinei Goularte. Content selection criteria for news multi-video summarization based on human strategies. International Journal on Digital Libraries, 22(1):1–14, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-020-00281-9.

deSiqueira:2020:PAH

[dSCGL20] Gustavo Oliveira de Siqueira, Sérgio Canuto, Marcos André Gonçalves, and Alberto H. F. Laender. A pragmatic approach to hierarchical categorization of research expertise in the presence of scarce information. International Journal on Digital Libraries, 21(1):61–73, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0260-z.

Deora:2006:EBQ

[DSGF06] V. Deora, J. Shao, W. A. Gray, and N. J. Fiddian. Expectation-based quality of service assessment. International Journal on Digital Libraries, 6(3):260–269, June 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-006-0004-3.

daSilva:2019:RDC

[dSRL19] João Rocha da Silva, Cristina Ribeiro, and João Correia Lopes. Ranking Dublin Core descriptor lists from user interactions: a case study with Dublin Core Terms using the Dendro platform. International Journal on Digital Libraries, 20(2):185–204, June 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0238-x.

Torres:2006:DLF

[dSTMGF06] Ricardo da S. Torres, Claudia Bauzer Medeiros, Marcos André Gonçcalves, and Edward A. Fox. A digital library framework for biodiversity information systems. International Journal on Digital Libraries, 6(1):3–17, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-005-0124-1.

Doyle:2009:LTD

[DVP09] Julie Doyle, Herna Viktor, and Eric Paquet. Long-term digital preservation: preserving authenticity and usability of 3-D data. International Journal on Digital Libraries, 10(1):33–47, May 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-009-0051-7.

Daranyi:2012:UWA

[DWD12] Sándor Darányi, Peter Wittek, and Milena Dobreva. Using wavelet analysis for text categorization in digital libraries: a first experiment with Strathprints. International Journal on Digital Libraries, 12(1):3–12, July 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-012-0079-y.

Elekes:2020:TMN

[EESB20] Ábel Elekes, Adrian Englhardt, Martin Schäler, and Klemens Böhm. Toward meaningful notions of similarity in NLP embedding models. International Journal on Digital Libraries, 21(2):109–128, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0237-y.

Elena:2010:HRA

[EKV^{+}10] Torou Elena, Akrivi Katifori, Costas Vassilakis, George Lepouras, and Constantin Halatsis. Historical research in archives: user methodology and supporting tools. International Journal on Digital Libraries, 11 (1):25–36, March 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-010-0062-4.

Edmunds:2017:EID

[ELH^{+}17] Scott C. Edmunds, Peter Li, Christopher I. Hunter, Si Zhe Xiao, Robert L. Davidson, Nicole Nogoy, and Laurie Goodman. Experiences in integrated data and research object publishing using GigaDB. International Journal on Digital Libraries, 18 (2):99–111, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0174-6.

Esposito:1998:AML

[EMS^{+}98] Floriana Esposito, Donato Malerba, Giovanni Semeraro, Nicola Fanizzi, and Stefano Ferilli. Adding machine learning and knowledge intensive techniques to a digital library service. International Journal on Digital Libraries, 2(1):3–19, October 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050033.

Englhardt:2020:ISC

[EWSB20] Adrian Englhardt, Jens Willkomm, Martin Schäler, and Klemens Böhm. Improving semantic change analysis by combining word embeddings and word frequencies. International Journal on Digital Libraries, 21(3):247–264, September 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-019-00271-6.

Fader:2002:OSI

[Fad02] Christina A. Fader. Optimizing sunk investments in e-commerce: a quality assurance challenge for small businesses. International Journal on Digital Libraries, 3 (4):279–283, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100042.

Frantzi:2000:ARM

[FAM00] Katerina Frantzi, Sophia Ananiadou, and Hideki Mima. Automatic recognition of multiword terms:. the C-value/NC-value method. International Journal on Digital Libraries, 3 (2):115–130, August 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007999900023.

Fantozzi:2017:TMA

[FBPC17] Carlo Fantozzi, Federica Bressan, Niccolò Pretto, and Sergio Canazza. Tape music archives: from preservation to access. International Journal on Digital Libraries, 18(3):233–249, September 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0208-8.

Ferreira:2007:IDS

[FBR07] Miguel Ferreira, Ana Alice Baptista, and José Carlos Ramalho. An intelligent decision support system for digital preservation. International Journal on Digital Libraries, 6(4):295–304, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0013-x.

FragomeniPadron:2020:EIL

[FCR20] Marcos Fragomeni Padron, Fernando William Cruz, and Juliana Rocha de Faria Silva. Extending the IFLA Library Reference Model for a Brazilian popular music digital library. International Journal on Digital Libraries, 21(3):289–306, September 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00277-5.

Ferguson:1998:AID

[FD98] Innes A. Ferguson and Edmund H. Durfee. Artificial intelligence in digital libraries: moving from chaos to (more) order. International Journal on Digital Libraries, 2(1):1–2, October 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050032.

Frew:2000:ADL

[FFF^{+}00] J. Frew, M. Freeston, N. Freitas, L. Hill, G. Janée, K. Lovette, R. Nideffer, T. Smith, and Q. Zheng. The Alexandria Digital Library architecture. International Journal on Digital Libraries, 2(4):259–268, May 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/PL00021470.

Fafalios:2020:BQS

[FHKN20] Pavlos Fafalios, Helge Holzmann, Vaibhav Kasturia, and Wolfgang Nejdl. Building and querying semantic layers for web archives (extended version). International Journal on Digital Libraries, 21(2):149–167, June 2020. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0251-0.

Figuera:2021:FAE

[Fig21] Marianna Figuera. A fuzzy approach to evaluate the attributions reliability in the archaeological sources. International Journal on Digital Libraries, 22(3):289–296, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-020-00284-6.

Fisichella:2021:UAR

[Fis21] Marco Fisichella. Unified approach to retrospective event detection for event-based epidemic intelligence. International Journal on Digital Libraries, 22(4):339–364, December 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-021-00308-9.

Fafalios:2020:THE

[FISN20] Pavlos Fafalios, Vasileios Iosifidis, Kostas Stefanidis, and Eirini Ntoutsi. Tracking the history and evolution of entities: entitycentric temporal analysis of large social media archives. International Journal on Digital Libraries, 21(1):5–17, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0257-7.

Farber:2020:CRA

[FJ20] Michael Färber and Adam Jatowt. Citation recommendation: approaches and datasets. International Journal on Digital Libraries, 21(4):375–405, December 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00288-2.

Fuhr:2019:IFI

[FKRN19] Norbert Fuhr, László Kovács, Thomas Risse, and Wolfgang Nejdl. Introduction to the focused issue on the 20th International Conference on Theory and Practice of Digital Libraries (TPDL 2016). International Journal on Digital Libraries, 20 (2):123–124, June 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-019-00265-4.

Fox:2018:GEI

[FKX18] Edward A. Fox, Martin Klein, and Zhiwu Xie. Guest editors’ introduction to the special issue on web archiving. International Journal on Digital Libraries, 19(1):1–2, March 2018. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0203-5.

Farag:2018:FCE

[FLF18] Mohamed M. G. Farag, Sunshin Lee, and Edward A. Fox. Focused crawler for events. International Journal on Digital Libraries, 19(1):3–19, March 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0207-1.

Felicetti:2017:SMC

[FM17] Achille Felicetti and Francesca Murano. Scripta manent: a CIDOC CRM semiotic reading of ancient texts. International Journal on Digital Libraries, 18(4): 263–270, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0189-z.

Fernando:2018:ACE

[FMN18] Zeon Trevor Fernando, Ivana Marenzi, and Wolfgang Nejdl. ArchiveWeb: collaboratively extending and exploring web archive collections — how would you like to work with your collections? International Journal on Digital Libraries, 19(1):39–55, March 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0206-2.

Fuchs:2004:DLK

[FMNH04] M. Fuchs, C. Muscogiuri, C. Niederée, and M. Hemmje. Digital libraries in knowledge management: an e-learning case study. International Journal on Digital Libraries, 4(1):31–35, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0060-x.

Ferilli:2023:HGB

[FRD23] Stefano Ferilli, Domenico Redavid, and Davide Di Pierro. Holistic graph-based document representation and management for open science. International Journal on Digital Libraries, 24(4):205–227, December 2023. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-022-00328-
z.

Freire:2014:WOB

[Fre14] Nuno Freire. Word occurrence based extraction of work contributors from statements of responsibility. International Journal on Digital Libraries, 14(3-4):141–148, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-014-0113-3.

Freire:2020:CHM

[FRH^{+}20] Nuno Freire, Glen Robson, John B. Howard, Hugo Manguinhas, and Antoine Isaac. Cultural heritage metadata aggregation using web technologies: IIIF, Sitemaps and Schema.org. International Journal on Digital Libraries, 21(1):19–30, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0259-5.

Frommholz:2010:GEI

[FRS10] Ingo Frommholz, Andreas Rauber, and Fabrizio Sebastiani. Guest Editors’ introduction to the focussed issue on the 14th European Conference on Digital Libraries (ECDL 2010). International Journal on Digital Libraries, 11(2):57–58, June 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0068-6.

Foscarin:2021:DQA

[FRT21] Francesco Foscarin, Philippe Rigaux, and Virginie Thion. Data quality assessment in digital score libraries. International Journal on Digital Libraries, 22(2):159–173, June 2021. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-021-00299-
                                        7.

Faulstich:2000:BHW

[FS00] Lukas C. Faulstich and Myra Spiliopoulou. Building HyperView wrappers for publisher Web sites. International Journal on Digital Libraries, 3(1):3–18, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/PL00021472.

Fast:2010:IEP

[FS10] Karl V. Fast and Kamran Sedig. Interaction and the epistemic potential of digital libraries. International Journal on Digital Libraries, 11(3):169–207, September 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-011-0066-8.

Fuhr:2007:EDL

[FTA^{+}07] Norbert Fuhr, Giannis Tsakonas, Trond Aalberg, Maristella Agosti, Preben Hansen, Sarantos Kapidakis, Claus-Peter Klas, László Kovács, Monica Landoni, András Micsik, Christos Papatheodorou, Carol Peters, and Ingeborg Sølvberg. Evaluation of digital libraries. International Journal on Digital Libraries, 8(1):21–38, November 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0011-z.

Furuta:2019:IFI

[FW19] Richard Furuta and Michele C. Weigle. Introduction to the focused issue on the 2016 ACM/IEEE-CS Joint Conference on Digital Libraries JCDL 2016. International Journal on Digital Libraries, 20(4):285, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-019-00276-1.

Gahegan:2007:BRS

[GABD07] Mark Gahegan, Ritesh Agrawal, Tawan Banchuen, and David DiBiase. Building rich, semantic descriptions of learning activities to facilitate reuse in digital libraries. International Journal on Digital Libraries, 7 (1-2):81–97, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0021-x.

Guillem:2017:PCT

[GBR17] Anais Guillem, George Bruseker, and Paola Ronzino. Process, concept or thing? Some initial considerations in the ontological modelling of architecture. International Journal on Digital Libraries, 18(4):289–299, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-016-0188-0.

Ghosh:2000:IEA

[GCAN00] Hiranmay Ghosh, Santanu Chaudhury, Chetan Arora, and Paramjeet Nirankari. ImAge: an extensible agent-based architecture for image retrieval. International Journal on Digital Libraries, 2(4):236–250, May 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/PL00021468.

Gibson:2024:UVI

[GCC24] Ryan Colin Gibson, Sudatta Chowdhury, and Gobinda Chowdhury. User versus institutional perspectives of metadata and searching: an investigation of online access to cultural heritage content during the COVID-19 pandemic. International Journal on Digital Libraries, 25(1):105–121, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00385-y.

Goncalves:2008:TDL

[GFW08] Marcos André Gonçalves, Edward A. Fox, and Layne T. Watson. Towards a digital library theory: a formal digital library ontology. International Journal on Digital Libraries, 8(2):91–114, April 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0033-1.

Grossner:2023:BPD

[GGM23] Karl Grossner, Susan Grunewald, and Ruth Mostern. Bringing places from the distant past to the present: a report on the World Historical Gazetteer. International Journal on Digital Libraries, 24(3): 159–162, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00341-2.

Gerken:2009:LLD

[GHJ^{+}09] Jens Gerken, Mathias Heilig, Hans-Christian Jetter, Sebastian Rexhausen, Mischa Demarmels, Werner A. König, and Harald Reiterer. Lessons learned from the design and evaluation of visual information-seeking systems. International Journal on Digital Libraries, 10(2-3):49–66, August 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-009-0052-6.

Giunchiglia:2012:CMM

[GMA12] Fausto Giunchiglia, Vincenzo Maltese, and Aliaksandr Autayeu. Computing minimal mappings between lightweight ontologies. International Journal on Digital Libraries, 12 (4):179–193, September 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0083-2.

Gotz:2005:FSD

[GMP05] David Gotz and Ketan Mayer-Patel. A framework for scalable delivery of digitized spaces. International Journal on Digital Libraries, 5(3):205–218, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0106-8.

Gudelj:2023:AMT

[GMP^{+}23] Danijel Gudelj, Mirta Matošić, Ljiljana Poljak, Vicko Tomić, Ana Marušić, and Matko Marušić. Accessibility of master’s theses at SEA–EU Alliance universities in open access repositories. International Journal on Digital Libraries, 24(4): 229–241, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00340-3.

Gatos:2000:APA

[GMPT00] B. Gatos, S. L. Mantzaris, S. J. Perantonis, and A. Tsigris. Automatic page analysis for the creation of a digital library from newspaper archives. International Journal on Digital Libraries, 3(1):77–84, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/PL00021477.

Geel:2016:UFH

[GN16] Matthias Geel and Moira C. Norrie. Using a file history graph to keep track of personal resources across devices and services. International Journal on Digital Libraries, 17(3):175–187, September 2016. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0181-7.

Griffin:2005:TNG

[GPT05] Stephen Griffin, Carol Peters, and Costantino Thanos. Towards the new-generation digital libraries: recommendations of the NSF/ EU-DELOS working groups. International Journal on Digital Libraries, 5(4):253–254, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-004-0093-9.

Goldman:2005:ASW

[GRB^{+}05] Jerry Goldman, Steve Renals, Steven Bird, Franciska de Jong, Marcello Federico, Carl Fleischhauer, Mark Kornbluh, Lori Lamel, Douglas W. Oard, Claire Stewart, and Richard Wright. Accessing the spoken word. International Journal on Digital Libraries, 5 (4):287–298, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0101-0.

Gossen:2020:TEE

[GRD20] Gerhard Gossen, Thomas Risse, and Elena Demidova. Towards extracting event-centric collections from Web archives. International Journal on Digital Libraries, 21 (1):31–45, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0258-6.

Greene:1997:MUM

[Gre97] Stephan Greene. Metadata for units of measure in social science databases. International Journal on Digital Libraries, 1(2): 161–175, September 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050012.

Griffin:2015:ISI

[Gri15] Stephen Griffin. Introduction to the special issue on digital scholarship. International Journal on Digital Libraries, 16(1):1– 2, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0142-6.

Griffin:2023:DLE

[Gri23a] Stephen M. Griffin. Digital libraries, epigraphy and paleography: Bring records from the distant past to the present: Part II. International Journal on Digital Libraries, 24(3):139–147, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00381-2.

Griffin:2023:SIE

[Gri23b] Stephen M. Griffin. Special issue: Epigraphy and paleography: Bringing records from the distant past to the present. International Journal on Digital Libraries, 24(2):77–85, June 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00371-4.

Golub:2019:RAK

[GST19] Koraljka Golub, Rudi Schmiede, and Douglas Tudhope. Recent applications of Knowledge Organization Systems: introduction to a special issue. International Journal on Digital Libraries, 20(3):205–207, September 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0264-8.

Gooding:2017:ILC

[GT17] Paul Gooding and Melissa Terras. Inheriting library cards to Babel and Alexandria: contemporary metaphors for the digital library. International Journal on Digital Libraries, 18(3):207–222, September 2017. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-016-0194-2.

Greene:1999:EZH

[GTP^{+}99] Stephan Greene, Egemen Tanin, Catherine Plaisant, Ben Shneiderman, Lola Olsen, Gene Major, and Steve Johns. The end of zerohit queries: query previews for NASA’s Global Change Master Directory. International Journal on Digital Libraries, 2(2-3):79–90, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050039.

Grierson:2008:ESD

[GWB^{+}08] Hilary Grierson, Andrew Wodehouse, Caroline Breslin, William Ion, David Nicol, and Neal Juster. An evaluation study of a digital library of ideas: Workflow Model and classroom use. International Journal on Digital Libraries, 9(1):29–39, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-008-0038-9.

Hasselbring:1997:FIR

[Has97] Wilhelm Hasselbring. Federated integration of replicated information within hospitals. International Journal on Digital Libraries, 1(3):192–208, December 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050016.

Hinze:2016:LTM

[HB16] Annika Hinze and David Bainbridge. Locationtriggered mobile access to a digital library of audio books using Tipple. International Journal on Digital Libraries, 17(4):339–365, November 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0165-z.

Hinze:2019:CLC

[HBC^{+}19] Annika Hinze, David Bainbridge, Sally Jo Cunningham, Craig Taube-Schock, Rangi Matamua, J. Stephen Downie, and Edie Rasmussen. Capisco: low-cost conceptbased access to digital libraries. International Journal on Digital Libraries, 20(4):

307–334, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0232-3.

Harb:2006:ABD

[HC06a] Hadi Harb and Liming Chen. Audio-based description and structuring of videos. International Journal on Digital Libraries, 6 (1):70–81, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0120-5.

Hunter:2006:PIA

[HC06b] Jane Hunter and Sharmin Choudhury. PANIC: an integrated approach to the preservation of composite digital objects using Semantic Web services. International Journal on Digital Libraries, 6(2):174–183, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-005-0134-z.

Hunter:2007:PEG

[HC07] Jane Hunter and Kwok Cheung. Provenance Explorer — a graphical interface for constructing scientific publication packages from provenance trails. International Journal on Digital Libraries, 7(1-2):99–107, October 2007. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-007-0018-5.

Hong:2004:TIS

[HCH04] Jen-Shin Hong, Bai-Hsuan Chen, and Sheng-Hao Hung. Toward intelligent styling for digital museum exhibitions: modularization framework for aesthetic hypermedia presentations. International Journal on Digital Libraries, 4(1):64–68, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0059-3.

Hong:2005:TID

[HCHH05] Jen-Shin Hong, Bai-Hsuan Chen, Sheng-Hao Hung, and Jieh Hsiang. Toward an integrated digital museum system — the Chi Nan experiences. International Journal on Digital Libraries, 5(3):231–251, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0116-1.

Hiebel:2017:CSE

[HDE17] Gerald Hiebel, Martin Doerr, and Øyvind Eide. CRMgeo: a spatiotemporal extension of CIDOC–CRM. International Journal on Digital Libraries, 18(4):271–279, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-016-0192-4.

Hamborg:2019:AIM

[HDG19] Felix Hamborg, Karsten Donnay, and Bela Gipp. Automated identification of media bias in news articles: an interdisciplinary literature review. International Journal on Digital Libraries, 20(4):391–415, December 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0261-y.

Hong:2005:E

[HDH05] Jen-Shin Hong, Martin Doerr, and Jieh Hsiang. Editorial. International Journal on Digital Libraries, 5(3):151–152, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-005-0115-2.

Hiebel:2021:FDP

[HGG^{+}21] Gerald Hiebel, Gert Goldenberg, Caroline Grutsch, Klaus Hanke, and Markus Staudt. FAIR data for prehistoric mining archaeology. International Journal on Digital Libraries, 22(3):267–277, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00282-8.

Hiemstra:2000:PJU

[Hie00] Djoerd Hiemstra. A probabilistic justification for using tf × idf term weighting in information retrieval. International Journal on Digital Libraries, 3(2):131–139, August 2000. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007999900025.

Haslhofer:2009:LAF

[HJK^{+}09] Bernhard Haslhofer, Wolfgang Jochum, Ross King, Christian Sadilek, and Karin Schellner. The LEMO annotation framework: weaving multimedia annotations with the Web. International Journal on Digital Libraries, 10(1):15–32, May 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-009-0050-8.

Heutelbeck:2012:IFI

[HK12] Dominic Heutelbeck and Claus-Peter Klas. IJDL focussed issue on persistent archives. International Journal on Digital Libraries, 12 (1):1, July 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-012-0093-0.

Huurdeman:2015:LFF

[HKS^{+}15] Hugo C. Huurdeman, Jaap Kamps, Thaer Samar, Arjen P. de Vries, Anat Ben-David, and Richard A. Rogers. Lost but not forgotten: finding pages on the unarchived web. International Journal on Digital Libraries, 16(3-4):247–265, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0153-3.

Hunt:2005:IVA

[HLZ05] Leta Hunt, Marilyn Lundberg, and Bruce Zuckerman. InscriptiFact: a virtual archive of ancient inscriptions from the Near East. International Journal on Digital Libraries, 5(3):153–166, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0102-z.

Hamborg:2020:BAN

[HMG20] Felix Hamborg, Norman Meuschke, and Bela Gipp. Bias-aware news analysis using matrix-based news aggregation. International Journal on Digital Libraries, 21 (2):129–147, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0239-9.

Hiary:2007:SSE

[HN07] Hazem Hiary and Kia Ng. A system for segmenting and extracting paper-based watermark designs. International Journal on Digital Libraries, 6(4):351–361, July 2007. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-007-0008-7.

Hermon:2021:FDC

[HN21] Sorin Hermon and Franco Niccolucci. FAIR Data and Cultural Heritage special issue editorial note. International Journal on Digital Libraries, 22(3):251–255, September 2021. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-021-00309-
                                        8.

Hlaoua:2010:RFR

[HPSB10] Lobna Hlaoua, Karen Pinel-Sauvagnat, and Mohand Boughanem. Relevance feedback revisited: dealing with content and structure in XML documents. International Journal on Digital Libraries, 11(1):1–24, March 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-010-0061-5.

Henry:1997:RAU

[HR97] Geneva Henry and Jim Reimer. A radiology archive using a digital library. International Journal on Digital Libraries, 1 (3):231–240, December 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050019.

Hartson:2004:UID

[HSPQ04] H. Rex Hartson, Priya Shivakumar, and Manuel A. Pérez-Quiñones. Usability inspection of digital libraries: a case study. International Journal on Digital Libraries, 4 (2):108–123, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0074-4.

Holzmann:2015:NEE

[HTR15] Helge Holzmann, Nina Tahmasebi, and Thomas Risse. Named entity evolution recognition on the Blogosphere. International Journal on Digital Libraries, 15 (2-4):209–235, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0135-x.

Husevaag:2019:SSM

[Hus19] Anne-Stine Ruud Husevåg. From subtitles to substantial metadata: examining characteristics of named entities and their role in indexing. International Journal on Digital Libraries, 20(3):241–251, September 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0252-z.

Harrison:2017:IWP

[HWT17] Kathryn A. Harrison, Daniel G. Wright, and Philip Trembath. Implementation of a workflow for publishing citeable environmental data: successes, challenges and opportunities from a data centre perspective. International Journal on Digital Libraries, 18(2):133–143, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0175-5.

Ioannidis:2005:DLI

[IMA^{+}05] Yannis Ioannidis, David Maier, Serge Abiteboul, Peter Buneman, Susan Davidson, Edward Fox, Alon Halevy, Craig Knoblock, Fausto Rabitti, Hans Schek, and Gerhard Weikum. Digital library informationtechnology infrastructures. International Journal on Digital Libraries, 5(4):266–274, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-004-0094-8.

Inskip:2012:TDC

[IMR12] Charles Inskip, Andy Macfarlane, and Pauline Rafferty. Towards the disintermediation of creative music search: analysing queries to determine important facets. International Journal on Digital Libraries, 12(2-3):137–147, August 2012. CODEN ????

ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-012-0084-1.

Ioannidis:2008:D

[IMSS08] Yannis Ioannidis, Diego Milano, Hans-Jörg Schek, and Heiko Schuldt. DelosDLMS. International Journal on Digital Libraries, 9 (2):101–114, November 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0044-y.

Ioannidis:2005:DLC

[Ioa05] Yannis Ioannidis. Digital libraries at a crossroads. International Journal on Digital Libraries, 5(4):255–265, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0098-4.

Ishita:2024:SIS

[ISK^{+}24] Emi Ishita, Sue Yeon Syn, Hao-Ren Ke, Chei Sian Lee, and Kazunari Sugiyama. Special issue on selected papers from ICADL 2021. International Journal on Digital Libraries, 25(1):1–2, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-024-00399-0.

Ishida:2020:ACK

[ISY20] Youichi Ishida, Toshiyuki Shimizu, and Masatoshi Yoshikawa. An analysis and comparison of keyword recommendation methods for scientific data. International Journal on Digital Libraries, 21(3):307–327, September 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-020-00279-3.

Itao:2006:SLS

[ITSY06] Tomoko Itao, Satoshi Tanaka, Tatsuya Suda, and Atsushi Yamamoto. SpaceGlue: linking spaces for adaptive service location. International Journal on Digital Libraries, 6 (3):270–279, June 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-006-0005-2.

Jurik:2015:BGB

[JBFNMD15] Bolette Ammitzbøll Jurik, Asger Askov Blekinge, Rune Bruun Ferneke-Nielsen, and Per Møldrup-Dalum. Bridging the gap between real world repositories and scalable preservation environments. International Journal on Digital Libraries, 16(3-4): 267–282, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0152-4.

Jones:2000:TLA

[JCMB00] Steve Jones, Sally Jo Cunningham, Rodger McNab, and Stefan Boddie. A transaction log analysis of a digital library. International Journal on Digital Libraries, 3 (2):152–169, August 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007999900022.

Jaidka:2018:ICS

[JCRK18] Kokil Jaidka, Muthu Kumar Chandrasekaran, Sajal Rustagi, and Min-Yen Kan. Insights from CL–SciSumm 2016: the faceted scientific document summarization Shared Task. International Journal on Digital Libraries, 19(2-3):163–171, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0221-y.

Chen:2016:ACS

[jCZhC16] Shu jiun Chen, Marcia Lei Zeng, and Hsueh hua Chen. Alignment of conceptual structures in controlled vocabularies in the domain of Chinese art: a discussion of issues and patterns. International Journal on Digital Libraries, 17(1):23–38, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0163-1.

Jiang:2022:EBB

[JDAD22] Ming Jiang, Jennifer D’Souza, Sören Auer, and J. Stephen Downie. Evaluating BERTbased scientific relation classifiers for scholarly knowledge graph construction on digital library collections. International Journal on Digital Libraries, 23(2):197–215, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-021-00313-y.

Johnston:1997:RTD

[JGL^{+}97] William Johnston, Jin Guojun, Case Larsen, Jason Lee, Gary Hoo, Mary Thompson, Brian Tierney, and Joseph Terdiman. Real-time digital libraries based on widely distributed, high performance management of large data objects. International Journal on Digital Libraries, 1(3):241–256, December 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050020.

Jatowt:2024:SIS

[JKPC24] Adam Jatowt, Marie Katsurai, Muhammad Syafiq Mohd Pozi, and Ricardo Campos. Special issue on selected papers from ICADL 2022. International Journal on Digital Libraries, 25(1):73–74, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-024-00400-w.

Jatowt:2019:TFA

[JKT19] Adam Jatowt, Daisuke Kawai, and Katsumi Tanaka. Time-focused analysis of connectivity and popularity of historical persons in Wikipedia. International Journal on Digital Libraries, 20(4):287–305, December 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0231-4.

Jatowt:2022:SIS

[JMI^{+}22] Adam Jatowt, Atsuyuki Morishima, Emi Ishita, Natalie Pang, and Lihong Zhou. Special issue on selected papers from ICADL 2020. International Journal on Digital Libraries, 23(2):111–112, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00323-4.

Jones:1999:GUI

[JMS99] Steve Jones, Shona McInnes, and Mark S. Staveley. A graphical user interface for Boolean query specification. International Journal on Digital Libraries, 2(2-3):207–223, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990050048.

Jones:2018:ASW

[JNV18] Shawn M. Jones, Michael L. Nelson, and Herbert Van de Sompel. Avoiding spoilers: wiki time travel with Sheldon Cooper. International Journal on Digital Libraries, 19(1):77–93, March 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0200-8.

Janecek:2005:ESF

[JP05] Paul Janecek and Pearl Pu. An evaluation of semantic fisheye views for opportunistic search in an annotated image collection. International Journal on Digital Libraries, 5(1):42–56, March 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0108-6.

Johnson:1997:EIF

[JTCA97] David B. Johnson, Ricky K. Taira, Alfonso F. Cardenas, and Denise R. Aberle. Extracting information from free text radiology reports. International Journal on Digital Libraries, 1(3):297–308, December 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050024.

Kokuer:2019:CAC

[KAMKJ19] Münevver Köküer, Islah Ali-MacLachlan, Daithí Kearney, and Peter Jančovič. Curating and annotating a collection of traditional Irish flute recordings to facilitate stylistic analysis. International Journal on Digital Libraries, 20(1):107–121, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-019-00267-2.

Klein:2022:EAP

[KB22] Martin Klein and Lyudmila Balakireva. An extended analysis of the persistence of persistent identifiers of the scholarly web. International Journal on Digital Libraries, 23(1):5–17, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-021-00315-w.

Klein:2019:CPS

[KBFG19] Martin Klein, Peter Broadwell, Sharon E. Farb, and Todd Grappone. Comparing published scientific journal articles to their pre-print versions. International Journal on Digital Libraries, 20(4):335–350, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-018-0234-1.

Klose:2023:CLC

[KBG23] Annamarie C. Klose, Isaiah Beard, and Scott Goldstein. Coins in the library: the creation of a digital collection of Roman Republican coins. International Journal on Digital Libraries, 24(2):103–115, June 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00338-x.

Kortemeyer:2021:UTB

[KD21] Gerd Kortemeyer and Stefan Dröschler. A user-transaction-based recommendation strategy for an educational digital library. International Journal on Digital Libraries, 22 (2):147–157, June 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00298-8.

Kabongo:2024:OLS

[KDA24] Salomon Kabongo, Jennifer D’Souza, and Sören Auer. ORKG-Leaderboards: a systematic workflow for mining leaderboards as a knowledge graph. International Journal on Digital Libraries, 25(1):41–54, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-023-00366-1.

Kortemeyer:2014:HLU

[KDP14] Gerd Kortemeyer, Stefan Dröschler, and David E. Pritchard. Harvesting latent and usage-based metadata in a course management system to enrich the underlying educational digital library. International Journal on Digital Libraries, 14(1-2):1–15, April 2014. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-013-0107-6.

Ke:2015:ITT

[Ke15] Weimao Ke. Information-theoretic term weighting schemes for document clustering and classification. International Journal on Digital Libraries, 16(2):145–159, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-014-0121-3.

Kelly:2024:EUF

[Kel24] Mat Kelly. Exploiting the untapped functional potential of Memento aggregators beyond aggregation. International Journal on Digital Libraries, 25(1):93–104, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00391-0.

Klein:1998:ETD

[KF98] Bertin Klein and Peter Fankhauser. Error tolerant document structure analysis. International Journal on Digital Libraries, 1 (4):344–357, March 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050028.

Karadkar:2004:MSD

[KFRF^{+}04] Unmil P. Karadkar, Luis Francisco-Revilla, Richard Furuta, Frank Shipman, Avital Arora, Suvendu Dash, Pratik Dave, and Emily Luke. Metadocuments supporting digital library information discovery. International Journal on Digital Libraries, 4 (1):25–30, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0064-6.

Kodali:2004:AMM

[KFW04] Naren Kodali, Csilla Farkas, and Duminda Wijesekera. An authorization model for multimedia digital libraries. International Journal on Digital Libraries, 4(3):139–155, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-004-0080-1.

Kaur:2021:FBF

[KG21] Satinder Kaur and Sunil Kumar Gupta. A fuzzy-based framework for evaluation of website design quality index. International Journal on Digital Libraries, 22 (1):15–47, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00292-6.

Kumar:2023:EDU

[KGE23] Sandeep Kumar, Tirthankar Ghosal, and Asif Ekbal. DeepMetaGen: an unsupervised deep neural approach to generate template-based meta-reviews leveraging on aspect category and sentiment analysis from peer reviews. International Journal on Digital Libraries, 24(4):263–281, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00348-3.

Klampfl:2014:UDS

[KGJK14] Stefan Klampfl, Michael Granitzer, Kris Jack, and Roman Kern. Unsupervised document structure analysis of digital scientific articles. International Journal on Digital Libraries, 14(3-4):83–99, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0115-1.

Khazaei:2017:SAS

[KH17] Taraneh Khazaei and Orland Hoeber. Supporting academic search tasks through citation visualization and exploration. International Journal on Digital Libraries, 18 (1):59–72, March 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0170-x.

Kumara:2023:ORL

[KHVT23] P. P. N. V. Kumara, Annika Hinze, Nicholas Vanderschantz, and Claire Timpany. Online reading lists: a mixed-method analysis of the academic perspective. International Journal on Digital Libraries, 24 (1):23–44, March 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00344-z.

Koutrika:2004:RBQ

[KI04] Georgia Koutrika and Yannis Ioannidis. Rulebased query personalization in digital libraries. International Journal on Digital Libraries, 4(1):60–63, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0068-2.

Kamdar:2005:UIW

[KJ05] Tapan Kamdar and Anupam Joshi. Using incremental Web log mining to create adaptive web servers. International Journal on Digital Libraries, 5(2):133–150, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0057-5.

Kakimoto:1999:BFT

[KK99] Toshihiro Kakimoto and Yahiko Kambayashi. Browsing functions in three-dimensional space for digital libraries. International Journal on Digital Libraries, 2(2-3):68–78, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990050038.

Kan:2008:SDL

[KLL08] Min-Yen Kan, Dongwon Lee, and Ee-Peng Lim. Scholarly digital libraries at scale: introduction to the special issue on very large digital libraries. International Journal on Digital Libraries, 9(2):81–82, November 2008. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-008-0042-0.

Kostkova:2013:WID

[KM13] Patty Kostkova and Gemma Madle. What impact do healthcare digital libraries have? An evaluation of national resource of infection control at the point of care using the Impact– ED framework. International Journal on Digital Libraries, 13(2):77–90, March 2013. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-012-0101-4.

Karimi:2019:ANS

[KMM19] Fariba Karimi, Philipp Mayr, and Fakhri Momeni. Analyzing the network structure and gender differences among the members of the Networked Knowledge Organization Systems (NKOS) community. International Journal on Digital Libraries, 20(3): 231–239, September 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0243-0.

Kapidakis:2016:ETS

[KMW16] Sarantos Kapidakis, Cezary Mazurek, and Marcin Werla. Editorial for the TPDL 2015 special issue. International Journal on Digital Libraries, 17(3):157–158, September 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-016-0190-6.

Klein:2014:MGE

[KN14] Martin Klein and Michael L. Nelson. Moved but not gone: an evaluation of real-time methods for discovering replacement web pages. International Journal on Digital Libraries, 14(1-2):17–38, April 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0108-0.

Kramer:1997:TFL

[KNH97] Ralf Kramer, Ralf Nikolai, and Corinna Habeck. Thesaurus federations: loosely integrated thesauri for document retrieval in networks based on Internet technologies. International Journal on Digital Libraries, 1 (2):122–131, September 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050009.

Kortemeyer:2016:SCQ

[Kor16] Gerd Kortemeyer. Scalable continual quality control of formative assessment items in an educational digital library: an empirical study. International Journal on Digital Libraries, 17(2):143–155, June 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0145-3.

Kramer:2021:DSB

[KPC^{+}21] Thomas Krämer, Andrea Papenmeier, Zeljko Carevic, Dagmar Kern, and Brigitte Mathiak. Data-seeking behaviour in the social sciences. International Journal on Digital Libraries, 22(2):175–195, June 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00303-0.

Kroll:2024:DSN

[KPK^{+}24] Hermann Kroll, Jan Pirklbauer, Jan-Christoph Kalo, Morris Kunz, Johannes Ruthmann, and Wolf-Tilo Balke. A discovery system for narrative query graphs: entityinteraction-aware document retrieval. International Journal on Digital Libraries, 25(1): 3–24, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-023-00356-3.

Klein:2015:GEI

[KR15] Martin Klein and Andreas Rauber. Guest editors’ introduction to the special issue on the Digital Libraries Conference 2014. International Journal on Digital Libraries, 16 (3-4):181–182, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0161-3.

Kulmukhametov:2021:IDQ

[KRB21] Artur Kulmukhametov, Andreas Rauber, and Christoph Becker. Improving data quality in large-scale repositories through conflict resolution. International Journal on Digital Libraries, 22(4):365–383, December 2021. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-021-00311-

Koopman:2018:TOS

[KRZ18] Bevan Koopman, Jack Russell, and Guido Zuccon. Task-oriented search for evidencebased medicine. International Journal on Digital Libraries, 19(2-3):217–229, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-017-0209-7.

Kreutz:2022:SPR

[KS22] Christin Katharina Kreutz and Ralf Schenkel. Scientific paper recommendation systems: a literature review of recent publications. International Journal on Digital Libraries, 23 (4):335–369, December 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00339-w.

Koutsomitropoulos:2009:SWE

[KSAP09] Dimitrios A. Koutsomitropoulos, Georgia D. Solomou, Andreas D. Alexopoulos, and Theodore S. Papatheodorou. Semantic Web enabled digital repositories. International Journal on Digital Libraries, 10(4): 179–199, December 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-010-0059-z.

Kern:2010:HSG

[KSG10] Roman Kern, Christin Seifert, and Michael Granitzer. A hybrid system for German encyclopedia alignment. International Journal on Digital Libraries, 11(2):75–89, June 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-011-0069-5.

Kengeri:1999:USD

[KSH^{+}99] Rekha Kengeri, Cheryl D. Seals, Hope D. Harley, Himabindu P. Reddy, and Edward A.

Fox. Usability study of digital libraries: ACM, IEEE–CS, NCSTRL, NDLTD. International Journal on Digital Libraries, 2 (2-3):157–169, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050044.

Kapidakis:2000:FPM

[KTS00] Sarantos Kapidakis, Sotirios Terzis, and Jakka Sairamesh. A framework for performance monitoring, load balancing, adaptive timeouts and quality of service in digital libraries. International Journal on Digital Libraries, 3(1):19–35, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00021473.

Kurakawa:2014:RNR

[KTT^{+}14] Kei Kurakawa, Hideaki Takeda, Masao Takaku, Akiko Aizawa, Ryo Shiozaki, Shun Morimoto, and Hideki Uchijima. Researcher Name Resolver: identifier management system for Japanese researchers. International Journal on Digital Libraries, 14(1-2):39–58, April 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-014-0109-z.

Kitchin:2017:ADS

[KVZ17] John R. Kitchin, Ana E. Van Gulick, and Lisa D. Zilinski. Automating data sharing through authoring tools. International Journal on Digital Libraries, 18(2):93–98, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-016-0173-7.

Kahn:2006:FDD

[KW06] Robert Kahn and Robert Wilensky. A framework for distributed digital object services. International Journal on Digital Libraries, 6(2):115–123, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0128-x.

Keil:2019:ARD

[KW19] Klaus Keil and Jennifer A. Ward. Applications of RISM data in digital libraries and digital musicology. International Journal on Digital Libraries, 20(1):3–12, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0205-3.

Kreutz:2022:SDA

[KWK^{+}22] Christin Katharina Kreutz, Michael Wolz, Jascha Knack, Benjamin Weyers, and Ralf Schenkel. SchenQL: in-depth analysis of a query language for bibliographic metadata. International Journal on Digital Libraries, 23(2):113–132, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-021-00317-8.

Kashyap:2023:SDP

[KYK23] Abhinav Ramesh Kashyap, Yajing Yang, and Min-Yen Kan. Scientific document processing: challenges for modern learning methods. International Journal on Digital Libraries, 24(4):283–309, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00352-7.

Leuski:2000:SBI

[LA00] Anton Leuski and James Allan. Strategybased interactive cluster visualization for information retrieval. International Journal on Digital Libraries, 3(2):170–184, August 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007999900016.

Landon:2015:PDR

[Lan15] George V. Landon. A pipeline for digital restoration of deteriorating photographic negatives. International Journal on Digital Libraries, 16(1):79–88, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0133-z.

Lazovik:2006:PME

[LAP06] Alexander Lazovik, Marco Aiello, and Mike Papazoglou. Planning and monitoring the execution of web service requests. International Journal on Digital Libraries, 6 (3):235–246, June 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-006-0002-5.

Lopuszynski:2015:TRT

[ŁB15] Michał Łopuszyński and Łukasz Bolikowski. Towards robust tags for scientific publications from natural language processing tools and Wikipedia. International Journal on Digital Libraries, 16(1):25–36, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0132-0.

Littman:2018:ABS

[LCK^{+}18] Justin Littman, Daniel Chudnov, Daniel Kerchner, Christie Peterson, Yecheng Tan, Rachel Trent, Rajat Vij, and Laura Wrubel. APIbased social media collecting as a form of web archiving. International Journal on Digital Libraries, 19(1):21–38, March 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0201-7.

Lim:2004:IJD

[LCN^{+}04] Ee-Peng Lim, Hsinchun Chen, Erich J. Neuhold, Shigeo Sugimoto, and Jianzhong Li. Introduction to the [International] Journal on Digital Libraries special issue on Asian digital libraries. International Journal on Digital Libraries, 4(4):245–246, November 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-004-0092-x.

Lee:2007:PME

[Lee07] Dongwon Lee. Practical maintenance of evolving metadata for digital preservation: algorithmic solution and system support. International Journal on Digital Libraries, 6 (4):313–326, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0014-9.

Lesk:2015:NKI

[Les15] Michael Lesk. The new knowledge infrastructure. International Journal on Digital Libraries, 16(1):3–4, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0143-5.

Long:1997:OXR

[LGT97] L. Rodney Long, Gin-Hua Goh, and George R. Thoma. Online X-ray atlas as a reference tool. International Journal on Digital Libraries, 1(3):220–230, December 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050018.

Lin:2008:IIP

[LHD08] Chia-Hung Lin, Jen-Shin Hong, and Martin Doerr. Issues in an inference platform for generating deductive knowledge: a case study in cultural heritage digital libraries using the CIDOC CRM. International Journal on Digital Libraries, 8(2):115–132, April 2008. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-008-0034-0.

Lin:2017:MMD

[LKA^{+}17] Xia Lin, Michael Khoo, Jae-Wook Ahn, Doug Tudhope, Ceri Binding, Diana Massam, and Hilary Jones. Mapping metadata to DDC classification structures for searching and browsing. International Journal on Digital Libraries, 18(1):25–39, March 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0197-z.

Litoiu:2002:PSV

[LKR02] M. Litoiu, D. Krishnamurthy, and J. Rolia. Performance stress vectors and capacity planning for e-commerce applications. International Journal on Digital Libraries, 3(4):309–315, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100045.

Li:2016:LAS

[LLS16] Ting Li, Yuhua Lin, and Haiying Shen. A locality-aware similar information searching scheme. International Journal on Digital Libraries, 17(2):79–93, June 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0128-9.

Li:2008:WAS

[LLSG08] Huajing Li, Wang-Chien Lee, Anand Sivasubramaniam, and C. Lee Giles. Workload analysis for scientific literature digital libraries. International Journal on Digital Libraries, 9(2):139–149, November 2008. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-008-0043-z.

Li:2018:CLL

[LMZ^{+}18] Lei Li, Liyuan Mao, Yazhao Zhang, Junqi Chi, Taiwen Huang, Xiaoyue Cong, and Heng Peng. Computational linguistics literature and citations oriented citation linkage, classification and summarization. International Journal on Digital Libraries, 19 (2-3):173–190, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0219-5.

Lagoze:2006:FAC

[LPSW06] Carl Lagoze, Sandy Payette, Edwin Shin, and Chris Wilper. Fedora: an architecture for complex objects and their relationships. International Journal on Digital Libraries, 6(2):124–138, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0130-3.

Lorenzini:2021:AEQ

[LRT21] Matteo Lorenzini, Marco Rospocher, and Sara Tonelli. Automatically evaluating the quality of textual descriptions in cultural heritage records. International Journal on Digital Libraries, 22(2):217–231, June 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-021-00302-1.

Lanagan:2012:VDL

[LS12] James Lanagan and Alan F. Smeaton. Video digital libraries: contributive and decentralised. International Journal on Digital Libraries, 12(4):159–178, September 2012. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-012-0078-z.

Le:2019:AQA

[LSC19] Long T. Le, Chirag Shah, and Erik Choi. Assessing the quality of answers autonomously in community question–answering. International Journal on Digital Libraries, 20(4): 351–367, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-019-00272-5.

Lee:2013:DVG

[LTCC13] Jin Ha Lee, Joseph T. Tennis, Rachel Ivy Clarke, and Michael Carpenter. Developing a video game metadata schema for the Seattle interactive media museum. International Journal on Digital Libraries, 13 (2):105–117, March 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-013-0103-x.

Liesaputra:2009:CGT

[LW09] Veronica Liesaputra and Ian H. Witten. Computer graphics techniques for modeling page turning. International Journal on Digital Libraries, 10(2-3):93–121, August 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-009-0055-3.

Mayer:2015:UOC

[MAC^{+}15] Rudolf Mayer, Gonçalo Antunes, Artur Caetano, Marzieh Bakhshandeh, Andreas Rauber, and José Borbinha. Using ontologies to capture the semantics of a (business) process for digital preservation. International Journal on Digital Libraries, 15 (2-4):129–152, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0141-7.

Marchionini:2004:BES

[Mar04] Gary Marchionini. A briefing on the evolution and status of the Open Video digital library. International Journal on Digital Libraries, 4(1):36–38, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0058-4.

Metilli:2019:WBT

[MBM19] Daniele Metilli, Valentina Bartalesi, and Carlo Meghini. A Wikidata-based tool for building and visualising narratives. International Journal on Digital Libraries, 20(4): 417–432, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-019-00266-3.

Mi:2021:BAD

[MBS21] Xiying Mi, Richard Bernardy, and LeEtta Schmidt. Building an archaeological data repository: a digital library and digital humanities collaboration at the University of South Florida. International Journal on Digital Libraries, 22(1):135–145, March 2021. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-020-00297-
                                        1.

Moraes:2018:IRS

[MBVL18] Luis Moraes, Shahryar Baki, Rakesh Verma, and Daniel Lee. Identifying reference spans: topic modeling and word embeddings help IR. International Journal on Digital Libraries, 19(2-3):191–202, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0220-z.

McDonough:2006:MSE

[McD06] Jerome P. McDonough. METS: standardized encoding for digital library objects. International Journal on Digital Libraries, 6 (2):148–158, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0132-1.

Mayr:2018:ISI

[MFC^{+}18] Philipp Mayr, Ingo Frommholz, Guillaume Cabanac, Muthu Kumar Chandrasekaran, Kokil Jaidka, Min-Yen Kan, and Dietmar Wolfram. Introduction to the special issue on bibliometric-enhanced information retrieval and natural language processing for digital libraries (BIRNDL). International Journal on Digital Libraries, 19 (2-3):107–111, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0230-x.

Mariani:2018:RPS

[MFP18] Joseph Mariani, Gil Francopoulo, and Patrick Paroubek. Reuse and plagiarism in speech and natural language processing publications. International Journal on Digital Libraries, 19(2-3):113–126, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0211-0.

Mittal:2006:ACB

[MG06] Ankush Mittal and Sumit Gupta. Automatic content-based retrieval and semantic classification of video content. International Journal on Digital Libraries, 6 (1):30–38, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0119-y.

Mardis:2008:NFU

[MHM08] Marcia A. Mardis, Ellen S. Hoffman, and Todd E. Marshall. A new framework for understanding educational digital library use: re-examining digital divides in U.S. schools. International Journal on Digital Libraries, 9(1):19–27, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0035-z.

Marks:2005:ADL

[MHML05] Linn Marks, Jeremy A. T. Hussell, Tamara M. McMahon, and Richard E. Luce. ActiveGraph: a digital library visualization tool. International Journal on Digital Libraries, 5(1):57–69, March 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0110-z.

Momeni:2015:SUC

[MHTH15] Elaheh Momeni, Bernhard Haslhofer, Ke Tao, and Geert-Jan Houben. Sifting useful comments from Flickr Commons and YouTube. International Journal on Digital Libraries, 16(2):161–179, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0123-1.

McMartin:2008:UOD

[MIW^{+}08] Flora McMartin, Ellen Iverson, Alan Wolf, Joshua Morrill, Glenda Morgan, and Cathryn Manduca. The use of online digital resources and educational digital libraries in higher education. International Journal on Digital Libraries, 9(1):65–79, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0036-y.

Munster:2018:ILT

[MKFK18] Sander Münster, Christina Kamposiori, Kristina Friedrichs, and Cindy Kröber. Image libraries and their scholarly use in the field of art and architectural history. International Journal on Digital Libraries, 19(4):367–383, November 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-018-0250-1.

Muhling:2022:VVI

[MKPI^{+}22] Markus Mühling, Nikolaus Korfhage, Kader Pustu-Iren, Joanna Bars, Mario Knapp, Hicham Bellafkir, Markus Vogelbacher, Daniel Schneider, Angelika Hörth, Ralph Ewerth, and Bernd Freisleben. VIVA: visual information retrieval in video archives. International Journal on Digital Libraries, 23 (4):319–333, December 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00337-y.

Maurer:2013:CWT

[MM13] Hermann Maurer and Heimo Mueller. Can the Web turn into a digital library? International Journal on Digital Libraries, 13 (2):65–75, March 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0097-9.

Marshall:2020:IFI

[MMJ20] Catherine C. Marshall, Ian Milligan, and Adam Jatowt. Introduction to the focused issue on the 2017 ACM/IEEE-CS Joint Conference on Digital Libraries JCDL 2017. International Journal on Digital Libraries, 21(2):89–91, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00286-4.

Marketakis:2017:XMF

[MMK^{+}17] Yannis Marketakis, Nikos Minadakis, Haridimos Kondylakis, Konstantina Konsolaki, Georgios Samaritakis, Maria Theodoridou, Giorgos Flouris, and Martin Doerr. X3ML mapping framework for information integration in cultural heritage and beyond. International Journal on Digital Libraries, 18 (4):301–319, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0179-1.

Muhling:2019:CBV

[MMK^{+}19] Markus Mühling, Manja Meister, Nikolaus Korfhage, Jörg Wehling, Angelika Hörth, Ralph Ewerth, and Bernd Freisleben. Contentbased video retrieval in historical collections of the German Broadcasting Archive. International Journal on Digital Libraries, 20(2):167–183, June 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0236-z.

Mendelzon:1997:QWW

[MMM97] Alberto O. Mendelzon, George A. Mihaila, and Tova Milo. Querying the World Wide Web. International Journal on Digital Libraries, 1(1):54–67, April 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050004.

Mandreoli:2004:DCS

[MMT04] Federica Mandreoli, Riccardo Martoglia, and Paolo Tiberio. A document comparison scheme for secure duplicate detection. International Journal on Digital Libraries, 4 (3):223–244, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0079-7.

Malliari:2022:CMA

[MNZD22a] Afrodite Malliari, Ilias Nitsos, Sofia Zapounidou, and Stavros Doropoulos. Correction to: Mapping audiovisual content providers and resources in Greece. International Journal on Digital Libraries, 23 (3):229, September 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00334-1.

Malliari:2022:MAC

[MNZD22b] Afrodite Malliari, Ilias Nitsos, Sofia Zapounidou, and Stavros Doropoulos. Mapping audiovisual content providers and resources in Greece. International Journal on Digital Libraries, 23(3):217–227, September 2022. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-022-00321-
6.

Martin:2002:MDS

[MPLR02] Patrick Martin, Wendy Powley, Hoi-Ying Li, and Keri Romanufa. Managing database server performance to meet QoS requirements in electronic commerce systems. International Journal on Digital Libraries, 3 (4):316–324, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100046.

Marlet:2021:WER

[MR21] Olivier Marlet and Xavier Rodier. A way to express the reliability of archaeological data: data traceability at the Laboratoire Archéologie et Territoires (Tours, France). International Journal on Digital Libraries, 22(3):257–266, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-019-00274-3.

Mundur:2006:KBV

[MRY06] Padmavathi Mundur, Yong Rao, and Yelena Yesha. Keyframe-based video summarization using Delaunay clustering. International Journal on Digital Libraries, 6 (2):219–232, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0129-9.

Munyaradzi:2014:SHQ

[MS14] Ngoni Munyaradzi and Hussein Suleman. A system for high quality crowdsourced indigenous language transcription. International Journal on Digital Libraries, 14(3-4):117–125, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0112-4.

Mutschke:2017:GEI

[MSB^{+}17] Peter Mutschke, Andrea Scharnhorst, Nicholas J. Belkin, André Skupin, and Philipp Mayr. Guest editors’ introduction to the special issue on knowledge maps and information retrieval (KMIR). International Journal on Digital Libraries, 18(1):1–3, March 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0204-4.

McKay:2004:EBD

[MSHC04] Dana McKay, Preeti Shukla, Rachael Hunt, and Sally Jo Cunningham. Enhanced browsing in digital libraries: three new approaches to browsing in Greenstone. International Journal on Digital Libraries, 4 (4):283–297, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0088-6.

Maull:2012:AAC

[MSS12] Keith E. Maull, Manuel Gerardo Saldivar, and Tamara Sumner. Automated approaches to characterizing educational digital library usage: linking computational methods with qualitative analyses. International Journal on Digital Libraries, 13 (1):51–64, December 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0096-x.

Meuss:2005:VER

[MSW^{+}05] Holger Meuss, Klaus U. Schulz, Felix Weigel, Simone Leonardi, and François Bry. Visual exploration and retrieval of XML document collections with the generic system X^{2}. International Journal on Digital Libraries, 5(1):3–17, March 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0109-5.

Meghini:2010:DMD

[MSY10] Carlo Meghini, Nicolas Spyratos, and Jitao Yang. A data model for digital libraries. International Journal on Digital Libraries, 11(1):41–56, March 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-011-0064-x.

MacIntyre:2000:NPD

[MT00] Ross MacIntyre and Simon Tanner. Nature: a prototype digital archive. International Journal on Digital Libraries, 3(1):67–76, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00021476.

Marketakis:2009:DMD

[MT09] Yannis Marketakis and Yannis Tzitzikas. Dependency management for digital preservation using Semantic Web technologies. International Journal on Digital Libraries, 10 (4):159–177, December 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-010-0058-0.

Mayr:2016:RAK

[MTC^{+}16] Philipp Mayr, Douglas Tudhope, Stella Dextre Clarke, Marcia Lei Zeng, and Xia Lin. Recent applications of Knowledge Organization Systems: introduction to a special issue. International Journal on Digital Libraries, 17(1):1–4, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0167-x.

Makela:2017:WAC

[MTLH17] Eetu Mäkelä, Juha Törnroos, Thea Lindquist, and Eero Hyvönen. WW1LOD: an application of CIDOC–CRM to World War 1 linked data. International Journal on Digital Libraries, 18(4):333–343, November 2017. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-016-0186-2.

Micsik:2015:EPM

[MTT15] András Micsik, Sándor Turbucz, and Zoltán Tóth. Exploring publication metadata graphs with the LODmilla browser and editor. International Journal on Digital Libraries, 16(1):15–24, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0130-2.

Ngimwa:2013:DRL

[NA13] Pauline Ngimwa and Anne Adams. The different roles of ‘Design Process Champions’ for digital libraries in African higher education. International Journal on Digital Libraries, 13(3-4):119–133, September 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-013-0104-9.

Nanba:2008:AEC

[NAO08] Hidetsugu Nanba, Natsumi Anzen, and Manabu Okumura. Automatic extraction of citation information in Japanese patent applications. International Journal on Digital Libraries, 9(2):151–161, November 2008. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-008-0045-x.

Neuhold:2004:DLE

[NF04] Erich Neuhold and Ingo Frommholz. Digital libraries as experienced by the editors of the journal. International Journal on Digital Libraries, 4(1):1–2, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0051-y.

Niccolucci:2016:RGP

[NH16] Franco Niccolucci and Sorin Hermon. Representing gazetteers and period thesauri in four-dimensional space–time. International Journal on Digital Libraries, 17 (1):63–69, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0159-x.

Niccolucci:2017:ERC

[NH17] Franco Niccolucci and Sorin Hermon. Expressing reliability with CIDOC CRM. International Journal on Digital Libraries, 18 (4):281–287, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0195-1.

Noei:2021:QSL

[NHP^{+}21] Ehsan Noei, Tsahi Hayat, Jessica Perrie, Recep Çolak, Yanqi Hao, Shankar Vembu, Kelly Lyons, and Sam Molyneux. A qualitative study of large-scale recommendation algorithms for biomedical knowledge bases. International Journal on Digital Libraries, 22(2):197–215, June 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00300-3.

Niccolucci:2017:DAS

[Nic17a] Franco Niccolucci. Documenting archaeological science with CIDOC CRM. International Journal on Digital Libraries, 18(3): 223–231, September 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0199-x.

Niccolucci:2017:EMF

[Nic17b] Franco Niccolucci. Extending, mapping, and focusing the CIDOC CRM. International Journal on Digital Libraries, 18(4): 251–252, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0198-y.

Nentidis:2021:WAN

[NKTP21] Anastasios Nentidis, Anastasia Krithara, Grigorios Tsoumakas, and Georgios Paliouras. What is all this new MeSH about? International Journal on Digital Libraries, 22 (4):319–337, December 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00304-z.

Nelson:2007:UWI

[NMSK07] Michael L. Nelson, Frank McCown, Joan A. Smith, and Martin Klein. Using the web infrastructure to preserve web pages. International Journal on Digital Libraries, 6 (4):327–349, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0012-y.

Neugebauer:2010:AMC

[NMT10] Tomasz Neugebauer, Corina MacDonald, and Felicity Tayler. Artexte metadata conversion to EPrints: adaptation of digital repository software to visual and media arts documentation. International Journal on Digital Libraries, 11(4):263–277, December 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-011-0077-5.

Nevill-Manning:1999:LGS

[NMWP99] Craig G. Nevill-Manning, Ian H. Witten, and Gordon W. Paynter. Lexically-generated subject hierarchies for browsing large collections. International Journal on Digital Libraries, 2(2-3):111–123, September 1999. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050041.

Nanni:2020:TCE

[NPD20] Federico Nanni, Simone Paolo Ponzetto, and Laura Dietz. Toward comprehensive event collections. International Journal on Digital Libraries, 21(2):215–229, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0246-x.

Novati:2005:AMI

[NPS05] Gianluca Novati, Paolo Pellegri, and Raimondo Schettini. An affordable multispectral imaging system for the digital museum. International Journal on Digital Libraries, 5(3):167–178, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0103-y.

Nikolaou:2000:TAS

[NS00] C. Nikolaou and C. Stephanidis. In the tradition of Alexandrian Scholars: Special issue of the International Journal on Digital Libraries. International Journal on Digital Libraries, 3(1):1–2, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00010717.

Nelson:2006:ISI

[NV06] Michael L. Nelson and Herbert Van de Sompel. IJDL special issue on complex digital objects: Guest editors’ introduction. International Journal on Digital Libraries, 6 (2):113–114, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0127-y.

Ochoa:2009:AEM

[OD09] Xavier Ochoa and Erik Duval. Automatic evaluation of metadata quality in digital repositories. International Journal on Digital Libraries, 10(2-3):67–91, August 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-009-0054-4.

Owen:2000:PTA

[OFM^{+}00] Charles B. Owen, James Ford, Fillia Makedon, Tilmann Steinberg, and Christina Metaxaki-Kossionides. Parallel text alignment. International Journal on Digital Libraries, 3 (1):100–114, July 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050014.

Orgel:2015:MMM

[OHBR15] Thomas Orgel, Martin Höffernig, Werner Bailer, and Silvia Russegger. A metadata model and mapping approach for facilitating access to heterogeneous cultural heritage assets. International Journal on Digital Libraries, 15(2-4):189–207, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0138-2.

Oomen:2013:SBT

[OOKS13] Johan Oomen, Paul Over, Wessel Kraaij, and Alan F. Smeaton. Symbiosis between the TRECVid benchmark and video libraries at the Netherlands Institute for Sound and Vision. International Journal on Digital Libraries, 13(2):91–104, March 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0102-3.

Orio:2012:IFI

[ORR12] Nicola Orio, Andreas Rauber, and David Rizo. Introduction to the focused issue on music DLs. International Journal on Digital Libraries, 12(2-3):51–52, August 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0095-y.

Orio:2009:MTA

[OSCF09] Nicola Orio, Lauro Snidaro, Sergio Canazza, and Gian Luca Foresti. Methodologies and tools for audio digital archives. International Journal on Digital Libraries, 10(4): 201–220, December 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-010-0060-6.

Organisciak:2023:ANN

[OSD23] Peter Organisciak, Benjamin M. Schmidt, and Matthew Durward. Approximate nearest neighbor for long document relationship labeling in digital libraries. International Journal on Digital Libraries, 24(4): 311–325, December 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00354-5.

Ohta:2000:RES

[OTA00] Manabu Ohta, Atsuhiro Takasu, and Jun Adachi. Reduction of expanded search terms for fuzzy English-text retrieval. International Journal on Digital Libraries, 3 (2):140–151, August 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007999900014.

Oh:2016:RDL

[OYP^{+}16] Sanghee Oh, Seungwon Yang, Jeffrey P. Pomerantz, Barbara M. Wildemuth, and Edward A. Fox. Results of a digital library curriculum field test. International Journal on Digital Libraries, 17(4): 273–286, November 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0151-5.

Pomerantz:2009:TDL

[PAM09] Jeffrey Pomerantz, June Abbas, and Javed Mostafa. Teaching digital library concepts using digital library applications. International Journal on Digital Libraries, 10(1):1–13, May 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-008-0049-6.

Papazoglou:2006:JSI

[Pap06] Michael P. Papazoglou. JDL special issue on service-oriented computing: advanced user-centered concepts. International Journal on Digital Libraries, 6(3):233–234, June 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-006-0001-6.

Parker:2004:CDI

[Par04] Julie J. S. Parker. Commercial digital image libraries, digital images and digital discontent. International Journal on Digital Libraries, 4(2):124–136, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0075-3.

Pinto:2020:APS

[PB20] José María González Pinto and Wolf-Tilo Balke. Assessing plausibility of scientific claims to support high-quality content in digital collections. International Journal on Digital Libraries, 21(1):47–60, March 2020. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0256-8.

Pertsas:2017:SOM

[PC17] Vayianos Pertsas and Panos Constantopoulos. Scholarly ontology: modelling scholarly practices. International Journal on Digital Libraries, 18(3):173–190, September 2017. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-016-0169-3.

Pontes:2022:CMM

[PCDM^{+}22a] Elvys Linhares Pontes, Luis Adrián Cabrera-Diego, Jose G. Moreno, Emanuela Boros, Ahmed Hamdi, Antoine Doucet, Nicolas Sidere, and Mickaël Coustaty. Correction to: MELHISSA: a multilingual entity linking architecture for historical press articles. International Journal on Digital Libraries, 23(2):161–165, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00320-z. See

[PCDM^{+}22b].

Pontes:2022:MME

[PCDM^{+}22b] Elvys Linhares Pontes, Luis Adrián Cabrera-Diego, Jose G. Moreno, Emanuela Boros, Ahmed Hamdi, Antoine Doucet, Nicolas Sidere, and Mickaël Coustaty. MEL-HISSA: a multilingual entity linking architecture for historical press articles. International Journal on Digital Libraries, 23 (2):133–160, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00319-6. See

correction [PCDM^{+}22a].

Palacio:2010:EGI

[PCSH10] Damien Palacio, Guillaume Cabanac, Christian Sallaberry, and Gilles Hubert. On the evaluation of Geographic Information Retrieval systems. International Journal on Digital Libraries, 11(2):91–109, June 2010. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-011-0070-z.

Page:2019:GEI

[PD19] Kevin R. Page and J. Stephen Downie. Guest editors’ introduction to the special issue on digital libraries for musicology. International Journal on Digital Libraries, 20(1):1–2, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-019-00268-1.

Proctor:2015:DFS

[PEC^{+}15] James D. Proctor, Kristen Eshleman, Tim Chartier, Lora Taub-Pervizpour, Kristin Bott, Juliane L. Fry, Chris Koski, and Tony Moreno. Digital field scholarship and the liberal arts: results from a 2012–13 sandbox. International Journal on Digital Libraries, 16(1):5–13, May 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0126-y.

Palaiologk:2012:ABC

[PETS12] Anna S. Palaiologk, Anastasios A. Economides, Heiko D. Tjalsma, and Laurents B. Sesink. An activity-based costing model for long-term preservation and dissemination of digital research data: the case of DANS. International Journal on Digital Libraries, 12(4):195–214, September 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0092-1.

Pentaris:2004:MO

[PI04] Fragkiskos Pentaris and Yannis E. Ioannidis. Mapping objects. International Journal on Digital Libraries, 4(1):52–55, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-003-0076-2.

Prasad:2018:NPD

[PKK18] Animesh Prasad, Manpreet Kaur, and Min-Yen Kan. Neural ParsCit: a deep learningbased reference string parser. International Journal on Digital Libraries, 19(4): 323–337, November 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0242-1.

Peponakis:2019:EMP

[PMKD19] Manolis Peponakis, Anna Mastora, Sarantos Kapidakis, and Martin Doerr. Expressiveness and machine processability of Knowledge Organization Systems (KOS): an analysis of concepts and relations. International Journal on Digital Libraries, 20(4): 433–452, December 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-019-00269-0.

Paraizo:2020:PED

[PP20] Tiago Antônio Paraizo and Denilson Alves Pereira. PVAF: an environment for disambiguation of scientific publication venues. International Journal on Digital Libraries, 21(4):407–421, December 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00289-1.

Purcell:1997:DEC

[PRS97] Gretchen P. Purcell, Glenn D. Rennels, and Edward H. Shortliffe. Development and evaluation of a context-based document representation for searching the medical literature. International Journal on Digital Libraries, 1(3):288–296, December 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990050023.

Papachristopoulos:2019:DSI

[PTB^{+}19] Leonidas Papachristopoulos, Giannis Tsakonas, Moses Boudourides, Michalis Sfakakis, Nikos Kleidis, Sergios Lenis, and Christos Papatheodorou. Discovering the structure and impact of the digital library evaluation domain. International Journal on Digital Libraries, 20(2):125–141, June 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0222-x.

Pouchard:2005:DGD

[PWB05] Line Pouchard, Andrew Woolf, and David Bernholdt. Data Grid discovery and Semantic Web technologies for the earth sciences. International Journal on Digital Libraries, 5(2):72–83, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0085-9.

Patel:2005:MRD

[PWM^{+}05] Manjula Patel, Martin White, Nicholaos Mourkoussis, Krzysztof Walczak, Rafal Wojciechowski, and Jacek Chmielewski. Metadata requirements for digital museum environments. International Journal on Digital Libraries, 5(3):179–192, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0104-x.

Rigaux:2012:DIN

[RAA^{+}12] P. Rigaux, L. Abrouk, H. Audéon, N. Cullot, C. Davy-Rigaux, Z. Faget, E. Gavignet, D. Gross-Amblard, A. Tacaille, and V. Thion-Goasdoué. The design and implementation of Neuma, a collaborative Digital Scores Library. International Journal on Digital Libraries, 12(2-3):73–88, August 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0089-9.

Rademaker:2015:LOD

[RBdP^{+}15] Alexandre Rademaker, Dário Augusto Borges Oliveira, Valeria de Paiva, Suemi Higuchi, Asla Medeiros e Sá, and Moacyr Alvim. A linked open data architecture for the historical archives of the Getulio Vargas Foundation. International Journal on Digital Libraries, 15(2-4):153–167, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0147-1.

Rezgui:2004:SPP

[RBE04] Abdelmounaam Rezgui, Athman Bouguettaya, and Mohamed Eltoweissy. SemWebDL: a privacy-preserving Semantic Web infrastructure for digital libraries. International Journal on Digital Libraries, 4(3): 171–184, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0081-0.

Rauber:2007:E

[RCD07] Andreas Rauber and Birte Christensen-Dalsgaard. Editorial. International Journal on Digital Libraries, 6(4):293–294, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-007-0027-4.

Reitsma:2010:CHM

[RD10] René F. Reitsma and Anne R. Diekema. Comparison of human and machine-based educational standard assignment networks. International Journal on Digital Libraries, 11(3):209–223, September 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0074-8.

Ryan:2015:LDA

[RGÓ^{+}15] Catherine Ryan, Rebecca Grant, Eoghan Ó Carragáin, Sandra Collins, Stefan Decker, and Nuno Lopes. Linked data authority records for Irish place names. International Journal on Digital Libraries, 15 (2-4):73–85, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0129-8.

Ross:2005:PRS

[RH05] Seamus Ross and Margaret Hedstrom. Preservation research and sustainable digital libraries. International Journal on Digital Libraries, 5(4):317–324, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0099-3.

Rizo:2019:MBS

[RM19] David Rizo and Alan Marsden. An MEIbased standard encoding for hierarchical music analyses. International Journal on Digital Libraries, 20(1):93–105, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0262-x.

Ronzino:2016:CBC

[RNFD16] Paola Ronzino, Franco Niccolucci, Achille Felicetti, and Martin Doerr. CRM ba a CRM extension for the documentation of standing buildings. International Journal on Digital Libraries, 17(1):71–78, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0160-4.

Ronzino:2017:HCC

[Ron17] Paola Ronzino. Harmonizing the CRMba and CRMarchaeo models. International Journal on Digital Libraries, 18(4):253–261, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-016-0193-3.

Risse:2015:IFI

[RPM^{+}15] Thomas Risse, Livia Predoiu, Annett Mitschick, Andreas Nürnberger, and Seamus Ross. Introduction to the focussed issue on Semantic Digital Archives. International Journal on Digital Libraries, 15(2-4):

71–72, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0148-0.

Richardson:2008:KDD

[RSF08] W. Ryan Richardson, Venkat Srinivasan, and Edward A. Fox. Knowledge discovery in digital libraries of electronic theses and dissertations: an NDLTD case study. International Journal on Digital Libraries, 9(2): 163–171, November 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-008-0046-9.

Rahat:2018:OIE

[RT18] Mahmoud Rahat and Alireza Talebpour. Open information extraction as an intermediate semantic structure for Persian text summarization. International Journal on Digital Libraries, 19(4):339–352, November 2018. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-018-0244-z.

Reiterer:2005:ICB

[RTM05] Harald Reiterer, Gabriela Tullius, and Thomas M. Mann. INSYDER: a contentbased visual-information-seeking system for the Web. International Journal on Digital Libraries, 5(1):25–41, March 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0111-y.

Reilly:2006:MCP

[RWS06] William Reilly, Robert Wolfe, and MacKenzie Smith. MIT’s CWSpace project: packaging metadata for archiving educational content in DSpace. International Journal on Digital Libraries, 6(2):139–147, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0131-2.

Rodrigues:2020:OMM

[RZ20] Marcos Wander Rodrigues and Luis Enrique Zárate. OrgBR-M: a method to assist in organizing bibliographic material based on formal concept analysis — a case study in educational data mining. International Journal on Digital Libraries, 21(4): 423–448, December 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00290-8.

Sumner:2005:LLG

[SAB^{+}05] Tamara Sumner, Faisal Ahmad, Sonal Bhushan, Qianyi Gu, Francis Molina, Stedman Willard, Michael Wright, Lynne Davis, and Greg Janée. Linking learning goals and educational resources through interactive concept map visualizations. International Journal on Digital Libraries, 5(1):18– 24, March 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-004-0112-x.

Saeednia:2000:HMB

[Sae00] Shahrokh Saeednia. How to maintain both privacy and authentication in digital libraries. International Journal on Digital Libraries, 2(4):251–258, May 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/PL00021469.

Stephanidis:2000:UID

[SAPN00] C. Stephanidis, D. Akoumianakis, A. Paramythis, and C. Nikolaou. User interaction in digital libraries: coping with diversity through adaptation. International Journal on Digital Libraries, 3(2):185–205, August 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007999900020.

Steiner:2014:EDH

[SAS^{+}14] Christina M. Steiner, Maristella Agosti, Mark S. Sweetnam, Eva-C. Hillemann, Nicola Orio, Chiara Ponchia, Cormac Hampson, Gary Munnelly, Alexander Nussbaumer, Dietrich Albert, and Owen Conlan. Evaluating a digital humanities research environment: the CULTURA approach. International Journal on Digital Libraries, 15 (1):53–70, November 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0127-x.

Sanderson:2012:SCM

[SASV12] Robert Sanderson, Benjamin Albritton, Rafael Schwemmer, and Herbert Van de Sompel. SharedCanvas: a collaborative model for digital facsimiles. International Journal on Digital Libraries, 13(1):3–16, December 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-012-0098-8.

Satoh:2006:LBS

[Sat06] Ichiro Satoh. Location-based services in ubiquitous computing environments. International Journal on Digital Libraries, 6 (3):280–291, June 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-006-0006-1.

Stelmaszewska:2004:PDC

[SB04] Hanna Stelmaszewska and Ann Blandford. From physical to digital: a case study of computer scientists’ behaviour in physical libraries. International Journal on Digital Libraries, 4(2):82–92, October 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0072-6.

Sautter:2014:IBR

[SB14] Guido Sautter and Klemens Böhm. Improved bibliographic reference parsing based on repeated patterns. International Journal on Digital Libraries, 14(1-2):59–80, April 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-014-0110-6.

Smith:2023:ADE

[SB23] Neel Smith and Christopher Blackwell. Analytical developments for the Homer Multitext: palaeography, orthography, morphology, prosody, semantics. International Journal on Digital Libraries, 24(3): 179–184, September 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00380-3.

Skogsrud:2004:TNS

[SBC04] Halvard Skogsrud, Boualem Benatallah, and Fabio Casati. A trust negotiation system for digital library Web services. International Journal on Digital Libraries, 4 (3):185–207, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-004-0083-y.

Silvello:2017:SRE

[SBF^{+}17] Gianmaria Silvello, Georgeta Bordea, Nicola Ferro, Paul Buitelaar, and Toine Bogers. Semantic representation and enrichment of information retrieval experimental data. International Journal on Digital Libraries, 18 (2):145–172, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0172-8.

Smeaton:2005:PRS

[SC05] Alan F. Smeaton and Jamie Callan. Personalisation and recommender systems in digital libraries. International Journal on Digital Libraries, 5(4):299–308, August 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0100-1.

Seales:2023:SST

[SC23] W. Brent Seales and Christy Y. Chapman. From stone to silicon: technical advances in epigraphy. International Journal on Digital Libraries, 24(2):129–138, June 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-023-00362-5.

Schmiede:2004:ADL

[Sch04] Rudi Schmiede. Advancing digital libraries in Germany and creating distributed scientific resources — the telos working group at Darmstadt University of Technology. International Journal on Digital Libraries, 4 (1):48–51, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0056-6.

Steinhauer:2014:EDB

[SDLÅ14] Jeremy Steinhauer, Lois M. L. Delcambre, Marianne Lykke, and Marit Kristine Ådland. Evaluating distance-based clustering for user (browse and click) sessions in a domain-specific collection. International Journal on Digital Libraries, 14(3-4):167–179, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0117-z.

Shipman:2015:IFI

[SF15] Frank Shipman and Richard Furuta. Introduction to the focused issue of awardnominated papers from JCDL 2013. International Journal on Digital Libraries, 16(2): 89, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0149-z.

Saier:2022:CLC

[SFT22] Tarek Saier, Michael Färber, and Tornike Tsereteli. Cross-lingual citations in English papers: a large-scale analysis of prevalence, usage, and impact. International Journal on Digital Libraries, 23 (2):179–195, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00312-z.

Strubulis:2014:CSP

[SFTD14] Christos Strubulis, Giorgos Flouris, Yannis Tzitzikas, and Martin Doerr. A case study on propagating and updating provenance information using the CIDOC CRM. International Journal on Digital Libraries, 15 (1):27–51, November 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0125-z.

Santana:2015:CDS

[SGLF15] Alan Filipe Santana, Marcos André Gonçalves, Alberto H. F. Laender, and Anderson A. Ferreira. On the combination of domain-specific heuristics for author name disambiguation: the nearest cluster method. International Journal on Digital Libraries, 16(3-4):229–246, September 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-015-0158-y.

Saikh:2022:ESN

[SGM^{+}22] Tanik Saikh, Tirthankar Ghosal, Amish Mittal, Asif Ekbal, and Pushpak Bhattacharyya. ScienceQA: a novel resource for question answering on scholarly articles. International Journal on Digital Libraries, 23(3): 289–301, September 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-022-00329-y.

Sable:2000:TBA

[SH00] Carl L. Sable and Vasileios Hatzivassiloglou. Text-based approaches for non-topical image categorization. International Journal on Digital Libraries, 3(3):261–275, October 2000. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s007990000038.

Srinivasan:2005:FOD

[SH05] Ramesh Srinivasan and Jeffrey Huang. Fluid ontologies for digital museums. International Journal on Digital Libraries, 5 (3):193–204, May 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0105-9.

Sumikawa:2021:FSC

[SI21] Yasunobu Sumikawa and Ryohei Ikejiri. Feature selection for classifying multi-labeled past events. International Journal on Digital Libraries, 22(1):63–83, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00293-5.

Song:2009:TAC

[SJ09] Sangchul Song and Joseph JaJa. Techniques to audit and certify the long-term integrity of digital archives. International Journal on Digital Libraries, 10(2-3):123–131, August 2009. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-009-0056-2.

Sumikawa:2021:AHR

[SJ21] Yasunobu Sumikawa and Adam Jatowt. Analyzing history-related posts in Twitter. International Journal on Digital Libraries, 22 (1):105–134, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00296-2.

Sugiyama:2015:CES

[SK15] Kazunari Sugiyama and Min-Yen Kan. A comprehensive evaluation of scholarly paper recommendation using potential citation papers. International Journal on Digital Libraries, 16(2):91–109, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0122-2.

Shaw:1999:IVI

[SKS^{+}99] Christopher D. Shaw, James M. Kukla, Ian Soboroff, David Ebert, Charles K. Nicholas, Amen Zwa, Ethan L. Miller, and D. Aaron Roberts. Interactive volumetric information visualization for document corpus management. International Journal on Digital Libraries, 2(2-3):144–156, September 1999. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050043.

Sugimoto:1998:SCP

[SKT98] Masanori Sugimoto, Norio Katayama, and Atsuhiro Takasu. A system for constructing private digital libraries through information space exploration. International Journal on Digital Libraries, 2(1):54–66, October 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s007990050036.

Smeaton:2004:ECF

[SLM04] Alan F. Smeaton, Hyowon Lee, and Kieran McDonald. Experiences of creating four video library collections with the Físchlár system. International Journal on Digital Libraries, 4(1):42–44, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0065-5.

Suchowolec:2019:EVO

[SLS19] Karolina Suchowolec, Christian Lang, and Roman Schneider. An empirically validated, onomasiologically structured, and linguistically motivated online terminology. International Journal on Digital Libraries, 20(3): 253–268, September 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0254-x.

Strle:2012:EDL

[SM12] Gregor Strle and Matija Marolt. The Ethno-Muse digital library: conceptual representation and annotation of ethnomusicological materials. International Journal on Digital Libraries, 12(2-3):105–119, August 2012. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-012-0086-z.

Sumikawa:2021:MGB

[SM21] Yasunobu Sumikawa and Tatsurou Miyazaki. Multilabel graph-based classification for missing labels. International Journal on Digital Libraries, 22(1):85–104, March 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00295-3.

Shum:2000:SOB

[SMD00] Simon Buckingham Shum, Enrico Motta, and John Domingue. ScholOnto: an ontology-based digital library server for research documents and discourse. International Journal on Digital Libraries, 3(3): 237–248, October 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990000034.

Smith:1997:MID

[Smi97] Terence R. Smith. Meta-information in digital libraries. International Journal on Digital Libraries, 1(2):105–107, September 1997. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s007990050007.

Skevakis:2014:MMI

[SMK^{+}14] Giannis Skevakis, Konstantinos Makris, Varvara Kalokyri, Polyxeni Arapi, and Stavros Christodoulakis. Metadata management, interoperability and Linked Data publishing support for Natural History Museums. International Journal on Digital Libraries, 14 (3-4):127–140, August 2014. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-014-0114-2.

Salatino:2022:CCS

[SOM22] Angelo Salatino, Francesco Osborne, and Enrico Motta. CSO Classifier 3.0: a scalable unsupervised method for classifying documents in terms of research topics. International Journal on Digital Libraries, 23 (1):91–110, March 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-021-00305-y.

Savino:2004:EDL

[SP04] Pasquale Savino and Carol Peters. ECHO: a digital library for historical film archives. International Journal on Digital Libraries, 4(1):3–7, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0062-8.

Saleh:2002:AHY

[SPLF02] K. Saleh, R. L. Probert, W. Li, and W. Fong. An approach for high-yield requirements capture for e-commerce and its application. International Journal on Digital Libraries, 3(4):302–308, May 2002. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990100041.

Suleman:2008:LCB

[SPO08] Hussein Suleman, Christopher Parker, and Muammar Omar. Lightweight componentbased scalability. International Journal on Digital Libraries, 9(2):115–124, November 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-008-0047-8.

Shaw:2016:SOD

[SRGK16] Ryan Shaw, Adam Rabinowitz, Patrick Golden, and Eric Kansa. A sharingoriented design strategy for networked knowledge organization systems. International Journal on Digital Libraries, 17(1):49–61, March 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-015-0164-0.

Stone:1998:SCP

[SS98] Harold S. Stone and Talal G. Shamoon. Search-by-content of partially occluded images. International Journal on Digital Libraries, 1(4):329–343, March 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050027.

Saleh:2006:DIG

[SS06] Bassam Saleh and Balqies Sadoun. Design and implementation of a GIS system for planning. International Journal on Digital Libraries, 6(2):210–218, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-005-0117-0.

Seara:2012:EOP

[SSB^{+}12] Everton F. R. Seára, Marcos S. Sunye, Luis C. E. Bona, Tiago Vignatti, Andre L. Vignatti, and Anne Doucet. Extending OAI– PMH over structured P2P networks for digital preservation. International Journal on Digital Libraries, 12(1):13–26, July 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-012-0080-5.

Senoo:2024:SPR

[SSK^{+}24] Ko Senoo, Yohei Seki, Wakako Kashino, Atsushi Keyaki, and Noriko Kando. Stance prediction with a relevance attribute to political issues in comparing the opinions of citizens and city councilors. International Journal on Digital Libraries, 25 (1):75–91, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-024-00396-3.

Schmidt:1997:LAR

[SSNM97] Joachim W. Schmidt, Gerald Schröder, Claudia Niederée, and Florian Matthes. Linguistic and architectural requirements for personalized digital libraries. International Journal on Digital Libraries, 1(1):89–104, April 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s007990050006.

Straccia:2004:OCV

[ST04] Umberto Straccia and Costantino Thanos. An open collaborative virtual archive environment. International Journal on Digital Libraries, 4(1):23–24, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0063-7.

Sikdar:2020:ESP

[STM^{+}20] Sandipan Sikdar, Paras Tehria, Matteo Marsili, Niloy Ganguly, and Animesh Mukherjee. On the effectiveness of the scientific peer-review system: a case study of the Journal of High Energy Physics. International Journal on Digital Libraries, 21 (2):93–107, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-018-0247-9.

Stolze:2000:SNE

[Sto00] Markus Stolze. Soft navigation in electronic product catalogs. International Journal on Digital Libraries, 3(1):60–66, July 2000. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/PL00021475.

Styliaras:2007:ASM

[STP07] Georgios D. Styliaras, Georgios K. Tsolis, and Chris M. Papaterpos. AssetCollector: a system for managing collections of cultural assets. International Journal on Digital Libraries, 8(1):61–78, November 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0028-3.

Samar:2018:QRB

[STvO^{+}18] Thaer Samar, Myriam C. Traub, Jacco van Ossenbruggen, Lynda Hardman, and Arjen P. de Vries. Quantifying retrieval bias in Web archive search. International Journal on Digital Libraries, 19(1):57–75, March 2018. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-017-0215-9.

Seki:2022:CBF

[SZOS22] Yohei Seki, Kangkang Zhao, Masaki Oguni, and Kazunari Sugiyama. CNN-based framework for classifying temporal relations with question encoder. International Journal on Digital Libraries, 23(2):167–177, June 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-021-00310-1.

Tahmasebi:2015:VOC

[TBC^{+}15] Nina Tahmasebi, Lars Borin, Gabriele Capannini, Devdatt Dubhashi, Peter Exner, Markus Forsberg, Gerhard Gossen, Fredrik D. Johansson, Richard Johansson, Mikael Kågebäck, Olof Mogren, Pierre Nugues, and Thomas Risse. Visions and open challenges for a knowledge-based culturomics. International Journal on Digital Libraries, 15(2-4):169–187, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-015-0139-1.

Thiel:2004:CCS

[TBF^{+}04] Ulrich Thiel, Holger Brocks, Ingo Frommholz, Andrea Dirsch-Weigand, Jürgen Keiper, Adelheit Stein, and Erich J. Neuhold. COLLATE

— a collaboratory supporting research on historic European films. International Journal on Digital Libraries, 4(1):8–12, August 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-003-0069-1.

Tsiknakis:1997:IAP

[TCK^{+}97] M. Tsiknakis, C. E. Chronaki, S. Kapidakis, C. Nikolaou, and S. C. Orphanoudakis. An integrated architecture for the provision of health telematic services based on digital library technologies. International Journal on Digital Libraries, 1(3): 257–277, December 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050021.

Tsakonas:2020:CRT

[TK20] Giannis Tsakonas and Jaap Kamps. Current research on theory and practice of digital libraries: best papers from TPDL 2017. International Journal on Digital Libraries, 21(1):1–3, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-020-00278-4.

Tzitzikas:2015:ADI

[TKM15] Yannis Tzitzikas, Yannis Kargakis, and Yannis Marketakis. Assisting digital interoperability and preservation through advanced dependency reasoning. International Journal on Digital Libraries, 15(2-4):103–127, April 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-014-0131-1.

Tian:2002:GBA

[TLJ^{+}02] Zengping Tian, Hongjun Lu, Wenyun Ji, Aoying Zhou, and Zhong Tian. An ngram-based approach for detecting approximately duplicate database records. International Journal on Digital Libraries, 3 (4):325–331, May 2002. CODEN ????

ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990100044.

Thanos:2022:EAA

[TMBC22] Costantino Thanos, Carlo Meghini, Valentina Bartalesi, and Gianpaolo Coro. An exploratory approach to archaeological knowledge production. International Journal on Digital Libraries, 23(3):231–239, September 2022. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-022-00324-3.

Tsoi:2007:DAA

[TMTA07] Ah Chung Tsoi, Jeff McDonell, Andrew Treloar, and Ian Atkinson. Dataset acquisition, accessibility, annotation, e-research technologies (DART) project. International Journal on Digital Libraries, 7(1-2):53–55, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0019-4.

Tahmasebi:2013:AWS

[TNZR13] Nina Tahmasebi, Kai Niklas, Gideon Zenz, and Thomas Risse. On the applicability of word sense discrimination on 201 years of Modern English. International Journal on Digital Libraries, 13(3-4):135–153, September 2013. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-013-0105-8.

Toszegi:2004:JND

[Tós04] Zsuzsanna Tószegi. The John von Neumann Digital Library (www.neumann-haz.hu). International Journal on Digital Libraries, 4(1):13–18, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-003-0077-1.

Tuarob:2015:GTM

[TPMG15] Suppawong Tuarob, Line C. Pouchard, Prasenjit Mitra, and C. Lee Giles. A generalized topic modeling approach for automatic document annotation. International Journal on Digital Libraries, 16(2):111–128, June 2015. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-015-0146-2.

Toepfer:2020:FAA

[TS20] Martin Toepfer and Christin Seifert. Fusion architectures for automatic subject indexing under concept drift. International Journal on Digital Libraries, 21(2):169–189, June 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-018-0240-3.

VandenEynden:2017:ARD

[VC17] Veerle Van den Eynden and Louise Corti. Advancing research data publishing practices for the social sciences: from archive activity to empowering researchers. International Journal on Digital Libraries, 18 (2):113–121, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0177-3.

Velegrakis:2000:ZWD

[VCC00] Yannis Velegrakis, Vassilis Christophides, and Panos Constantopoulos. On Z39.50 wrapping and description logics. International Journal on Digital Libraries, 3(3): 208–220, October 2000. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990000036.

VandeSompel:2008:AFA

[VCH08] Herbert Van de Sompel, Ryan Chute, and Patrick Hochstenbach. The aDORe federation architecture: digital repositories at scale. International Journal on Digital Libraries, 9(2):83–100, November 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-008-0048-7.

Verbert:2008:AGC

[VD08] Katrien Verbert and Erik Duval. ALO-COM: a generic content model for learning objects. International Journal on Digital Libraries, 9(1):41–63, August 2008. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-008-0039-8.

Vassallo:2021:TOC

[VF21] Valentina Vassallo and Achille Felicetti. Towards an ontological cross-disciplinary solution for multidisciplinary data: VI–SEEM data management and the FAIR principles. International Journal on Digital Libraries, 22(3):297–307, September 2021. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-020-00285-5.

Vanderschantz:2021:CQF

[VH21] Nicholas Vanderschantz and Annika Hinze. Children’s query formulation and search result exploration. International Journal on Digital Libraries, 22(4):385–410, December 2021. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-021-00316-
9.

vanKranenburg:2019:DSC

[vKdBV19] Peter van Kranenburg, Martine de Bruin, and Anja Volk. Documenting a song culture: the Dutch Song Database as a resource for musicological research. International Journal on Digital Libraries, 20(1):13–23, March 2019. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-017-0228-4.

Velios:2017:SCD

[VM17] Athanasios Velios and Aurelie Martin. Offthe-shelf CRM with Drupal: a case study of documenting decorated papers. International Journal on Digital Libraries, 18(4): 321–331, November 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0191-5.

Ven:2008:SIS

[VSVS08] Kris Ven, Kenneth Sörensen, Jan Verelst, and Marc Sevaux. Stimulating information sharing, collaboration and learning in operations research with libOR. International Journal on Digital Libraries, 8(2):79–90, April 2008. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-007-0032-7.

Wang:1999:HSA

[Wan99] Bing Wang. A hybrid system approach for supporting digital libraries. International Journal on Digital Libraries, 2(2-3): 91–110, September 1999. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050040.

Waters:2023:EDI

[Wat23] Donald J. Waters. The emerging digital infrastructure for research in the humanities. International Journal on Digital Libraries, 24(2):87–102, June 2023. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-022-00332-3.

Waugh:2006:DVE

[Wau06] Andrew Waugh. The design of the VERS encapsulated object experience with an archival information package. International Journal on Digital Libraries, 6(2):184–191, April 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-005-0135-y.

Weinstein:1998:COM

[WB98] Peter C. Weinstein and William P. Birmingham. Creating ontological metadata for digital library content and services. International Journal on Digital Libraries, 2 (1):20–37, October 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050034.

Warner:2007:EPA

[WBL^{+}07a] Simeon Warner, Jeroen Bekaert, Carl Lagoze, Xiaoming Liu, Sandy Payette, and Herbert Van de Sompel. Erratum to: Pathways: augmenting interoperability across scholarly repositories. International Journal on Digital Libraries, 6(4):363, July 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0025-6. See

[WBL^{+}07b].

Warner:2007:PAI

[WBL^{+}07b] Simeon Warner, Jeroen Bekaert, Carl Lagoze, Xiaoming Liu, Sandy Payette, and Herbert Van de Warner. Pathways: augmenting interoperability across scholarly repositories. International Journal on Digital Libraries, 7(1-2):35–52, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0016-7. See

erratum [WBL^{+}07a].

Wright:2012:JCD

[WC12] Michael J. Wright and Lillian Cassel. Joint conference on digital libraries (JCDL) 2011. International Journal on Digital Libraries, 13(1):1–2, December 2012. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-012-0099-7.

Witten:2004:TMD

[WDDT04] Ian H. Witten, Katherine J. Don, Michael Dewsnip, and Valentin Tablan. Text mining in a digital library. International Journal on Digital Libraries, 4(1):56–59, August 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-003-0066-4.

Westman:2010:EVV

[Wes10] Stina Westman. Evaluation of visual video summaries: user-supplied constructs and descriptions. International Journal on Digital Libraries, 11(2):125–140, June 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0071-y.

Wang:2006:MAR

[WGZ^{+}06] James Z. Wang, Kurt Grieb, Ya Zhang, Ching chih Chen, Yixin Chen, and Jia Li. Machine annotation and retrieval for digital imagery of historical materials. International Journal on Digital Libraries, 6 (1):18–29, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0121-4.

Walsh:2020:COM

[WHCF20] David Walsh, Mark M. Hall, Paul Clough, and Jonathan Foster. Characterising online museum users: a study of the National Museums Liverpool museum website.

International Journal on Digital Libraries, 21(1):75–87, March 2020. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0248-8.

White:2018:BWR

[Whi18]  Howard D. White.  Bag of works re-

trieval: TF*IDF weighting of works cocited with a seed. International Journal on Digital Libraries, 19(2-3):139–149, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s00799-017-0217-7.

Wong:1997:IAD

[WHK^{+}97] Stephen T. C. Wong, Kent Soo Hoo, R. C. Knowlton, R. A. Hawkins, K. D. Laxer, Donny Tjandra, and Marco A. Abundo. Issues and applications of digital library technology in biomedical imaging. International Journal on Digital Libraries, 1(3): 209–219, December 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s007990050017.

Watanabe:2024:HLL

[WIMM24] Shingo Watanabe, Hiroyoshi Ito, Masaki Matsubara, and Atsuyuki Morishima. Humanin-the-loop latent space learning for bibliorecord-based literature management. International Journal on Digital Libraries, 25 (1):123–136, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00389-8.

Witt:2007:POP

[Wit07] Michael Witt. Providing an OAI–PMH interface to the Storage Resource Broker with OAISRB. International Journal on Digital Libraries, 7(1-2):57–58, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-007-0017-6.

Weibel:1997:ESS

[WL97] Stuart L. Weibel and Carl Lagoze. An element set to support resource discovery. International Journal on Digital Libraries, 1(2):176–186, September 1997. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s007990050013.

Wang:2004:TWM

[WLC04] Jenq-Haur Wang, Wen-Hsiang Lu, and Lee-Feng Chien. Toward Web mining of crosslanguage query translations in digital libraries. International Journal on Digital Libraries, 4(4):247–257, November 2004. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-004-0091-y.

Weigl:2019:PSA

[WLC^{+}19] David M. Weigl, David Lewis, Tim Crawford, Ian Knopke, and Kevin R. Page. On providing semantic alignment and unified access to music library metadata. International Journal on Digital Libraries, 20 (1):25–47, March 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-017-0223-9.

Wang:2006:DID

[WLS06] Chaokun Wang, Jianzhong Li, and Shengfei Shi. The design and implementation of a digital music library. International Journal on Digital Libraries, 6(1):82–97, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-005-0125-0.

Wolf:2004:LEE

[WPS^{+}04] Patrick Wolf, Wolfgang Putz, Avare’ Stewart, Arnd Steinmetz, Matthias Hemmje, and Erich Neuhold. LectureLounge — experience education beyond the borders of the classroom. International Journal on Digital Libraries, 4(1):39–41, August 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0053-9.

Wenige:2018:RRU

[WR18] Lisa Wenige and Johannes Ruhland. Retrieval by recommendation: using LOD technologies to improve digital library search. International Journal on Digital Libraries, 19 (2-3):253–269, September 2018. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-017-0224-8.

Willkomm:2023:CBU

[WRSB23] Jens Willkomm, Markus Raster, Martin Schäler, and Klemens Böhm. CH-Bench: a user-oriented benchmark for systems for efficient distant reading (design, performance, and insights). International Journal on Digital Libraries, 24(4):243–261, December 2023. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-023-00347-
4.

Wright:2007:CDL

[WSMK07] Michael Wright, Tamara Sumner, Reagan Moore, and Traugott Koch. Connecting digital libraries to eScience: the future of scientific scholarship. International Journal on Digital Libraries, 7(1-2):1–4, October 2007. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-007-0030-9.

Wright:2017:MCE

[WTH17] Daniel G. Wright, Philip Trembath, and Kathryn A. Harrison. Meeting the challenge of environmental data publication: an operational infrastructure and workflow for publishing data. International Journal on Digital Libraries, 18(2):123–132, June 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-016-0176-4.

Wu:2016:SIH

[Wu16] Shao-Chun Wu. Systems integration of heterogeneous cultural heritage information systems in museums: a case study of the National Palace Museum. International Journal on Digital Libraries, 17(4): 287–304, November 2016. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-015-0154-2.

Wang:1998:CBI

[WWFW98] James Ze Wang, Gio Wiederhold, Oscar Firschein, and Sha Xin Wei. Contentbased image indexing and searching using Daubechies’ wavelets. International Journal on Digital Libraries, 1(4):311–328, March 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s007990050026.

Williams:2005:SWW

[WZ05] Hugh E. Williams and Justin Zobel. Searchable words on the Web. International Journal on Digital Libraries, 5(2):99–105, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

                                        https://link.springer.com/article/10.
                                        1007/s00799-003-0050-z.

Xu:2024:STI

[XI24] Zhewei Xu and Mizuho Iwaihara. Selftraining involving semantic-space finetuning for semi-supervised multi-label document classification. International Journal on Digital Libraries, 25(1):25–39, March 2024. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-023-00355-4.

Yang:2017:CKM

[YG17] Bin Yang and Jean-Gabriel Ganascia. Creating knowledge maps using Memory Island. International Journal on Digital Libraries, 18(1):41–57, March 2017. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-016-0196-0.

Yada:2020:ITM

[YKP20] Shuntaro Yada, Kyo Kageura, and Cecile Paris. Identification of tweets that mention books. International Journal on Digital Libraries, 21(3):265–287, September 2020. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-019-00273-
4.

Yang:2005:AID

[YRW^{+}05] Yanyan Yang, Omer F. Rana, David W. Walker, Roy Williams, Christos Georgousopoulos, Massimo Caffaro, and Giovanni Aloisio. An agent infrastructure for ondemand processing of remote-sensing archives. International Journal on Digital Libraries, 5(2):120–132, April 2005. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-003-0054-8.

Yao:2006:BES

[YSZ06] Shu-Yuen D. Yao, Cyrus Shahabi, and Roger Zimmermann. BroadScale: Efficient scaling of heterogeneous storage systems. International Journal on Digital Libraries, 6 (1):98–111, February 2006. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-005-0118-z.

Zhang:2010:SVD

[ZBS^{+}10] Li Zhang, Michael Bieber, Min Song, Vincent Oria, and David E. Millard. Supplementing virtual documents with just-in-time hypermedia functionality. International Journal on Digital Libraries, 11(3): 155–168, September 2010. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-011-0065-9.

Zhou:2004:NFT

[ZG04] Shuigeng Zhou and Jihong Guan. A new full-text indexing model with low space overhead for Chinese text retrieval. International Journal on Digital Libraries, 4(4): 272–282, November 2004. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

                                        tronic). URL https://link.springer.com/
                                        article/10.1007/s00799-004-0089-5.

Zimmerman:2007:MAU

[Zim07] Ann Zimmerman. Not by metadata alone: the use of diverse forms of knowledge to locate data for reuse. International Journal on Digital Libraries, 7(1-2):5–16, October 2007. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

                                        (electronic). URL https://link.springer.
                                        com/article/10.1007/s00799-007-0015-8.

Zhang:1998:SMP

[ZJ98] Aidong Zhang and Thomas V. Johnson. Scheduling multimedia presentations in educational digital libraries. International Journal on Digital Libraries, 1(4):386–400, March 1998. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (electronic). URL

https://link.springer.com/article/10.
1007/s007990050031.

Zeng:2019:KOS

[ZM19] Marcia Lei Zeng and Philipp Mayr. Knowledge Organization Systems (KOS) in the Semantic Web: a multi-dimensional review. International Journal on Digital Libraries, 20(3):209–230, September 2019. CODEN ???? ISSN 1432-1300 (print), 1432-5012 (elec-

tronic). URL https://link.springer.com/
article/10.1007/s00799-018-0241-2.

Zhu:2012:DRP

[ZMM^{+}12] Bing Zhu, Richard Marciano, Reagan Moore, Laurin Herr, and Jurgen Schulze. Digital repository: preservation environment and policy implementation. International Journal on Digital Libraries, 12(1):41–49, July 2012. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/s00799-012-0082-3.

Zhu:2000:ALS

[ZYZ^{+}00] Huican Zhu, Tao Yang, Qi Zheng, David Watson, Oscar H. Ibarra, and Terence Smith. Adaptive load sharing for clustered digital library servers. International Journal on Digital Libraries, 2(4):225–235, May 2000. CO-DEN ???? ISSN 1432-1300 (print), 1432-5012

(electronic). URL https://link.springer.
com/article/10.1007/PL00021467.